A Complete Bibliography of Publications in The Computer Journal: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

30 May 2024 Version 2.07

Title word cross-reference

A [PW69]. Ax = b [Loi68]. Ax = λBx [Fet65, PW69]. B [PW69]. e [Ber69, Sal68]. LL^{T} [OK63]. LR [Fra61, Wil65c]. N [Phi67c, CD69, MLB63]. π [Tee67c]. QR [Fra61, Fra62, GRB68, OK63, Wat08, Wil65d, Wil65c]. σ_{1} [DM64, Eva64, Lyn64]. τ [Riv68, Wra66]. x [Loi68]. y′ = f(x)y + g(x) [Day65].

-dimensional [Phi67c]. -method [Riv68]. -methods [Wra66]. -ordered [Eva64, Lyn64].

1 [Ano67g, EAL66, Ful66a, TS68]. 100s [Wil67c]. 13 [HJBH67]. 19 [Rud67b]. 1905 [Ham66b]. 1960 [Her61b]. 1960-61 [Ano61a]. 1962 [Ano62c, Ano62h, Ano62j, BT63, GT63b, Hou63b, Nee63b, PCSG63, RP63]. 1963 [BT63, PCSG63]. 1967 [Ano67g].

2 [BH63, Boo79, Gea64a, Ove67c, PHJ67]. 26 [Hil69, Woo70].

3 [Pal62, RP63]. 34 [Ham69d]. 35 [Cob69, PP69].

4 [BW69a]. 40 [Sco69]. 42 [Sal70]. 44 [Fie70, Red71]. 4th [Ano62c].

60 [BBG^{+}63, Dij62, Dun62b, Dun63, GS63, Hoc62, McE67, Ryd64, Sam63c, Woo60].

7 [PHJ67]. 7th [Ano62c].

802 [Coo60b].

9 [Dun63].

A.D.I. [FM66]. A.D.P. [Lav64, Pol60, Str61]. Ability [KV63]. ABS12 [Hoc62]. abscissas [OSS69]. Accelerated [Aha65]. Accelerating [Wri63, Car61]. acceleration [Gou68]. Acceptance [BS61, Pai65]. access [Fra68, WR67, Wil67a, WN68]. according [Boo67g, Boo67h, Hil69, Woo70]. Accounting [Cas63, Moo63, Hop63, Wil63b]. accumulator [Col63a]. Accuracy [MP62]. activity [Tur68]. Adaptation [Cha62]. Adaptive [AA68]. Additive [MP68]. Address [Weg62, Col63a, Gea64a, Las68, Mic64]. Addressing [EAL66, Elc65]. ADI [ML65]. adjustment [Hio67]. Administration [Sam65, WKRJ60]. Adopts [Hin60]. Advanced [Mor69, Pri64]. adversity [Mus66b]. Advertisers [Ano65e, Ano66f]. aeroplane [Tay61a]. against [OL68]. age [Geo63]. Ageing [DGM68]. Agricultural [AS63]. Aid [BD61, FFMR66, Bro69b, HK67]. Aided [Kin69b, Gol69]. aids [Geo63]. Air [Dav63, Lan65]. Aitken [Eva63, Eva64]. ALAM [d’I69]. Algebra [Bar67, Cli62c, BBB68, Bur67]. Algebraic [FR63, Hus61, Ull67, Wil66a, Wil60, d’I69, Boh68, Ros67]. Algébraiques [Wil62b, Wil63a]. ALGOL [Dun62a, BBG^{+}63, Bar64, Dun63, FK68, GS63, Hig63, Hoa63b, Hoc62, HB67, McE67, Sam63c, Tay64, Tru66, Wat63, Bar65b, Dij62, Dun62b, Hoa62b, Mil68b, Pul64, Ryd64, Sam65, Sip69, Woo60, Zar69, Hoa64]. Algorithm [Ada69, Bae62, Boo67k, Boo67e, Boo67g, Boo67i, Boo67j, Boo67b, Boo67h, Boo67a, Boo67c, Boo67d, Boo67f, Boo68a, Bro69c, Dak65, Dav68, Day69, GRB68, Ham67a, HP67, HP79, Le 69, Mak68a, Ove67a, Ove67b, PP79, Pit67, Pre68, Rud67a, Spä68, Spä69b, Spä69a, Til69, Wat08, Wat11, Wat67, Woo69d, Alm66, AM65, BH68, Bar65a, BP69, Deu69, Flo64, Kni68, May66, MB66, OW69, Rob63b, Rya66, Tem69b, Wil65d, Yul68, Boo68b, Boo79, Cob69, Fie70, Ham69d, HJBH67, Ove67c, PHJ67, PP69, Red71, Rud67b, Sal70, Sco69]. algorithmic [BBG^{+}63]. algorithmically [DP66]. Algorithms [Ano66a, Ano67a, Ano67b, Ano67c, Ano68a, Ano68b, Ano68c, Ano68d, Ano69a, Ano69b, Ano69c, Ano69d, Boo67l, Boo67m, Boo67n, Ham66a, Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c, Phi68a, CP66b, PR64, RS69, Wil65c, Ano66i, Ham66c, Hil69, Woo70].

Allen [Nee63b]. allied [Cur65, Scr64b]. Allocation [IJ62, Lam63]. Almost [Fos68a]. ALP [CW62]. Alphacode [DH60]. alphanumeric [Tho64a]. Alston [Ano69e]. alternate [Jon66]. alternating [Eva65a, FM64, Had68]. alternative [Bak64]. among [Kin69a]. Analogue [Bis66, Fra61, HB66, Tho61, Tho68c, Tiz60, BBS66, Bea68, BW69b, Cra67, DP66, Fir61, Lan65, Lew65a, RWT64, RW68, Tho68a, Car63, Clu63a, Her61b]. analyse [LW66a]. Analyser [Sim69, Mar69]. analysi [LW65]. Analysing [Jar68]. Analysis [Bae66a, Boo68b, Cle63, Dre66, Gri62, Hio68b, KP60, Kin69b, Lin68, Liv60, OW66, Osb66, Sla62b, Wil67c, YS60, Bae66c, FK68, Gen69, Osb64a, Pen66, SW68, Smi66c, TW64, YS61, YGS63, Cle68]. analyst [AT70, Fel69, Fel70]. Analysts [Tra66]. Analytic [Ros67, Sch65, Ede65]. analyzed [Fos64]. analyzers [CND66, Wal64]. angular [Ede65]. Annual [Ano61a, RP63]. Appeal [Sam68a]. appears [OM64]. Appendix [Ano62a, Ano62b]. applicable [Woo69c]. Application [Ake66, CR63, Ras66, Tho62, Wis64, ABN61, Bar65c, Col63b, For66, FGR67, May69, OS65b, OM64, PW67, Smi66c, TS64, TB63, WP67]. application-oriented [For66, FGR67]. Applications [Pai60, FK68, Gea61b, Gos60, RP63]. applied [Eva64, Lew65a, Lyn64]. applying [Pla61]. appreciation [MM69]. Approach [Ano63c, Ano67d, Gre67, Ull65, Ano63a, For66, FGR67, GlM68, Hum62, Lar64, May64, MW68, Ros66b, Thr69]. approaches [JJ67]. Approximation [CP66a, OW67, Tee63c, Ash68, BY66, Cob69, Osb64b, OW69, PP69, Pre68, Riv68, Tem69b, Cle66b, Rei69]. Approximations [Bel63, Eve63, Mil69, CP66b, CO66, Osb63, Osb65, Phi68a, Phi68b, Pow67]. April [Ano62g, Ano62i, Ano62h, Ano62j]. Architect [FFMR66]. Architectural [New66]. arcsin [Mak68a]. Areas [Ada69]. argument [OL68]. Arithmetic [How67b, CL66a, CL66b]. Arithmetical [Ham69d]. Army [Sla60, TS64]. arrangement [Bea65, Gil66, RW67]. array [Boo67g, Boo67h, Boo67f, Hil69, Woo70]. Arrays [Fos68a]. Aspects [Dem60, Dru63]. assembler [Gea64a]. Assembly [BM60, BM61b, Cle67b, LWS69a]. Assessment [BD61, JP69]. Assignment [CP64, Pag63b]. Assisted [Geo66]. associated [GL64]. Association [LW65]. Assurance [Gil64, Sut60]. asymmetric [Duk66b, Jon66, Rac65]. Asymptotic [Cha68b, Jar67b]. asynchronous [Che64]. Atlas [BRC66, BMR67a, CL66b, CP62, Dol66, EAL66, HB67, HPS61, HJW62, KHPS61, McE67, MR67, SW68, d’I69]. Atoms [WJ61]. Attempt [WY62]. Audit [Ano67d, Tho60]. Auditing [Goo64, Tho60, dP60]. auditor [Fri65, dP60]. Augmentation [Bae66b]. Author [Boo67l, Boo67m, Boo67n]. Authorization [TS64]. Autocode [Gib68, And66a, BM63, BRC66, BMR67a, CL66b, CW62, Gea61b, Gea65, Gib61, McE67, YGS63]. Automata [Bel67b, Bel69c, Coo68b, Fos65b]. Automatic [Ano60d, Ano60f, CP61, DN64, Dun67, FR63, Fra67, Gil65, Hin60, Hou63b, Lei62, Pai60, RP63, Wes64, d’A62, Ano60a, GL70, Gib60, JP69, Lyn69, Ros60]. Automation [Ano61b, Ash67b, Bur68, GT63b, Pow65b]. AUTOSTAT [DM60]. Award [Ano69e]. Axially [Abb64, Ran68a].

B [Cle68]. Babbage [OS65a]. Bacchus [Ede63]. background [Rob63b, Spä68]. backing [Cha62]. Backus [Roh68]. backward [Rya66]. Bairstow [Han62, Wat67]. Band [Rei67, PW69]. bandwidth [AM65]. Bank [Hin60]. Banks [Hin61]. Bankwesen [Pow65b]. Bartee [PCSG63]. based [Bar65a, DP66, Ros66b]. Basic [Coo68a, Ham69d, Sol68b, Ull65, YS61, Ili69]. Basis [AS63, Dou65]. batchwise [You60]. BCL [Hou69a, Hou69b]. be [Ano61h, Fin60, Hil67]. benefits [Pai69, WP69]. Bessel [Sto68]. Best [Cob69, Fle68, OW67, PP69, Pre68, Phi68b]. between [Boo67c, Boo67d, Cor65, Nic66, Obr67, Tiz60]. bibliography [DGM68]. Bickley [MM69]. Biharmonic [Gre67, Tee63c, FM64]. binary [Flo64, Hur68]. Biology [Pay66]. Biomedical [Wri66c]. Bishop [Wil67c]. BLM [Ili69]. block [Car66, OS65b]. Blocks [Sam65]. Board [Ano65b, Ano65c, Ano66c, Ano67h, Ano68e, Ano69g, Ano69h]. Book [Alb67, Alb68, Ale69, And66b, Ano60b, Ars65, Ash67a, Bae64b, Bae64a, Bae66b, Bae67, Bar62, Bel67a, Bel68, Bel69a, Bel69b, Box67, Bra63a, BT63, Bri65, Bri66, Bro68, Bur66a, Car63, Cha68a, Che63, Chu69, Cla62, Cle67a, Cle63, Cle65, Cle66b, Cle68, Clu63b, Clu63a, Col64c, Coo69, Cra66a, Cur68, Cur67, Dav67a, Dir68, Dor67, Dor68, Dou69, Dou66, Eve68, Eve69, Fie64, Flo68a, Flo68b, Fox61a, Fox68, Fra69a, Ful66a, Ful66b, Ful67, Gat67, Gea61a, Gea64b, Gea67, Gee66, Gee67, Gil62, Gil63a, Gil69b, Gil68, Goo60, Goo63, GT63a, Gre65, Hag69, Ham66d, Har65a, Har63, Hea61, Her61a, Her61b, Hio68a, Hoa63a, Hoa64, Hop62, Hou63a, Hou63b, How62, How65, How68b, Hug67, Hum67, Ili64, Jon64, Jon65b, Jon69, Ken65, Kir63, LW62, Lan68, Lav68, Lay65, Lev63, Llo68, Llo69]. Book [Mag64, Mag65, Mar62, May69, McM68, Meg62, Mil62, Mil64, Mit69, Mur68, Nee61a, Nee61b, Nee63a, Nee63b, Nel68, New65a, Nix64, Oat63, OS65a, OS67a, OS67b, OS68b, Osb66, Pag63a, Pay66, Per63, PCSG63, Pig64, Pow65b, Pri67, Pri68b, Pri61, Pri63, Rei69, Ric63, RP63, Rob67, Rob63a, Rog65, Ros62, Ros68, Sam63a, Sam68b, Sha67, Sla62a, Smi63a, Spe65, Str62, Str64, Tay61b, Tee63a, Tee67b, Tiz62, Tiz64, Tiz65, Toc64, TW61, Tur66, Vaj60, Vaj64, Vic67a, Vic67b, Wal62, Wat69, Wat68a, Wes64, Wes69, Wil61a, Wil62a, Wil66a, Wil69, Wil62b, Wil63a, Wil67c, Wil64c, Wil66b, Wis61, Wol66, Wol68, Woo62, Woo66a, Wri66c, You63b, You68a, You68b, Sam68a]. Boolean [Mar69]. Boor [GT63b]. both [Bae66c, OSS69, Pen66]. bound [WP67]. boundaries [LW66b]. Boundary [Osb66, AH69, Bic68, Bic69, Fyf69, Has61, JM68, KM66, Osb69, Pri68a]. boundary-value [Pri68a]. Bounds [LM68, Wil61c]. Brain [Ano67l, dF66]. Brains [Goo65]. Britain [TB63]. British [Ano62c, Ano63b, Ano69n, Pla61]. Bureau [Dou65, Lew65b, Kin62]. Business [Che62b, Pai60, Gea61b].

C [PCSG63]. Calculating [Pow64, Pow65a, CP66b, Goo61, Zan67]. Calculation [Eve63, GRB68, Pow68, SD62, Wil65b, And69b, Ber69, May66, Sal68]. calculations [Pal63]. calculus [All66a, Pak67b]. Cambridge [Wil67c, JJ67]. can [Fin60]. Canada [Ano65f]. Capacity [PCSG63]. Car [Lei62]. Cardiff [Ano62c]. Carlo [CR63, KB62]. case [OW69, Tem69b]. Cavity [Gre69]. census [TB63]. Centre [Buc60, TS64]. Certain [Coo66, Lyn65, Bak64, BM61b, EF63b, GM65, Jar66a, Jar67b, MP65a, MP66, Sal70, Spä69b]. Certification [Boo79]. Change [Pap60]. Change-Ringing [Pap60]. Character [CP61, GB61, Hin61, MN61, New61, Pai66, Sha65, Bar65d, Bau65b, cGR61, Ros64b, Str61]. Characterization [Mic63]. characters [Coo68a]. Charles [OS65a]. Chart [OB68a]. charts [How67a]. Chebyshev [Ano64a, Bea65, Cha67, CN63, Coo67, Eg69, Ell63, Fox62, Gou68, Nor64, Oli69, OW67, Scr65, Tee63b, Wol69, Wri64, Wri63, Cle68]. check [Mag64]. Checking [LEN63]. chemical [Cam69, You60]. chemin [Fos64]. chemin-de-fer [Fos64]. choosing [cGR61]. Chromatic [Col67, WP67]. Cie [Wil62b, Wil63a]. Circuits [Bel69c, AA68, BH68, OS65b, Vic65]. circular [MP65a, MP66]. cited [Chu68]. Clason [GT63b]. class [Fox67, Gro68]. Classification [LW67c, Ros64a, WB68, Con66, Hal68, JJ67, LW65, LW66a]. classifications [JS68]. classificatory [LW67a, LW67b, LW68]. Clenshaw [OS68a]. Cleo [Ric64]. Closed [Tha66]. clump [JJ67]. clump-finding [JJ67]. Cluster [SW68]. Clustering [LW67b]. co [RW61]. co-diagonalization [RW61]. COBOL [Cli62a, Gil69a, Hum62, Kin62, Gol64]. Code [CH60, Bar65d, cGR61, Ros64b]. codes [Mag64]. codiagonal [SF61]. Coding [Bra63b, Pai60, Woo66c, Col63a, Ker63b, Tee67a]. Coefficients [SS63a, Cha67, Coo67, Gen69]. Collection [Wri61a]. College [Pri64]. collocation [Wri64]. colouring [Woo69c]. column [Tew67]. Columns [Boo67b]. Combination [Tee63b, Ano64a]. Combinatorial [Lyn65, Mil68a]. combined [Lar64]. comes [Mus67]. Comments [Deu69, Fra70, Kin67b, New61]. commerce [Bag61b]. Commercial [Lew65b, d’A62, Gri66, Pai65, Pla61, HSP66]. commissioning [Ano60a]. Committee [Smi63b]. Communication [Smi66b, Bri67, PR64]. communications [Pea63, Wil66b]. Compact [Sim69, Jen66]. compaction [HW67]. Comparison [Jar68, Box65, Box66a, DGM68, Goo61, Win69]. Competition [Ano61a, WR67]. compilation [Gea64a, Rya66]. Compile [Hex67]. Compile-Time [Hex67]. Compiler [BMR67a, BMR67b, Ryd64, BC67, Cor65, CGH^{+}68, Fin68, For65, Jon65a, MR67, Sch65, Tay68b]. compilers [CP62]. Compiling [Hus61, Wis68]. COMPL [FS66]. Complex [Boo68a, Boo68b, Rud67a, Rud67b, Wri63, Che65b, CL66a, Gui68, Lar64, Mak68a, Sea69]. complexity [Und63]. Composition [Gil63b]. Comprehensive [Sol60]. compressed [Lan65]. Compression [Fra67]. Computability [Coo69]. Computation [Ano60d, BB69, Duk66a, Han62, Her61b, Jon69, Rac65, BW69b, DP66, Duk66b, Jon66, Lew65a, NPW68, Sum65, Ful66a, Ful66b, Cra66a]. Computational [BT63, Pap68b, Per65, Tew68, FM66, Pap68a, Ral61].

Computations [Coo66]. compute [Rya66]. Computed [Wil61c, Loi68]. Computer [Ano61e, Ano61h, Ano62c, Ano63b, Ano67e, Ano69f, Ano69k, Ano69n, AS63, Bag62, Bar67, BD61, Buc60, Buc65, Bur66b, Che62b, Cla66, Cle60, Coo60a, Cud67, Dou65, Ede63, FFMR66, Fox69, Geo66, Gil63b, Gol69, HG67, Hug61, Ili64, Kin69b, LW65, LW66a, LR67, LEN63, Lew65b, McO66, MP65b, Mil69, Mut67, OB68a, Ong69, Pag62, Pal63, Pay66, Phi66, Smi64, TS64, Tho61, Tho62, War60, WY62, Wil67b, WH68, ABN61, Bae66a, Bae66c, Bar65c, Bea68, BS61, Bri67, Car64, Cha62, Che62a, Che64, Col64a, Cor65, DN64, Dol66, Fin60, Fir61, For66, FGR67, Fos64, FL68, Gla69, Goo61, Gos63, GB61, Gro68, HK67, Hil67, Jan63, JN65, Joh61, Kni68, Lan65, MFM64, MW68, ML65, Pal62, Pen66, Pla61, PW67, Red60, RWT64]. computer [Rig64, RW61, SW68, Smi66c, Smi68, Sut60, Tay61a, TB63, Wil61b, Wil67a, WN68, You63a, dP60, Bae66b, Bur66a]. Computer-Aided [Kin69b, Gol69]. Computer-Oriented [Bag62, ML65]. Computers [Ano62d, Ano67d, BD61, BHRS69, Cha63, Dou61, Geo68, GT63b, HRS68, KV62, Mas63, Nix64, Pap60, Pla64, Pri66, RP63, Tho68c, Weg62, Wex69, Yat62, Cle67b, Col63a, Cra67, Edm60, Gea61b, Gow62, Gri66, Mel66, OS65b, Pin61, RW68, cGR61, Tho68a, dP60, Clu63a, Wri66c, Car63]. Computing [Ano65f, Bis66, Fox61b, Mor69, Nud69, PCSG63, Sla66, Tho60, Bag61b, Cla60, EAL66, Gen69, Gil65, Pit65, PL68, Pri64, WM69, Woo68]. concluded [Ano60b]. Conditional [May69]. Conditions [CP66a, Has61, JM68, KM66, Kin69a, Osb69]. Conduction [MP62, Bir68, Osb69]. Conference [Ano62c]. Conferences [Ano62m]. configuration [AM65]. Congress [Ano61g]. congruential [DR67, MP68]. Conjugate [FR64]. Connolly [Her61b]. conservative [RM69]. consideration [Hil67]. Considerations [Pag60, cGR61, Bon69, Gil66]. consisting [Wil68b]. console [PL68]. Constant [SS63a, Jar67b]. constants [Tee67a]. constrained [Box65, Box66a, Gui68]. constraints [Ord64]. constructing [Alm66]. construction [And66a, Bar65c, CO66, FL68, How67a, JS68, OS65b]. containing [Scr64c]. Contributions [Pol60, Dru63, Str60]. Control [BL62, Cle66a, Cud67, GT63b, Ham66d, Hou63b, Lei62, Nig68, Wes69, Bri67, For66, FGR67, Fra68, GE69, Hug61, JP69, Lam63, Pal62, TS64, Wes64, SS63b, May69]. controllability [GE69]. convention [Ros63b]. Convergence [Bar61, Sto68, Wat68b, Wil65c, Aha65, CP66b, Erd65, Eva65a, GL70, Lyn69]. Convergent [FP63]. conversational [Col66b]. Conversion [Tiz60, Flo64]. Convex [BH69b, Col66a]. Cooper [RS69]. Corps [TS64, Sla60]. Correction [Ano64a]. Corrections [Sac62]. Corrector [Hal67, For61, Ral61]. Correlated [Wal64]. Correspondence [AT70, And69a, And69b, Ano60c, Ano61c, Ano62e, Ano62f, Ano64b, Ano66b, AMKM66, Bae66c, Bar64, Bar66a, Bau65a, Bau65b, Bem69, Ber69, Ber61, Bic69, Bou65, BSPI65, Box66b, BM61a, BH69a, BH69b, BBdD66, Bux63, Car70, CL66a, Col67, CG65b, CG65a, Cra67, DD70, Deu69, Duk66b, Eva66, Fai61, Fel63, Fel64, Fel69, Fel70, FY63, Fra70, GL70, Goo68a, GB69, HL70, HB61, Hig65, HLOS65, Hig67, Hil67, HSP66, Ker63a, Ker63b, KK68, Kin71, Kiz66a, LWS69b, LWS69c, LO65, MF61, Mer61, Moo65, MR65, Myl66, O’B68b, Obr67, Owe68, OL68, Pak67a, Pak67b, Pap68a, Pay63, PI65, PB69, Pri60, Red71, Ros66a, Rut63, Sam63b, Sam63c, Sam66a, Scr64a, Sel72, Sol68a, Sol68b, Str65a, Str65b, Tay64, Tay68a, Tay68b, Tee67a, Tee67c, Tem69a, Tem69b, Tho64b, Tho68a, Tho68b, Tho66, WD63]. Correspondence [Wat61, Wel70, WP69, Woo61, Wri66a, Ano67g, Bar69a, Chu68, Gui68, Kin67b, O’C69, Sam68c, Wil67d]. Corrigendum [Ano65c, Ano67f]. COSEC [PL68]. COSMOS [O’B67]. cost [Bon69]. Cotes [Sac62]. Counter [Col64b]. Courtauld [Cra66b, O’B67]. covariance [Goo61]. Cowper [RWT64]. CPL [BBH^{+}63]. CPL1 [CGH^{+}68]. Crank [Dav67b, KM66, Ros66a]. Creed [GS63, Sam63c]. CRESTS [Cra66b]. criteria [GL70, Lyn69, Pow67]. critical [DGM68]. Crout [Tew69a]. CRT [Gol69]. crystallographer [Mat67]. CSL [Bux66]. Cubature [Mei66]. Cubic [AH69, Bic68, Bic69, Fyf69]. cumulative [Smi66a]. Current [JJ67, d’A62, Box66a]. Curtis [OS68a]. Curve [CW61, Cle60, Sne66, Ada69, Hay69, Le 69, OSS69]. Curves [Ano66e, Kiz66b, BH69b, Kiz66a, Wri66a]. Cybernetics [New65b, New65a]. cycles [Hun60].

D [PCSG63, Wil67c]. dama [Fin60]. damped [Til69]. Data [Ano63c, Ano65f, Ano65a, Dav63, Fra67, Har65a, Har65b, Hin60, Kin69b, LW67c, Llo64, LT61, Lon61, Rus63, Sla60, Smi63b, Tho60, Wil61b, Wil61d, WKRJ60, Wri61a, Ano63a, BFH61, Dac63, DM60, FE69, Fri65, Gat65, Gil64, Pit65, Sol68b, Spä68, Str61, Til69, Wil63b, Wis68, dP60, Ano61d, SS63b, Gea64b]. Davies [PCSG63]. day [Ano67g]. dealing [Bro60]. Debugging [Bar69b, JP69, Bar69a]. Decision [Gri68, Kin67a, Car70, Kin69a, Kin71, LW66b]. decoding [Wil64c]. Decomposition [BHS65, BH68]. Deferred [GlM68, May64]. defined [Pow67]. definite [MT61, MP65a, MP66, OS69]. Definition [Kin69b, Hou69a, Thr69]. degree [Sco69, Spä69a]. delay [GE69]. Delays [KV63, Tho68c, RW68, Tho68a]. Dependent [Wol66]. depending [Tom67]. derivative [All66b, KM66, Mil60, Scr64c]. Derivatives [Fle65, Pow64, Pow65a, LM63a, Zan67]. derived [NE67]. Descent [FP63]. Description [Ree67, HPS61, ML65]. Design [BW69a, BD61, Dem60, Gri62, New66, PCSG63, BH68, Bur66b, Cam69, Lav64, OS65b, Pin61, Vin64, Wil67a, WN68, Wis64, Her61b]. designational [Wat63]. designing [Gri66]. Desirable [Pri64]. determination [Car61]. determining [Boo67f, Bye67, Hil69, Rya66, Woo70]. detonation [Pal63]. DEUCE [DH60, Red60]. development [Hur65a, Hur65b, Lum64, PL68, Smi64, Wil67d]. Developments [d’A62]. diagnosis [BH63, Ler69]. Diagonal [Wil62c, Bak64, Cay61, EF63b]. Diagonalisation [Sea69]. diagonalization [RW61]. dialects [HB67]. Dictionary [GT63b, Kin69b]. Difference [MP62, Tee63c, Eva63, Eva65b, Eva66, Osb63, Osb64a, Osb64b, Osb65, Ran68a]. differentiable [BH69b]. Differential [Joh66, Kha66, SS63a, All66b, Cay61, CN63, CG65a, Dav67b, Day65, Day66, Day67, Eg69, Eng69, Eva65b, Eva66, For61, Fox62, Fox63, Gib60, Hai69, Has61, JS67, Kar66, Mak68b, May64, Mil60, Nor64, Oli69, Osb63, Osb64a, Osb64b, OM64, Osb65, OW68, PC64, RM69, Ros63a, Scr64c, Scr65, Tho68b, Wal64, WP68, Wri64]. Differentiation [FR63, Hun61, Sch65]. diffraction [MP65a, MP66]. Digital [AS63, Cle60, Ful66a, Ful66b, HB66, PCSG63, RP63, Tiz60, Bar65c, BBS66, Cla60, DP66, Edm60, Fin60, GB61, JN65, Joh61, Pit67, Tay61a, Wal64, PCSG63, Wil66b]. digits [Ber69, Sal68]. dimensional [Che65b, Phi67c]. Dimensions [CD69, Phi67b, Hal68, MLB63]. dioded [Vic65]. Direct [Bra63b, Jen61, Ker63b, MP65b, Osb65, All66a, Pak67b, Pal69]. directed [BC67, Ree67, Sch65]. direction [Eva65a, FM64, Had68, Rya66]. direction-independent [Rya66]. Dirichlet [FM66, JM68]. discretization [PC64, Wat68b]. Discussion [Ano62g, Ano62i, Ano62h, Ano62j, Ano63a, Ano67g, Bar69a, Chu68, Gui68, Kin67b, O’C69, Sam68c, Wil67d]. disk [Ros66b]. disk-based [Ros66b]. display [Gla69, Sip69, Wis68]. Distance [Cha63]. distinguish [Ros63b]. distributed [NPW68]. Distribution [BB69, But61, Ash68]. Distributions [YA61]. divisors [MST64]. Do [Nef65, Bae64c]. Do-it-yourself [Nef65]. Document [Hin61, Str61]. Documentation [Dou61, Kin69b, Vin64]. double [GRB68]. Douglas [FM64]. Douglas-Rachford [FM64]. down [Dow69]. drawing [Pit67]. driven [FE69]. dual [Tur68]. Dumping [KV62]. Durand [Wil62b, Wil63a]. during [Rya66, Smi66c]. Dynamic [IJ62, Pri63, Lan65, MW68, Sla64a]. dynamics [NPW68].

E.D.P. [Hic60]. E.S.P. [Wil64b]. each [Boo67f, Hil69, Woo70]. Early [Cor62, Hic60]. Eastern [Hag69]. Economics [KV62]. Economy [Sla66]. edited [Cle66b, PCSG63, RP63]. Editor [Ano67k, Ano69k, Tho68c, Ano69l, Ano69m, Bar65d, Hur65a, O’B68c, Pap68b, Par67, RW62, Sam68a, Str65e]. Editorial [Ano62k, Ano65b, Ano65c, Ano66c, Ano67h, Ano68e, Ano69g, Ano69h]. Editors [Ano69f]. Editorship [Mut67]. EDSAC [BH63]. Effect [KV63, GL70, Lyn69, Mal61]. Effective [Coo69]. Efficiency [Hil67]. Efficient [Pow64, Sam65, RW61]. Egdon [Bar66a, Sam66a, BHJV66]. eigenfunctions [Bye67]. Eigenproblem [Wil60]. Eigensystems [Wil61c]. Eigenvalue [Wil66a, Boh68, Erd65, GM65, Osb64c, OM64]. Eigenvalues [PW69, Wri63, Bye67, FJ66, GRB68, Par67, Wil65d]. Eigenvectors [Ano64a, Tee63b]. Election [MM60]. Electric [Dun62a, Dun62b, KP60, RW67]. Electrical [Ras66]. electricity [Bur66b]. Electronic [Ano61e, Gat65, KV62, Edm60, Fri65, Geo63, Goo61, TB63]. element [Boo67f, Hil69, Thr69, Woo70]. Elementary [Dre66, MST64, BT63]. Elements [Ili69, Boo67k, Boo67g, Boo67i, Boo67j, Boo67h, Boo67f, Con66, HW67, Hil69, Ros64a, Woo70]. Elimination [Wil62c]. eliminations [SF61]. Elliott [Coo60b, Hoa62b, Hoa63b, Wil64b]. ellipses [Pit67]. Elliptic [Osb66, Eva63]. Elsevier [GT63b]. empirical [Loi68]. encountered [MP65a, MP66, Smi66c]. end [Boo67e, Boo67c]. energy [Jon66, Rac65]. Engineering [Her61b, Ras66, Ler69]. English [Dun62a, Ker63b, Bra63b, Dun62b, HSP66, Pai65]. enquiry [Smi68]. entry [Car70, Kin69a, Kin71]. enumeration [BP69]. Environment [Bar69b, Bar69a]. EOL [Luk67]. Equation [MP62, Wat67, Bir68, Bye65, Day65, Eva65b, Eva66, FM64, Fet65, Had68, JS67, JM68, Mal61, Mil60, Osb69, Ran68b, Tho68b]. Equations [Ano60g, Boy60, Bro69d, Cle65, Fle68, Jar66b, Joh66, O’B68c, Rei67, SS63a, Wri63, All66b, BFMW64, Bak64, Bar65a, BSD60, Bro69c, Car66, Cay61, CN63, CG65a, DM64, Dav67b, Day66, Day67, DS67, Eg69, Ell63, Eng69, Eva63, EF63b, Eva64, Fie70, For61, Fox62, Fox63, GJM68, Gib60, Hai69, Has61, Hur68, JS67, JN65, Jar66a, Jar67b, Jen66, Kar66, LW66b, Lin69, LM68, Lyn64, Mak68b, MT61, May64, MP67, Nor64, O’B68b, Oli69, Osb63, Osb64a, Osb64b, OM64, Osb65, OW68, PC64, RM69, Ros63a, Scr64c, Scr65, Tew69b, Wat68b, WP68, Wol69, Wri64, Wil62b, Wil63a, Kha66]. equipment [GS63, Hop63, Sam63c]. Equivalence [Coo66, RS69, Lyn64]. Errata [Ano62l, Ano64c, Ano64d, Ano65d, Ano66d, Ano67i, Ano68f, Ano68g, Dun62a]. Erratum [Ano67j, MP66, Sav70]. Error [Ano66e, Eng69, Kiz66a, Kiz66b, OS68a, Wil61c, Wri66a, dW64, BH63, Cha68b, Cur65, Gen69, Hai69, Jar67b, Loi68, OSS69, Oli69, Osb64a, Phi68b, Scr64b, Smi66a, Ful66a, Ful66b]. errors [PC64, Pow67, Wal64]. essay [Las68]. established [Wil67d]. establishment [Tay61a]. Estimate [Phi68b, Hai69, Loi68, Smi66a]. estimates [Cha68b, Eng69, dW64]. Estimating [Gos63]. Estimation [Cur65, Eva65a, Scr64b, Cha67, OS68a, Oli69, Ran68b]. Europe [Hag69]. Evaluating [Fle65, HG67, Tew68, GE69, Gro68]. Evaluation [CR63, Fox69, Lan64, Mak68a, MP65a, MP66, Coo67, CP68, Del68, Mil60, MB66, OS69, Win69]. Event [Nud69]. EVERYBODY [Hig63]. evolution [Pin61]. examination [Col64a, FL68, Woo68]. Examinations [Ong69, Gro68]. exchange [CP66b]. executive [Ros66b]. Exercises [TS68]. Exhibition [Ano61e, Ano61h]. expansion [Cha67, Coo67, You63a]. Expansions [Sto68]. expenditure [Red60]. Experience [BMR67b, Cor62, Dac63, Dij62, Gle62, Lum64, Red60, Tay61a, Wil63b, Ayr64, FGR67, Gol64, Kin62, Mil63, Nef65, Pla61, Ric64, Sut60]. Experiences [BHS65, Hic60, Sla64b]. Experimental [New66, Gea61b]. Experiments [Mic63, Abb64, YGS63]. explicit [Bir68, Eva65b, Eva66]. explicitly [Scr64c]. explosive [Pal63]. exponential [Spä68]. exponentially [Til69]. Expressing [Tho62]. Expressions [Hus61, Lan64, Col63a, Sip69, Wat63]. Extended [Cle66a, Gea65, BM63, And66a]. extension [Hoc62, Hur68, OS65b]. Extensions [For61, Yul68]. External [Tho60, dP60]. extract [Cra66b]. extrapolated [Eva63, Eva64]. Extrapolation [Wri63, dW64]. Extremum [Bae62].

Facilities [LR67, Wil61d, BMR67a, Nap67, PL68, Pri64, You63a, For66]. Fact [Hur65b, Cli62b, Hur65a, Lum64]. Factor [Hio68b, Car61, Eva62, Eva65a, LM68]. factored [Woo69b]. factorial [RS69]. faculty [Alm69, DD70]. Family [Mei66, Red60]. Fast [Boo67k, Wis64]. fault [OS65b]. features [BBH^{+}63, BRC66]. feedback [GE69]. fer [Fos64]. Fibonacci [Boo79, Ove67c, PHJ67, PP79]. field [Aha65, Gea64a, Wil61b, Jon69]. fields [Ran68a]. file [Wis68]. Filing [Fra69b, Wil64a]. Filtering [Woo66c]. filters [Bea68]. Finding [Nic66, Pow62, Pow64, Che65b, FJ66, JJ67, Lar64, Par67, Rei66, Ros60]. Finite [Tee63c, Eva65b, Eva66, Osb63, Osb64a, Osb64b, Osb65, Ran68a, Thr69]. Finite-Difference [Tee63c, Eva65b, Eva66, Osb63, Osb64a, Osb64b, Osb65, Ran68a]. Firm [BHRS69, HRS68]. First [War60, All66b, BFMW64, Che62a, Lin69, Mil60, RM69, Rya63, Scr64c, Sut60]. first-order [RM69]. Fitting [Cad61, CW61, Cle60, Spä68, Til69, Hay69, OSS69]. five [Chu68]. Fixed [Nie65]. floating [SS63b]. floating-point [SS63b]. Flow [Gre69, OB68a, How67a]. fluid [Abb64]. flying [GB61]. flying-spot [GB61]. Force [Dav63]. Forecasting [MM60, Hio67]. Forests [Win60]. Form [Sim69, Wil62c, Gla69, SF61, Woo69a, Roh68]. Formal [Ler69, Nap67]. format [Car70, Kin69a, Kin71]. formats [Zar69]. Formatting [Sip69]. forming [Col63b]. formula [Cha68b, JS67, KM66, OS68a, Ran68a]. Formulae [FR63, Mei66, MP62, Ste67, Bar61, LM63a, dW64]. Formulas [Sac62, Tha66]. Forthcoming [Ano61f]. FORTRAN [HS69, MP65b, TH64, Tay68b, Gle62, HL70, Hea68, Pul64, Wel70]. forum [Ano63a]. forward [Rya66]. Fourier [Bel63, Boo68a, Boo68b, Dol66, Gen69, Mat63, Sto68]. Fourth [Ede63, Eva65b, Eva66, Kar66]. Fourth-Generation [Ede63]. fourth-order [Eva65b, Eva66]. Fox [Cle68]. Fragments [Lev65]. Francis [Wat11]. Fredholm [BFMW64, Ell63]. frequently [MP65a, MP66]. friend [Mat67]. full [You60]. full-scale [You60]. Function [FR64, Fle65, Fox69, Jar66b, NM65, Pow62, Pow64, Boy63, Cha67, Hor60, Mil60, MR65, RS69, Ros60, Til69, Tom67, Zan67]. functional [Wat68b]. Functions [Coo69, Jar68, Pow65a, BY66, Bea68, Che65b, CO66, Goo61, GL64, JN65, Lar64, Ord64, Cle66b]. Fundamental [Tho62]. Further [Fra70]. Future [Ano66i, Ano69i, Ano69j, Ham66c, Wil61d, War60].

G [Ano64a, MM69, RP63, Wil67c]. game [Fin60, Fos64, Mic63]. game-learning [Mic63]. game-theoretic [Fos64]. Games [Hal66]. Gamma [Pal62]. Garabedian [Cle66b]. gas [Lan65]. Gauss [Cha68b]. Gaussian [Bar61, FR63, Tom67]. Gear [Lei62]. General [Rus63, YS60, And66a, Dol66, GE69, LW67a, LW67b, MP67, Ran68a, Ros63a, Str65c, Und63]. general-purpose [Und63]. generalised [Had68, RS69]. Generalized [Fle68, Tew68, FM64]. Generation [Ano61h, Ede63, FR63, OB68a, RW68, Tho68a, Tho68c, Edm60]. generator [BW69a, Hun60]. generators [DR67, MP68]. Genius [OS65a, Phi65]. Geoffrion [BP69]. geometry [MP65a, MP66]. Giro [Gra67]. Givens [RW61, RW62]. Gladwell [Wil67c]. Go [TW64]. God [Jam66]. Gödel [Goo68a, Goo68b]. Goertzel [Gen69]. GOLD [Ove67a]. GOLDSEC [Ove67b]. Golem [Jam66]. good [Mus67]. Goodman [RP63]. Goursat [Day66]. government [BS61]. Grader [Ber66]. Gradients [FR64]. Grading [TS68]. gradual [Pai65]. Graduated [Pol60, Dru63]. grammar [CND66]. graph [WP67, Woo69c]. graphic [Gla69]. graphical [Lar64, New68]. graphs [NE67]. Great [TB63]. greatest [Ros60]. grids [Boy63]. grist [Mus67]. Groups [Rei69]. growth [Und63].

H [Cle66b, Cor62]. Handbook [Ano60d, Ano65a]. Handling [Hin61, Mil68b, Gow62, Str61]. Hao [PCSG63]. hardware [GS63, OL68, Sam63c]. Harvard [RP63]. having [Bea68, Cha67]. Heat [MP62, Bir68, Osb69]. heights [MW68]. held [Ano67g]. Hessenberg [Han62]. heuristic [Bur66b, Yul68]. hierarchic [JS68]. hierarchical [LW66a, LW67a]. High [MP62, Dav67b, Mak68b, Pal63]. high-explosive [Pal63]. high-order [Dav67b]. higher [LM63a]. Honorary [Ano69f, Ano69k]. Householder [Ano69e, Rei67, RW62, Wil60]. Human [Bae66b, Geo68, dF66]. Humanistic [Wex69]. Hybrid [NPW68, Car64]. Hydro [KP60]. Hydro-electric [KP60]. hyperbolae [Pit67]. hyperbolic [Day66, GlM68]. Hypercubes [LM63b].

I.C.T. [Ham66b]. I.S.O. [Ros64b]. ICC [GT65, Dou66]. identification [Tho64a]. Ifip [Ano61g, GT65, Dou66]. IFIP-ICC [Dou66]. II [Wil63a, Ful66b, HPS61, LW67b]. III [Lew63, Pin61]. ILLIAC [Gea64a]. imbedding [Wol66]. Impact [Dou61, You63a]. Implementation [Dun62a, Dun62b, FE69, Bro69b, FK68, Ham66b, Hou69a, FGR67]. Implicit [Hai69, BP69, Had68, Mak68b, Ros63a]. Impossible [ApS65, AMKM66, BSPI65, HLOS65, Str65d, Str65e]. improved [FM66, Pal69]. Improving [Bag61a, Fos68b]. inadequate [GL70, Lyn69]. incompressible [Abb64]. independent [Rya66]. Index [Ano65e, Ano66f, Wri60, Wri61b, BC67]. Indices [Wat69]. Induction [Bur69b]. Industrial [Cud67, Hoc62]. industry [Bag61b]. inference [Ull67]. Information [Ano60e, Bur66a, Cli62c, CH60, Dou66, H. 68, Nee63b, WB68, Woo66b, Woo66c, Bon69, Col66a, Gat65, LW68]. information-statistic [LW68]. Initial [Bae64c, Car64]. initial-value [Car64]. Initialising [BP69]. initiation [Pal63]. inner [Wil68a]. Input [Dun63, Coo68a, Gla69, Hoa63b, New68, Sol68b, TS64, Zar69]. input/output [Hoa63b, Zar69]. Inquiries [Wri61b]. Instability [Wil62c, KM66]. Installation [Rus63]. Instruction [Geo66, OW66]. insurance [Dru63, MFM64, Ano60a, Ano60f]. Integer [Dak65, Law69]. integral [BFMW64, Eg69, Ell63, JS67, Lin69, Wol69]. Integrals [CR63, FR63, CP68, Del68, GL64, MP65a, MP66, OS69]. integrands [Fox67]. Integrated [Hop63, Dav63]. integrating [Wal64]. Integration [CG65a, LM63b, Phi67b, Tha66, Wri66b, Day65, Fox67, Gib60, Hai69, Phi67c, Ral61, RM69, Tom67, Cle65]. integrations [Ede65]. Integrity [Fra69b]. integro [Day67, Eg69, WP68]. integro-differential [Day67, WP68]. integrodifferential [Wol69]. Intelligence [Flo68c]. Interaction [Cor65]. Interchanges [Hea63b]. interconnections [Wis64]. Internal [Tho60, KHPS61]. International [Ano62m, Gat65, Cla66]. interpolating [Win69]. Interpolation [Lar67, Sal70, Spä69b, Bic68, Bic69, Hal68, Hun68, Pow67, Sco69, Spä69a]. Interpretation [Car70, Kin71, Kin69a]. interpreter [Fos65a, FE69]. Interpretive [KB62, Hor60, SS63b]. Interruption [Wet66]. interval [OS69]. Introduction [New65a, New65b, Pol60, Sla60, Tho61, Woo60, For65]. Introductory [Osb66]. Invariant [Wol66]. inventory [For66, FGR67, Lew65a]. Inverse [Fle68, Pow69b, Boh68, May66]. Inverses [Tew68]. inversion [NE67]. investigation [All66a, Pak67b]. Invoicing [Jen61, Pal62]. involving [LW66b, MP65a, MP66]. Irascible [Phi65, OS65a]. irregular [LM68]. Irving [PCSG63]. Irwin [PCSG63]. ISO [Bar65d]. Iteration [Jar66b, Jar68, Tee63b, Wri63, Ano64a, Car66, DM64, Eva63, Eva64, Pri68a, Wol69]. Iterative [EF63a, Eve63, Hun61, Jar67a, Kha63, Lew65a, MT61, Osb63, Pow62, Tha62, BW69b, Erd65, JN65, Jar66a, Lar64, Nor64, Tew69b].

J [Ano64a]. Jacobi [Cha62, Sea69, Wri63]. jet [Geo63]. job [NP68]. job-shop [NP68]. Jobs [Pag62]. Journal [Cla66, Ano69k, Ano69f, Ano67e, Kha66, Mut67]. journals [Chu68]. Junction [Sav69, Sav70].

KALDAS [DP66]. KDF [Dun63]. KDF9 [Bar66a, Dun62a, Sam66a, Bar65b, BHJV66, Dun62b, PL68, WR67, Hoa64]. KDF9ALGOL [Mul69]. Kent [Nee63b]. key [Col63b]. kind [BFMW64, Lin69]. know [Hig63]. Kutta [Cur65, Day66, Eng69, Kar66, Scr64b].

L [Cle68, Wil67c]. Laboratory [Ano67g]. Lamé [Wil65b]. Lanczo [Kiz66a, Kiz66b]. Lanczos [Wri66a, Ano66e, Wra66]. Language [Bra63b, BM60, BL62, Cle66a, Cor62, Ili69, JJ67, BBG^{+}63, Bag61a, BFH61, Bur67, Cam69, CW62, Cor65, CP62, DM60, Elc65, FS66, Gri66, Ker63b, KM65, LWS68, Luk67, Nap67, Smi68, Wil68a, Bag62]. language-compiler-computer [Cor65]. languages [HSP66, Pai65, Woo69b]. Lanuage [Dun67]. Laplace [Had68, JM68, LM68, Ran68b]. Large [Cha63, Dou65, Kha63, Lew65b, Liv60, Sla60, Wri60, Che64, Rig64, Smi66c, Sut60, Woo69c, PCSG63].

Large-Capacity [PCSG63]. Latent [Han62, DP64]. lawyers [HK67]. leading [Chu68]. LEAPS [Rya63]. Learning [Bur68, Mic63, And66b]. Least [Cad61, Fle68, Kha63, Rei67, JM68, Pow67, Ros60, Tew69b]. Least-Square [Kha63]. Lebow [PCSG63]. left [Woo69b]. Legal [McO66]. Legendre [Cha68b, Ede65, GL64]. Length [Ham69d, HW67]. LEO [Ano60a, Gos60, Pin61, Ano60f, Lew63]. let [Mus66a]. Letter [Ano69l, Bar65d, Hur65a, O’B68c, Pap68b, Par67, RW62, Str65e, Ros63b]. Letters [Ano67k, Ano69m, Tho68c]. Level [Bro60, LW67c, BH68, DN64, RW61]. levels [Duk66b, Duk66a, Jon66, Rac65]. lexicographic [Boo67i]. Library [Ash67b]. Liebmann [MST64]. Life [Gil64, Sut60]. limit [GlM68, May64]. limited [Car70, Kin69a, Kin71]. Line [Bar69b, Bar69a, Deu69, Eva62, Eva65a, Fra70, MW68, Phi68a, PL68, RW67, Spä68, WM69]. line-thinning [Fra70]. Linear [Ano60g, Boy60, Fie64, Fle68, Mor69, OW67, Pow65a, Rei67, SS63a, Tew69c, Aha65, Bak64, Bar65a, BSD60, Bur67, BP69, DS67, EF63b, Eva64, GlM68, Has61, JS67, Jen66, Law69, Lyn64, MT61, MB66, Nor64, Ord64, OW68, Scr65, Tho68b, Til69]. Linear-Programming [Mor69]. Linguistics [Hag69]. link [Pit65]. Liouville [GM65]. Lisp [d’I69]. list [Bar65b, Boo67e, Col63b, CW62, Fos65a, Hou69b, Tru66, Wis64]. list-processing [Bar65b, CW62, Fos65a, Hou69b, Wis64]. listed [Woo69d]. Lists [Wil65a, WJ61, Elc65]. literature [DGM68]. LITHP [Ham66b, Tru66]. Liver [WY62]. loading [Cle67b]. Lobatto [JS67]. Local [CD69]. Locating [Bae62, Jar67a]. logarithmic [Cha67]. Logic [Ede63, LEN63, AA68, Wis64, PCSG63]. logical [Wil68b]. London [Wil67c, Ano61h, CGH^{+}68, Dol66]. Lookup [OW66]. lower [Smi66a]. Lsix [Hou69a]. Ltd [Cle65].

M [Wil67c]. Machine [Flo68c, EAL66, Flo64, Ili69]. Machines [Coo68b, Fox61b, Goo65, PCSG63, And66b]. macro [Bro69b, Nap67]. macrogenerator [Str65c]. magnetic [BSD60, Cha62, DGM68, EAL66, Gil66, Goo64]. magnetic-tape [BSD60, Cha62, Goo64]. magnitudes [Boo67g, Hil69, Woo70]. main [BBH^{+}63, BRC66, Gla69]. Maintenance [Pag60, Ros66b]. man [Bri67, Mus66b, Mus66a, Mus67]. man-computer [Bri67]. Management [Gil63a, H. 68, Har65b, Lew65b, Har65a]. Manchester [HPS61, KHPS61]. manipulation [Luk67, Obr68, Ros67]. Manipulative [Bar67]. Manipulator [d’I69]. Manufacturing [Rus63]. many [Ber69, Sal68]. Mapping [Sla64b]. Market [Coo60a, Gos60]. Marking [Gro68]. Markov [May69]. Marshall [PCSG63]. Mass [Fra69b]. Masson [Wil62b, Wil63a]. Matching [Hex67, Wri61b]. Mathematical [Bel69c, BT63, Fos65b, Fox69, Mil64, PCSG63, Vaj64, FK68, Tee67a, Hag69, Hou63b]. Mathematics [BT63, Goo65, Cla66, Pay66]. Matrices [Kha63, Cay61, FJ66, GRB68, NE67, OK63, Par67, Sea69, Tew67, Tew69a, Wil65d]. Matrix [Boo67b, Han62, Pow69b, Ras66, Wil62c, Wil67c, Wri63, AM65, DP64, Fet65, MT61, May66, O’B67, RW61, SF61]. matters [Ral61]. maximum [Phi68b, Pow67]. May [Ano67g]. Measure [WB68]. Measures [Tiz60]. Measuring [Jan63, GT63b]. meat [Mus66b, Mus66a, Mus67]. Mechanical [Lan64]. Mechanization [FR63, Mic63]. Mechanized [Nee63b]. Mechanizing [Wri60]. medicine [Ler69, Pay66]. Members [Ano63b, Ano69n]. membranes [O’C69]. Memory [Ano67l, PCSG63, WR67]. Men [Ano61b]. Mercury [And66a, Gea65, Gib61, PM62]. mesh [Eva62]. Method [Ano66e, Bro69d, Cle65, CR63, FP63, Han62, Hun61, Jar67a, Kha63, Kiz66b, NM65, O’B68c, Pow62, Pow64, Pow65a, Sne66, Tew68, Tha62, Wat67, Wil60, Wil62c, Wri63, All66b, Bak64, Bar65a, Box65, Boy63, Bur66b, Bye67, Cha62, Col63b, Day65, Day66, Ell63, Eva63, Eva64, Eva65b, Eva66, FM64, FM66, For61, GJM68, Gen69, Gib60, Gou68, GRB68, Gui68, Had68, Hay69, Hun68, Jar67b, JM68, Jon66, Kar66, Kiz66a, MP67, ML65, O’B68b, Osb64b, Osb64c, Pri68a, Rei66, Riv68, RW61, Ros60, Sea69, Smi66a, Tew69b, Wri66a]. Methods [BW69b, Bis66, CW61, Fle68, Hal67, Hou63b, HB66, Nud69, Ped66, Pri66, Tew69c, Tho61, Wri66b, All66a, Box65, Box66a, Fox62, Goo61, Jar66a, Ler69, Lin69, Mak68b, MT61, Mil68a, OS65b, OK63, Osb64a, Osb65, OW68, Pak67b, Pal69, Pea69, Ral61, Wat68b, Wis64, WP68, Wra66, Wri64, Jon69]. metric [Pea69]. Metzger [BT63]. Michaelson [Wil67c]. Midsummer [Ano62m]. mill [Mus67]. Milne [Smi66a]. Mind [Geo68, Wal69]. Mini [Gil69a]. Mini-COBOL [Gil69a]. minimal [Col66a]. Minimax [CP66a, CP66b, CO66, OW69, Tem69b]. minimisation [Hur68, Pea69]. Minimization [Dav68, FP63, FR64, Fle65, NM65, Pow69a]. Minimizing [Pow65a, Zan67]. Minimum [Pow64]. Ministry [Ano60a, Ano60f]. MINX [PHJ67]. Mixed [Dak65, LWS68]. MLS [LWS68]. mode [Col66b]. model [Law69, Mic63]. Modelling [Wal69]. Modern [Hio68b, HS69, Ham66d]. modification [Ash68, Gui68]. Modifications [Pow69b]. modified [Eva63, Eva64, Sea69]. Molecular [CR63, FR63]. molecules [Duk66b, Jon66, Rac65]. Moment [Jon69]. Money [Ano61b]. Monitoring [Wet66]. monothetic [LW65]. Monte [CR63, KB62]. Montecode [KB62]. Morning [Ano62g, Ano62i]. most [Chu68]. motion [Abb64]. moving [LW66b]. Multi [Ham69d, LW67c, BH68, Fra68, Lam63, You60]. multi-access [Fra68]. Multi-Length [Ham69d]. Multi-Level [LW67c, BH68]. multi-project [Lam63]. multi-stage [You60]. Multidimensional [LM63b, CP68]. Multiple [CL66a, CL66b, Col64b, How67b, Hea63a, SS63b, Wil67a, WN68, Wil65d]. multiple-access [Wil67a, WN68]. multiple-precision [SS63b]. multiplication [Joh61]. Multiplicative [DR67]. Multipoint [Jar66a]. multiprogramming [HLN68]. multiway [Gow62]. Music [Gil63b, Woo66d].

N [Hal68]. Names [Bra63b, Elc65, Ker63b, Tho64a]. National [Ano60f, Dru63, Ano60a, Ano67g, Coo60b, Gra67]. National-Elliott [Coo60b]. natural [Flo64]. nature [Bri67]. Naur [Roh68]. Navier [Bye65].

NEBULA [BFH61, Rou62]. Necessary [CP66a]. needs [Cor65]. Netherlands [Ano61d]. nets [Wil68b, Cle65]. Network [Sol60, Boo67e, Boo67c, Boo67d, Car66, Nic66]. Networks [Hal66, Nud69, Phi67a, Bur66b, Mar69]. Neville [Hun68]. Newsletter [Tra66]. Newton [Sac62]. NF [Wil62b, Wil63a]. Nicolson [Dav67b, KM66, Ros66a]. node [Boo67e, Boo67c, Boo67d]. nodes [Boo67e, Boo67d]. Noise [Smi66b, Sum65]. Non [Fle68, Pow65a, Aha65, Bar65a, DS67, GlM68, Gri66, Has61, HB67, JS68, Kni68, Mic64, Nor64, Smi68, WP68, Wol69]. non-hierarchic [JS68]. Non-Linear [Fle68, Pow65a, Aha65, Bar65a, DS67, GlM68, Has61, Nor64]. non-paged [Kni68]. non-programmer [Smi68]. non-programming [Gri66]. non-scientist [Mic64]. non-singular [WP68, Wol69]. non-standard [HB67]. Nonlinear [Ano66g, Box66b, Box66c, Bro69d, Myl66, Pow69a, Bro69c, CN63, Fie70, OW69, Tem69b]. nonlinearly [OM64]. NOR [Wil68b]. Normal [But61, Ada69, Woo69a]. normalarea [HJBH67]. Notation [Ham62]. Note [Ano60a, Ano62k, Bae62, Ber69, Che64, Che65b, Cob69, CND66, Col63a, Col63b, Day67, Ede65, Elc65, Eva62, EF63b, FFMR66, Fet65, Fie70, Flo64, GL64, Ham69d, HJBH67, HP67, Hil69, Hun60, Jon66, LW67c, LW68, Mat63, Mil60, OW66, OW68, Ove67c, Pag63b, PHJ67, PP69, Ran68a, Red71, Rei67, Rud67b, Ryd64, Sal70, Sco69, SS63a, Tem69b, Win69, Woo70, And66a, BY66, Bar69a, Bar69b, Bea65, Boo67l, Boo67m, Boo67n, Boo68b, Cha67, Chu68, Ham66b, Hou69b, Jar67b, Ran68b, RM69, Roh68, Wis68]. Notes [Ano66h, Ano67m, Cla60, Pap68b, Per65, Pap68a]. Notice [Ano63b, Ano66i, Ano69n, Ham66c, Tra66]. Novel [Tee63c]. Number [BT63, CH60, BW69a, Cob69, DR67, Hun60, MP68, PP69, Pre68, WP67]. numbers [Col67, Coo68a, Edm60]. numeral [Ros63b]. Numerical [BFMW64, Boh68, Cle63, Cle68, Dre66, Gre67, Gre69, Hun61, JS67, Joh66, Lin69, MP62, MLB63, Osb66, Phi67c, Phi67b, Pri66, SS63a, Tra66, Abb64, Aha65, All66a, BY66, Day65, Day66, Day67, Del68, Ell63, Hai69, LW66b, Mak68b, Mil60, MP67, New68, Osb64a, OM64, Osb69, Pak67b, Ral61, RM69, Ros63a, Scr64c, Tom67, Wol69, Wra66]. Numériques [Wil62b, Wil63a].

O [Ros63b]. O.C.R. [Pai69]. Obituary [Ano64e, Boy68]. Obtaining [Bye65]. OCR [WP69]. Oettinger [RP63]. off [Wal64]. office [Ano60a, Sut60, Ano60f]. offset [Pri68a]. On-Line [Bar69b, Bar69a, PL68, WM69]. One [Ano67g, Mus66b, Mus66a, Mus67, Pow69b, Sim69, BP69, Col63a, Day65]. One-day [Ano67g]. one-step [Day65]. One-Track [Sim69]. Ontario [Lio67]. Open [AS63]. Open-Shop [AS63]. Operating [Clo65, Cor62, Dij62, Gle62, Kin62, FS66, HPS61, KHPS61, Tur68]. operation [Cle67b, Rig64]. Operational [Mil63, FGR67]. Operations [Ham67a, Ham69d, Owe65, O’B67]. Operator [Tee63c]. Operators [Rei69]. Optical [Pai66]. Optimal [May69, Ste67]. Optimisation [Bro69a, Sav69, BW69b, Pal69, Sav70, Wes69]. Optimization [Fir61, Gea64a, Box65, Box66a, Fin68, Gui68, Bri66]. Optimizing [EF63a, MW68]. Optimum [Joh61, Car61, LM68, Ran68b, RW67, Rei66]. Orbitals [FR63]. Order [Boo67g, Hil69, Woo70, All66a, All66b, Boo67i, Dav67b, Eva65b, Eva66, Kar66, Mak68b, Mil60, RM69, Scr64c]. Ordered [Fos68a, Col63b, Eva64, Lyn64]. Ordering [Jen61, CND66, DM64]. ordinary [Cay61, CN63, Eng69, For61, Fox62, May64, Nor64, Oli69, Osb64a, Osb64b, OW68, RM69, Wri64]. ordinates [OSS69]. Organization [Bag61b, Buc60, MN61, Cas63, Ili64, KHPS61, Pit65, Pla61]. Oriented [Bag62, Bag61a, For66, FGR67, KM65, ML65]. Orion [TH64, Tay68b]. Orthogonal [CW61, Pow68, And69b, Hay69]. orthogonalising [Pal69]. other [ABN61, Boo67d, Box65]. our [Mus67]. outer [Wil68a]. outlook [Wil61b]. output [Dun63, Hoa63b, Zar69]. Over-Relaxation [EF63a, Tee63b, Ano64a, Car61, Eva62, Eva65a, LM68, Rei66]. overrelaxation [Ran68b]. own [O’B67]. Oxford [Cle65, Cle68].

PACE [PM62]. Package [Wil64b]. paged [Kni68]. pages [Cle65, Cle68, Wil67c, Cle66b]. Paging [Hig67, LWS69c, PB69, OL68]. Paid [Pol60]. Papers [Ano66h, Ano67m, Ano69i, Ano69j, Chu68]. Parabolic [Cle65, Eva65b, Eva66, LW66b, PC64]. parallel [Cle67b, Mil63]. parameter [DS67, OM64, Ran68b, Rei66, Tom67]. parameters [Mal61, Mic63]. Paris [Wil62b, Wil63a]. Parity [Mag64]. Parity-check [Mag64]. Park [Lei62]. Parker [Cle68]. Part [Fra61, Fra62, HPS61, KHPS61, Mus66b, Mus66a, Mus67, PL68, WN68, All66a, For66, FGR67, HS69, Mic63, Woo69b]. Partial [Fox63, Joh66, Cay61, Day66, Eva65b, Eva66, Osb63, Osb65, PC64, TW64]. partitioning [Car66]. partly [Col63b]. path [Boo67e, Boo67c, Boo67d]. Pattern [Dor66, Hal68, Ull65, Ull67]. Paulson [Ash68]. Pay [Sla60, TS64, Cas63, Moo63]. Peaceman [Gou68, ML65]. Pegasus [BM63, Gea61b, Gib61, Hor60, Ryd64]. Pensions [Ano60f, Pol60, Ano60a]. performance [Gos63, Yat62]. Pergamon [Cle65]. peripheral [Cle67b]. Permutation [Boo67j, Boo67k, Boo67h, NP68, Tew67]. Permutations [Boo67i, Boo67b, Hea63b]. Persistent [PC64]. Perspectives [Che65a]. PERT [How67a, Rob63b]. Phrase [BM60]. Physical [Ano67g]. Physics [BT63]. pi [Tee67c]. PICC [Ano61f]. Picture [Ede63]. Piecewise [Bic68, Bic69, Phi68a]. pitfalls [Pai69, WP69]. PL [Ano67g, TS68]. PL/1 [Ano67g, TS68]. place [Str61]. Plane [Col66a, Che65b, Lar64]. Planning [Ash67b, Lam63]. plans [War60]. plant [Cam69, You60]. played [Fin60]. plotter [Pit67]. plotting [Che65b, Le 69]. point [AH69, Bic68, Bic69, Fyf69, Has61, Kiz66a, Pri68a, Rya66, Smi64, SS63b, Wri66a]. Points [Ano66e, Jar67a, Kiz66b, Sne66, Nic66, Ord64]. Policy [Ham67f, Lew65a]. Polish [Col63a, Ham62]. polyhedra [Col66a]. Polynomial [Ham67a, O’B68c, OSS69, Riv68, Wat67, Bea68, GJM68, O’B68b, Phi68b, Pow67, Win69]. Polynomials [Cle68, CD69, Wil65b, Ede65, Gou68, Hay69]. polythetic [LW66a]. population [TB63]. positive [MT61]. postage [Lun69, Sel72]. Power [SD62, RW67]. pp [Wil62b, Wil63a]. practical [Dac63]. Practice [Tho60, Wes69]. Precision [How67b, CL66a, CL66b, SS63b]. Predicting [YA61]. Prediction [Lin68]. Predictor [Hal67, For61, Ral61]. Predictor-Corrector [Hal67, For61, Ral61]. Preparation [Pai66, Col64a, Dav63]. preprocessor [HB67]. presence [Hio67, Sum65]. Present [Wil61d, All66b]. Presidential [Mic64]. Press [Cle65, Cle68, Wil67c]. pretend [Mus66a]. Previously [Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c]. Price [Sla64b]. Prime [CH60]. Primitive [Wat69]. principal [Del68]. Principles [Bag62, Tho62, Vic65, Woo66c]. printed [YS61]. Prizes [Ano61a, Gil65]. Problem [Bro69a, Pap68b, Per65, Bag61a, Boh68, BP69, Car64, Day66, FM66, Hop61, How68a, Lar64, Law69, Lun69, Obr68, PW67, Sel72, Wra66, Wil66a]. problem-oriented [Bag61a]. Problems [Ano66g, Box66c, CP64, Dak65, Gil64, Gre67, Gre69, Pag63b, Rus63, Sla60, Sol60, Tho60, Tho66, dP60, Aha65, AH69, All66a, ABN61, Bag62, Bic68, Bic69, Box66b, Box66a, Col67, Fir61, For66, FGR67, Fos64, Fyf69, Gil65, GM65, LW66b, LT61, MP65a, MP66, Myl66, Osb64c, OM64, Pak67b, Pap68a, Pri68a, Smi66c, Str60, WP67, Woo69c, Osb66]. Procedure [LM63b, TS68, Tho62, GE69, HW67, Le 69, NP68, Pal69, Ros66a]. Procedures [Ham69d, BY66, Osb63, Wat63]. Proceedings [RP63, Ano61f]. Process [EF63a, Nig68, Tee63b, Ano64a, Erd65, MST64, Sla64a]. Processes [Cud67, May69, Bye65, Cur65, Hai69, Mil68a, RW62, Ros63a, Scr64b, Wol66]. Processing [Ano60e, Ano61d, Ano65f, Dun67, Har65b, Hin60, LR67, MP65b, Rus63, Sla60, WKRJ60, Bar65b, BFH61, CW62, DM60, Fos65a, Fri65, Gat65, Gil64, Hop61, Hou69b, Mil63, Wis64, YS61, Dou66, Har65a, Gea64b]. Processor [WH68, Bro69b, GB69, Tru66]. producing [ABN61]. Production [War60, Hug61]. products [GL64]. profile [BW69b]. profitability [Jan63]. Program [ApS65, Bar69b, BHS65, BM60, Bur69a, Cad61, Hur65a, Hur65b, KB62, KV63, New66, Sol60, Wet66, YS60, And66a, AMKM66, BH63, Bar69a, BSPI65, BM61b, Col66b, Dol66, Fos68b, Gib60, HLOS65, How67a, LW68, Lio67, Lum64, Str65d, Str65e, SS63b, Und63]. Programmer [Wil64a, Smi68]. Programming [Ake66, Ano60g, Ano66g, Box66c, Boy60, Dak65, Fra67, Hal66, Hea63a, Mor69, Phi66, Pow69a, RP63, d’A62, Box66b, BFH61, Bur67, BP69, Cas63, Che65a, Fie64, Gri66, KM65, Law69, MW68, Myl66, Phi67a, Pri63, Sla64a, Vaj64, Wil68a, BT63, Cra66a]. Programming-Sequential [Pow69a]. Programs [Ber66, Bur69b, McO66, Cam69, DN64, Gib61, JP69, LW65, LW66a, Mul69, Nie65]. Progress [Bar66b, Owe65, Pol60, Sam66b, Rou62]. Progressive [LM63b]. Project [Nud69, Lam63]. Projection [Tew69c]. projections [Col66a]. promise [Yat62]. proof [Woo69a]. Properties [Bur69b, Lyn65, Bar61, Con66]. proposals [BM61b, Nap67]. proposed [CP62]. prospects [LT61]. Protection [McO66]. Prototype [Gre69]. Proving [Bur69b, Mel66]. pseudo [DR67, Edm60, Hun60, May66, MP68]. pseudo-inverse [May66]. pseudo-random [DR67, Edm60, Hun60, MP68]. psychology [Pla64]. Publication [Ano61f, Ano66i, Ham66c]. published [Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c]. punched [cGR61]. purpose [Jon65a, Str65c, Und63]. Purposes [Gri62]. Puzzles [Bur69a].

quadratic [Ord64]. Quadrature [Sac62, Ste67, Bar61, Cha68b, JS67, LM63a, MLB63, OS68a]. Quadratures [Tha62]. Quality [MN61]. quasi [Cay61]. quasi-diagonal [Cay61]. queuing [Str60]. Quicksort [Hoa62a]. quintic [Sal70, Spä69b].

R [RP63, Wil67c]. Rachford [ML65, FM64, Gou68]. radar [JP69]. radiational [MP65a, MP66]. radix [Woo69d]. RAMPS [Lam63]. Random [But61, KV63, BW69a, DR67, Edm60, Hun60, MP68]. Rank [Pow69b, Boo67f, Hil69, Woo70]. Rapid [OW66, Cra66b]. Rapidly [FP63]. Rapidwrite [Ayr64, Hum62]. rates [Eva65a]. Ratio [BB69, Ash68]. Rational [Jar66b, Lar67, Tha66, Cob69, CO66, JN65, PP69, Pre68]. reactions [Bae64c]. readability [Hum62]. Readings [Dun67]. Real [Gri62, Phi66, CL66a, Cob69, FJ66, GRB68, Hop61, Par67, Phi67a, PP69, Pre68, Wil65d]. Real-Time [Phi66, Phi67a]. realization [BM61b, Wat63]. really [Bae64c]. Rearrange [Boo67h]. Rearranging [Win60]. Recognition [CP61, Dor66, Hin61, New61, Pai66, Sha65, Ull65, Bau65b, GB61, Str61]. Record [Tho64a]. Recording [Jen61, Pol60, Dru63]. records [Gil66]. rectangular [Phi67c]. recurrence [MB66]. Recursive [Coo69]. Reducing [Wil62c, AM65, OS65b]. reduction [Boy63, SF61, Tew69a]. Reductions [Rei67]. redundant [OS65b]. Reed [PCSG63]. reflected [Flo64]. regarding [Gil66]. region [Phi67c]. regions [LM68]. Regression [Sla62b, Hea63a]. related [Wil65c]. relating [Ral61]. relationships [Kin69a]. relative [Loi68]. Relaxation [EF63a, Tee63b, Ano64a, Bye65, Car61, Eva62, Eva65a, LM68, Rei66]. Reliable [Sum65]. remarks [Fin60]. Remote [AS63, MB66]. renaming [Mel66]. repeating [Hun60]. replicated [YGS63]. Report [Hoa62b, Pol60, BBG^{+}63, Rou62]. representation [Bar64, GS63, Sam63c, Tay64]. requirements [Boy63, Sam68c]. Research [AS63, Cra66a, Fox61b, JJ67, Owe65, Wex69, Gos60, Las68, Nix64, Yat62, Wri66c]. reservations [Geo63]. resolution [Mel66]. Resource [Lam63]. Resourcing [Bae66b]. respect [Con66]. Result [Ano61a]. Results [MM60, FM66, Mel66]. Retrieval [CH60, Gee66, FE69, Llo64, Nee63b]. reverse [Col63a]. Review [Alb67, Alb68, Ale69, And66b, Ano60b, Ars65, Ash67a, Bae64a, Bae66b, Bae67, Bar62, Bel67a, Bel68, Bel69a, Bel69b, Box67, Bri65, Bri66, Bro68, Bur66a, Car63, Cha68a, Che63, Chu69, Cla62, Cle67a, Cle65, Cle66b, Cle68, Clu63a, Col64c, Coo69, Cra66a, Cur68, Cur67, Dav67a, Dir68, Dor67, Dor68, Dou69, Dou66, Eve68, Eve69, Fie64, Fle65, Flo68a, Flo68b, Fox61a, Fox68, Fra69a, Ful66a, Ful66b, Ful67, Gea61a, Gea64b, Gea67, Gee66, Gee67, Gil69b, Gil68, Goo60, GT63a, Gre65, Hag69, Ham66d, Har65a, Har63, Hea61, Her61a, Hio68a, Hou63a, How65, How68b, Hug67, Hum67, Ili64, Jon64, Jon65b, Jon69, Ken65, Kir63, Lan68, Lav68, Lay65, Lev63, Llo68, Llo69, Mag64, Mag65, May69, McM68, Meg62, Mil62, Mit69, Mur68, Nee61a, Nee61b, Nee63a, Nel68, New65a, Nix64, Oat63]. Review [OS65a, OS67a, OS67b, OS68b, Osb66, Pay66, Pig64, Pow65b, Pri67, Pri68b, Pri61, Rei69, RP63, Rob67, Rob63a, Rog65, Ros62, Ros68, Sam68a, Sam68b, Sha67, Sla62a, Spe65, Str64, Tay61b, Tee67b, Tiz62, Tiz64, Tiz65, Toc64, TW61, Tur66, Vaj60, Vaj64, Vic67a, Vic67b, Wat69, Wat68a, Wes64, Wes69, Wil61a, Wil62a, Wil66a, Wil69, Wil62b, Wil63a, Wil67c, Wil64c, Wil66b, Wis61, Wol66, Wol68, Woo62, Woo66a, Wri66c, You63b, You68a, You68b, Gil63a, GT63b, Her61b, Hoa64, Hou63b, Nee63b, Pri63]. Reviews [Bae64b, Bra63a, Cle63, Clu63b, Gat67, Gil62, Goo63, Hoa63a, Hop62, How62, LW62, Mar62, Mil64, Pag63a, Per63, Ric63, Sam63a, Smi63a, Str62, Tee63a, Wal62, BT63, PCSG63, RP63]. Revised [BBG^{+}63]. Revisited [Wat08]. Rhodesia [Che62a]. rigid [Duk66b, Jon66, Rac65]. Rigorous [Wil61c]. Ring [WH68]. Ringing [Pap60]. Robert [BT63]. Rogers [Her61b]. Romberg [Fox67]. Rome [Ano61f]. Roots [Eve63, Han62, DP64, Mal61, Wat69]. Rosenbrock [MR65, Pal69]. rotational [Duk66b, Duk66a, Jon66, Rac65]. Roth [BH68]. round [Wal64]. round-off [Wal64]. Route [Pap68b, Per65, Nic66, Pap68a]. Routine [Pag60, PM62]. Routines [BMR62, GL70, Lyn69, Sol68b]. Row [Tew67]. Row-column [Tew67]. Rows [Boo67b]. Royal [Dav63, Sla60, TS64]. rules [CND66]. Run [KV63]. Runge [Cur65, Day66, Eng69, Kar66, Scr64b]. Running [Gib61]. runs [Dow69]. Russian [GT63b].

S [Ano69e, PCSG63, Wil67c]. S.O.R. [LM68, O’C69]. S.S.O.R. [DM64]. SABER [Hop61]. salesman [Obr68]. Sampling [But61, Tur66]. SAS [Geo63]. Satellite [Pea63]. satisfying [Mil60]. Saul’yev [Cle65]. Scale [Sla60, Woo69c, You60]. scaling [PM62]. Scanner [MN61, GB61]. Scattered [Sne66]. Schedule [KV63]. Scheduling [Pag62, Tho66, ABN61, HJW62, Kni68, Lam63, Lio67, NP68, PW67]. Scheme [Bar67, IJ62, Pol60, Bar65b, DM64, Fra70, Hor60, Jen66, RW61]. schemes [Win69]. School [Ano62d, ABN61, Law69, Lio67]. Science [Pay66, Mic64]. Sciences [Bur66a]. scientist [Mic64]. Search [Dak65, Boo79, Ove67c, Pal69, PHJ67, PP79]. searching [Day69]. Seasonal [Hio67]. secant [Bar65a]. second [All66b, Hug61, Mil60, Scr64c, All66a]. second-order [All66b, Scr64c]. Section [Tho60, Boo67g, Boo67h, Boo67f, Hil69, Woo70, dP60]. Seeking [Sne66]. Segmentation [Hig67, Las68, DN64]. Selected [Ano66e, Kiz66b, Kiz66a, Wri66a]. Selective [Mat63]. selfmodifying [Nie65]. Selling [Bro69a]. Semi [Tee63b, Ano64a, Rei69]. Semi-Groups [Rei69]. Semi-Iteration [Tee63b, Ano64a]. Separable [Ake66, Osb63, Osb65]. separated [Osb69]. September [Ano62c]. sequence [MB66]. Sequential [Coo68b, Ped66, Pow69a, LW66b]. Series [Boo68a, GM65, Jar68, Mat63, Wri66b, Bea65, Cha67, CN63, Ede65, Ell63, Gib60, Nor64, Oli69, Pin61, Scr65, Wol69]. service [Bag61b, Kin62, Pit65]. Session [Ano62g, Ano62i, Ano62h, Ano62j]. set [Con66, Ros64a]. Several [Pow62, Pow64, Box66a]. shared [Bae66a, Bae66c, LWS69a, Pen66]. sharing [Coo60b, Lew63, Mil63]. Shop [AS63, NP68]. Shortest [Boo67c, Boo67d, Pap68b, Per65, Boo67e, Nic66, Pap68a]. should [Hig63, Hil67]. shown [Ano61h]. Signals [Smi66b]. significant [Ber69, Sal68]. similarity [LW66a, Ull67]. simple [BBB68, Col66b, Fos65a]. Simplex [NM65]. Simulate [WY62]. Simulated [Gri62]. Simulation [BL62, Cle66a, HG67, HB66, LEN63, Lei62, Wil67b, You60, Bae66c, Bea68, Cam69, Che64, DP66, Lew65a, Pen66, RWT64]. Simulations [KB62, Bux66]. Simulator [Dem60, Wil64b, BBS66]. Simultaneous [Ano60g, Boy60, Bro69d, Wri63, BSD60, Bro69c, Fie70, Jen66]. single [Col63a, PL68]. single-address [Col63a]. singular [DP64, Fox67, May66, Tew69b, WP68, Wol69]. singularity [Cha67]. sites [MW68]. Siting [Lev65]. size [Eva62]. slip [MP67]. Small [BHRS69, Che62b, Coo60a, MP65b, WH68, Col64a, Eva62, HRS68]. small-store [Col64a]. Smoothing [CD69, dW64]. SNAP [Nap67]. Society [Ano61d, Ano62c, Ano63b, Ano65f, Ano69n]. SODA [Tur68]. sofa [How68a]. Software [Mul69, Bro69b, Nef65, Sam68c, Wil67d]. soldier [Cas63, Moo63]. Solution [BSD60, Bro69c, Cay61, DS67, Fie70, Fle68, Joh66, MP62, Rei67, SS63a, Tho68b, Wat67, Wil60, Aha65, BFMW64, Bak64, Bir68, Boh68, Car64, CN63, Day66, Day67, Eg69, Ell63, EF63b, Eva65b, Eva66, Fir61, For61, Fyf69, GM65, Hai69, Has61, JS67, JN65, Jen66, Kar66, LW66b, LM68, Loi68, Mak68b, Nor64, Oli69, Osb64a, Osb64c, OM64, Osb65, Osb69, Ran68a, Ros63a, Scr64c, Scr65, WP68, Wol69, Wra66]. solutions [AH69, Bye65, Eng69, Wil62b, Wil63a]. Solving [Bro69d, Bur69a, Jar66b, Kha63, O’B68c, Tew69c, Wri63, All66a, All66b, Bar65a, Dav67b, Eva63, FM64, GJM68, Had68, Jar66a, Jar67b, MP67, O’B68b, Osb63, OW68, Pak67b, Pri68a]. Some [Bon69, Bro60, BM61b, CW61, CR63, Dem60, Dru63, FM66, Fin60, Gil66, Lar67, LR67, Nap67, New61, Pap68a, Pap68b, Per65, Ral61, Ros63a, Sla64b, WP68, Abb64, All66a, Chu68, FK68, Gea61b, Goo61, Kin67b, Mel66, Pak67b]. SOR [Lyn64]. Sort [Boo67f, Hil69, Woo70, Cra66b, Woo69d]. Sorting [Fos68a, Col63b, LW67a, LW67b]. source [Elc65]. source-language [Elc65]. sources [DGM68]. Spanning [Obr68]. Sparse [Kha63, Tew69c, NE67, Tew67, Tew69a]. Special [Boo68b, Coo68a, GB61, Jon65a]. special-purpose [Jon65a]. specialized [YGS63]. specifying [Gri66]. SPECOL [Smi68]. Spectra [SD62]. Speech [LR67, Woo66d]. speeding [Erd65]. Spline [Sco69, Spä69a, AH69, BY66]. splines [Fyf69, Sal70, Spä69b]. spot [GB61]. Square [Eve63, Kha63]. Squares [Cad61, Fle68, Pow65a, Rei67, JM68, Pow67, Tew69b]. SSOR [Lyn64]. Stability [Hal67, Kar66, MP67, Ros66a, Wat68b]. Stable [Mak68b, Eva65b, Eva66]. Staff [YA61]. stage [You60]. stamp [Lun69, Sel72]. standard [HB67]. Standards [Har65a, Har65b]. start [Boo67e, Boo67c, Boo67d]. Starting [Eve63]. State [Sla66, Cla60]. Statement [Ham67f]. Statements [MP65b]. Stationary [Ord64, Pow62]. statistic [LW68]. statistical [DM60]. Statistics [Gri62, Ped66, O’B67]. steady [Abb64]. steelworks [Mas63]. Stefan [Wra66]. step [Day65, GRB68]. stereographic [Col66a]. stochastic [Mil68a]. stock [Pal62]. stock-control [Pal62]. Stokes [Bye65]. Storage [Bro60, Fra69b, IJ62, Boy63, HW67, Jen66, Kni68, Lan65]. store [BSD60, Cha62, Col64a, DN64, RW61]. storing [Hay69]. Stove [RWT64]. straight [Phi68a, Spä68]. strategies [LW67a, LW67b]. strategy [NE67]. stratifying [Bag61a]. String [Mil68b]. Structural [Bur69b, Liv60]. Structure [BM60, WH68, Bri67, HLN68, Wis68]. Student [TS68]. Studies [Gre69]. Study [Hou63b, KV63, Car64, Lan65]. Sturm [GM65]. subdivision [OS69]. subject [OSS69, Osb69]. Submission [Ano66h, Ano67m]. subroutine [Coo68a]. subscript [Rya66]. subscripts [Boo67g, Boo67h, Hil69, Woo70]. Successive [EF63a, Tee63b, Ano64a, Boy63, Car61, Ran68b, Rei66]. Suitable [Kha63]. Sum [Bur69a, Pow65a]. SUMFAC [Boo67a]. Summation [Mat63]. supervisor [HLN68, LWS69a]. Supplement [Ano66a, Ano67a, Ano67b, Ano68a, Ano68c, Ano69a, Ano69b, Ano69c, Ano69d, GT63b, Ham66a, Ano67c, Ano68b, Ano68d, Ham67b, Ham67c, Ham67d, Ham67e, Ham68a, Ham68b, Ham68c, Ham69a, Ham69b, Ham69c, Ano66i, Ham66c]. supply [Bur66b, Dav63]. Surface [Cad61, CW61, Hay69, Thr69]. surfaces [MP67]. Survey [Lon61, PCSG63, And66a, Red60]. Surveys [Coo60a, YS60, YS61]. Swann [Pal69]. Switching [Bel69c, Ste66]. symbol [Luk67]. Symmetric [EF63a, Abb64, FJ66, Jen66, MT61, OK63, Par67, PW69, Ran68a, RW61, Sea69, Wil65d]. symmetrical [AM65]. Symposium [Ano61f, RP63, Ano67g, Ano61h]. Syntax [Sim69, CND66, Fos68b, FE69, FK68, Ree67, Sch65, Wil68a, Zar69]. syntax-analyzers [CND66]. syntax-directed [Ree67, Sch65]. syntax-driven [FE69]. Synthesis [Mar69, Woo66d, Dol66, Wil68b]. System [Ano67l, Fra69b, Gee66, Gri62, Hic60, Jen61, KP60, Rei67, Wil64a, Bae66a, Bae66c, Bar66a, BBB68, BHJV66, Cle67b, Cra66b, Dav63, For66, FS66, Fra68, Hoa63b, HPS61, HJW62, Jan63, KHPS61, LWS69a, LWS68, Mal61, MR67, NPW68, O’B67, Pal62, Pen66, Ros66b, Sam66a, TH64, Tur68, WM69, Wis64, Woo68, BW69a]. Systematic [Ano66j]. Systematics [BBdD66, Gri66, Gri68, Kin67b]. Systems [Ano63c, Bur68, Clo65, Cra66a, Fle68, H. 68, HG67, Kin69b, Liv60, Nig68, Phi66, Rus63, Smi66c, Smi66b, Tew69c, Tur66, Ano63a, Bak64, BS61, Bri67, Cay61, Cor65, DS67, Eng69, EF63b, Eva64, For61, Goo64, GE69, GlM68, Gri66, Kar66, LW67a, LW67b, Lav64, Lyn64, Mak68b, Rig64, RM69, Ros66b, Str61, Vin64, Wil67a, WN68, Hou63b, PCSG63].

T [Her61b]. Table [OW66, Car70, Kin69a, Kin71, Mil60]. Tables [Gri68, Col64a, Gow62, Kin67a, Mil64, YS61]. tabulate [Cra66b]. TANT [Mar69]. tape [BSD60, Cha62, DGM68, EAL66, Gil66, Goo64]. tapes [cGR61]. target [CP62]. Taylor [Gib60]. Teaching [Fox61b]. team [Bri67]. Technique [CP61, Gil63b, OB68a, Sav69, Wet66, Che65b, Dav67b, Eva65a, Fin68, Fos64, MW68, New68, Oli69, Sav70, Woo69c]. Techniques [ABN61, BH63, Bro60, How67b, Hus61, Lar67, Mor69, PCSG63, Pow69a, Wil67b, Hur68, Str60, PCSG63]. Technology [Pri64]. Tee [Ano64a]. TELCOMP [Mul69]. tele [Hop61]. tele-processing [Hop61]. telecommunication [Geo63]. teleprinter [GS63, Sam63c]. Television [Bro69a]. term [GB69, Rya66]. Terminal [Sac62]. terminology [GT65]. terms [Gat65, MB66]. ternary [Hur68]. Test [Ano66g, Box66c, Box66b, Dow69, Hun60, Myl66]. testing [BW69a, Tay61a]. tests [Gro68]. Textbook [Nee63b]. Their [RP63, ABN61, Elc65, May69, NE67, You63a]. Theorem [Mel66, Pow69b, Goo68a, Woo69a, Goo68b]. Theorem-proving [Mel66]. theoretic [Fos64]. Theoretical [Pag60, Bae66c, Pen66, Ral61]. Theory [Bel67b, Bel69c, Bur68, Coo68b, May69, Ras66, Ste66, Tho60, Woo66c, Fos65b, LW67a, LW67b, Mil68a, Woo69b, Bri66, Coo69, Gee66, Ham66d, PCSG63, Tur66]. Thinking [Wal69]. thinning [Deu69, Fra70]. Thomas [PCSG63]. Three [Phi67b, Che65b, Rya63, Sco69, Spä69a]. three-dimensional [Che65b]. Threshold [Wil64c]. Time [Bro69a, Coo60b, Hex67, Lew63, Phi66, Tho68c, Wol66, Bae66a, Bae66c, Col64a, GE69, Hop61, Joh61, LWS69a, Mil63, Pen66, Phi67a, RW68, Tho68a]. Time-Dependent [Wol66]. time-shared [Bae66a, Bae66c, LWS69a, Pen66]. Time-sharing [Coo60b]. time-tables [Col64a]. Times [Nud69]. timetable [Alm69, DD70]. Timetables [Ano62d, Alm66, ABN61, Bar65c, FL68, Woo68, Yul68]. Timetabling [Tho66, Col67, Law69, WP67, Woo69c]. timing [Che64]. Titan [LWS68]. tool [Lam63]. top [Duk66b, Jon66, Rac65]. topology [Las68]. towers [MW68, RW67]. Track [Sim69]. Trains [KV63]. transfer [Bea68]. Transformation [Fra61, Fra62]. transformations [Box66a]. Transforms [Bel63]. translate [Mul69]. Translation [Ham62]. Translator [Hoa62b, DH60, McE67, Pul64, Ree67]. Transmission [Ano63c, Ano65a, Lon61, Rus63, Smi63b, Wil61d, Woo66b, Wri61a, Ano63a, Dac63, Dav63, LT61, MW68, RW67, Str61, Wil61b]. transmitting [Wil63b]. Transport [Wol66, PW67]. Transportation [Hal66]. travelling [Obr68]. Tree [Dak65, Obr68]. Trees [BMR62, Lyn65, Win60, Obr67]. trend [Hio67]. Tri [Wil62c, Bak64, EF63b]. Tri-Diagonal [Wil62c, Bak64, EF63b]. trials [BS61]. tridiagonal [FJ66, OK63, Par67]. trigonometric [Hun68]. truncation [Cur65, Scr64b, Smi66a]. TSAS [LWS69a]. tunnel [Vic65]. turbines [Lan65]. Turning [Jar67a]. twenty [Chu68]. twenty-five [Chu68]. Two [Bro60, Str60, AH69, Bic68, Bic69, DN64, Fyf69, Has61, Nic66, Obr67, OW68, Phi67b, Pri68a, RS69, RW61, HP67]. Two-Level [Bro60, DN64, RW61]. two-point [AH69, Bic68, Bic69, Fyf69, Has61, Pri68a]. TWOBYTWO [HP79]. Type [Cle65, Hex67, Sac62, Eng69, FM64, LM68, PC64]. Type-Matching [Hex67]. types [Obr67]. typesetting [Gol69].

U.K. [Wil67d]. U.S.S.R [BT63]. U.S.S.R. [Cla60]. Unconstrained [Pow69a]. underground [Lan65]. Unit [JJ67]. Unitary [Fra61]. units [Gla69, Wil68b]. Universal [Bag62]. universities [Sam68c, You63a]. University [Buc60, Buc65, Cle68, DD70, Dol66, HPS61, KHPS61, Wil67c, WKRJ60, Woo68, Alm66, Alm69, Yul68]. upper [WP67]. Use [AS63, BBS66, Cha63, Clo65, Dou65, Gri68, Jar68, BH68, BS61, Box66a, Col66b, Cra66a, Dac63, Fyf69, Gea61b, GB69, Hil67, Hoc62, JN65, LM63a, Wra66]. used [Gat65, Hay69]. Useful [Wil65a]. User [Fra68, Gol64, Ric64, Bae64c, Cor65, HPS61, KM65, Lav64, Smi64, Ayr64]. uses [Mus66b]. Using [BHS65, Bro69b, Wat67, BSD60, Col64a, Ede65, FK68, GS63, Gib60, GRB68, GB61, Hop63, JS67, Mar69, Pal62, Red60, Sam63c, Sch65, Sea69, Tho64a]. USSOR [Lyn64]. utilisation [Gla69]. Utility [Wil64a].

V [Cle65, Hou63b]. Value [Osb66, AH69, Bon69, Car64, Del68, Fyf69, Pri68a, Ros60]. Values [Pow62, Tee67a]. Variable [Pea69, HW67, Tho64a]. variable-length [HW67]. Variables [Pow62, Pow64]. Variance [BB69, Dav68, Ash68]. Variate [Col64b]. variation [DS67]. variations [All66a, Pak67b]. variety [Hop63, Pai65]. vector [Aha65, Boo67k, Boo67i, Boo67j, Pri68a]. Vectors [Pow68, And69b, DP64, Pal69]. versus [Nie65]. VI [HS69, Hea68, HL70, Wel70]. Vibration [Wil67c]. view [Lav64, Smi64]. Views [Smi63b]. virtual [Las68]. viscous [Abb64]. Vocabulary [Dou66, Gat65]. Volterra [JS67, Lin69, WP68]. Volume [Ful66a, Ful66b, RP63, Wil63a].

W [BT63, GT63b, Her61b, PCSG63]. Wang [PCSG63]. want [Bae64c]. Watt [Gen69]. wave [Bea68]. which [Fin60, OM64]. William [MM69]. Windscale [WM69]. within [Gri68]. Without [Pow65a, Fle65, Pow64, Zan67]. Word [Bur69a]. words [Coo68a]. Work [Ano61h, Che62b, GT65]. Working [Ham69d]. world [HS69]. Writing [Bux66, For65].

year [Hug61, Sut60, War60]. years [Rya63]. yourself [Nef65]. Yovits [PCSG63].

Zero [Weg62, BP69, Ros63b]. Zero-Address [Weg62]. zero-one [BP69]. zeros [Che65b, Lar64]. Zubov [Hou63b].

References

Aleksander:1968:ALC

[AA68] I. Aleksander and R. C. Albrow. Adaptive logic circuits. The Computer Journal, 11(1):65–71, May 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/65.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110065.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
65.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/66.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/67.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
68.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/70.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
71.tif.

Abbott:1964:ASS

[Abb64] M. R. Abbott. Axially symmetric steady motion of a viscous incompressible fluid: some numerical experiments. The Computer Journal, 7(1):47–53, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/1/47.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070047.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/48.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/50.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/51.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/52.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/
53.tif.

Appleby:1961:TPS

[ABN61] J. S. Appleby, D. V. Blake, and E. A. Newman. Techniques for producing school timetables on a computer and their application to other scheduling problems. The Computer Journal, 3(4):237–245, January 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/237.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030237.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
237.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/238.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/239.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/240.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/241.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/242.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/243.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/244.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/245.tif.

Adams:1969:AAU

[Ada69] A. G. Adams. Algorithm 39: Areas under the normal curve. The Computer Journal, 12(2):197–198, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/197.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/198.tif.

Albasiny:1969:CSS

[AH69] E. L. Albasiny and W. D. Hoskins. Cubic spline solutions to twopoint boundary value problems. The Computer Journal, 12(2): 151–153, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/151.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120151.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/151.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/152.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/153.tif.

Ahamed:1965:ACN

[Aha65] S. V. Ahamed. Accelerated convergence of numerical solution of linear and non-linear vector field problems. The Computer Journal, 8(1):73–76, April 1965. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/8/1/73.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_01/080073.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/75.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/76.
tif.

Akeroyd:1966:ASP

[Ake66] A. J. Akeroyd. An application of separable programming. The Computer Journal, 8(4):344–346, January 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/344.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080344.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
344.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/345.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/346.tif.

Albasiny:1967:BR

[Alb67] E. L. Albasiny. Book review. The Computer Journal, 10(3): 270, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/3/270.full.pdf+html.

Albasiny:1968:BR

[Alb68] E. L. Albasiny. Book review. The Computer Journal, 11(1):76, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/76.full.pdf+html.

Aleksander:1969:BR

[Ale69] I. Aleksander. Book review. The Computer Journal, 12(2):147, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/147.full.pdf+html.

Allen:1966:IDN

[All66a] B. T. Allen. An investigation into direct numerical methods for solving some calculus of variations problems. Part 1 — Second order methods. The Computer Journal, 9(2):205–210, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/2/205.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_02/090205.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/205.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/206.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/207.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/208.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/209.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

09/Issue_02/tiff/210.tif. See correspondence [Pak67b].

Allen:1966:NMS

[All66b] B. T. Allen. A new method of solving second-order differential equations when the first derivative is present. The Computer Journal, 8(4):392–394, January 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/392.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080392.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
392.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/393.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/394.tif.

Almond:1966:ACU

[Alm66] Mary Almond. An algorithm for constructing University timetables. The Computer Journal, 8(4):331–340, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/331.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080331.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
331.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/332.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/333.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_04/tiff/334.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/335.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/

tiff/336.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/337.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/338.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/339.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/340.tif.

Almond:1969:UFT

[Alm69] Mary Almond. A University faculty timetable. The Computer Journal, 12(3):215–217, August 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/3/215.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/120215.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/
215.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/216.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/217.tif.

See correspondence [DD70].

Alway:1965:ARB

[AM65] G. G. Alway and D. W. Martin. An algorithm for reducing the bandwidth of a matrix of symmetrical configuration. The Computer Journal, 8(3):264–272, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/264.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080264.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
264.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/265.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/266.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/268.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/269.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/270.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/271.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/272.tif.

ApSimon:1966:CIP

[AMKM66] H. G. ApSimon, W. D. Maurer, P. J. H. King, and C. H. R. Morris. Correspondence: An impossible program. The Computer Journal, 8(4):329–330, January 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/329.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/329.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/330.tif.

See [Str65e, ApS65, HLOS65, BSPI65].

Anderson:1966:NCG

[And66a] A. J. B. Anderson. A note on the construction of a general survey program in Extended Mercury Autocode. The Computer Journal, 8(4):312–314, January 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/312.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080312.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
312.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/313.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/314.tif.

Andrew:1966:BRB

[And66b] A. M. Andrew. Book review: Learning Machines. The Computer Journal, 9(1):26, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/1/26.full.pdf+html.

Andrew:1969:C

[And69a] A. M. Andrew. Correspondence. The Computer Journal, 12(4): 411, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/411.1.full.pdf+html.

Andrew:1969:CCO

[And69b] A. M. Andrew. Correspondence: The calculation of orthogonal vectors. The Computer Journal, 12(4):411, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/411.tif. See [Pow68].

Anon:1960:NCL

[Ano60a] Anon. Note on commissioning of LEO automatic office at the Ministry of Pensions and National Insurance. The Computer Journal, 2 (4):198, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_04/tiff/198.tif.

Anonymous:1960:BRC

[Ano60b] Anonymous. Book review (concluded). The Computer Journal, 3 (1):20, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/1/20.full.pdf+html.

Anonymous:1960:C

[Ano60c] Anonymous. Correspondence. The Computer Journal, 3(3):174, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/3/174.full.pdf+html.

Anonymous:1960:HAC

[Ano60d] Anonymous. Handbook for automatic computation. The Computer Journal, 2(4):169, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/2/4/169.full.pdf+html.

Anonymous:1960:IP

[Ano60e] Anonymous. Information processing. The Computer Journal, 3 (2):75, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/2/75.full.pdf+html.

Anonymous:1960:LAO

[Ano60f] Anonymous. Leo Automatic Office at the Ministry of Pensions and National Insurance. The Computer Journal, 2(4):198, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/2/
4/198.full.pdf+html.

Anonymous:1960:SEL

[Ano60g] Anonymous. Simultaneous equations and linear programming. The Computer Journal, 3(1):50, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/3/1/50.full.pdf+html.

Anonymous:1961:APR

[Ano61a] Anonymous. Annual prizes: Result of 1960-61 competition. The Computer Journal, 4(3):211, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/4/3/211.full.pdf+html.

Anonymous:1961:AMM

[Ano61b] Anonymous. Automation: Men and money. The Computer Journal, 3(4):201, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/4/201.full.pdf+html.

Anonymous:1961:C

[Ano61c] Anonymous. Correspondence. The Computer Journal, 4(3):229, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/3/229.full.pdf+html.

Anonymous:1961:DPS

[Ano61d] Anonymous. Data Processing Society of the Netherlands. The Computer Journal, 4(2):136, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/4/2/136.full.pdf+html.

Anonymous:1961:ECE

[Ano61e] Anonymous. Electronic computer exhibition. The Computer Journal, 4(1):63, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/63.full.pdf+html.

Anonymous:1961:FPP

[Ano61f] Anonymous. Forthcoming publication of the Proceedings of the 1960 PICC Symposium, Rome. The Computer Journal, 3(4):261, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/4/261.full.pdf+html.

Anonymous:1961:IC

[Ano61g] Anonymous. Ifip congress 62. The Computer Journal, 4(1):19, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
1/19.full.pdf+html.

Anonymous:1961:WNG

[Ano61h] Anonymous. Work of ‘new generation’ to be shown at London Computer Exhibition and Symposium. The Computer Journal, 3 (4):245, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/4/245.full.pdf+html.

Anonymous:1962:Aa

[Ano62a] Anonymous. Appendix. The Computer Journal, 5(2):106, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
2/106.full.pdf+html.

Anonymous:1962:Ab

[Ano62b] Anonymous. Appendix 2. The Computer Journal, 5(3):199, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/199.1.full.pdf+html.

Anonymous:1962:BCS

[Ano62c] Anonymous. British Computer Society Conference, Cardiff: 4th to 7th September 1962. The Computer Journal, 5(2):156, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
2/156.full.pdf+html.

Anonymous:1962:CST

[Ano62d] Anonymous. Computers and school timetables. The Computer Journal, 5(3):247, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/5/3/247.full.pdf+html.

Anonymous:1962:Ca

[Ano62e] Anonymous. Correspondence. The Computer Journal, 5(2):154a– 155, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/2/154-a.full.pdf+html.

Anonymous:1962:Cb

[Ano62f] Anonymous. Correspondence. The Computer Journal, 5(3):246– 247, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/3/246.full.pdf+html.

Anonymous:1962:DSAa

[Ano62g] Anonymous. Discussion: Session 1: 17 April 1962 (Morning). The Computer Journal, 5(2):119–125, August 1962. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/5/2/119.
full.pdf+html.

Anonymous:1962:DSAc

[Ano62h] Anonymous. Discussion: Session 2: 17 April 1962. The Computer Journal, 5(3):170–176, November 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/170.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/170.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_03/tiff/171.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_03/tiff/172.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/173.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/174.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/175.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/176.
tif.

Anonymous:1962:DSAb

[Ano62i] Anonymous. Discussion: Session 3: 18 April 1962 (Morning). The Computer Journal, 5(2):135–139, August 1962. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/5/2/135.
full.pdf+html.

Anonymous:1962:DSAd

[Ano62j] Anonymous. Discussion: Session 4: 18 April 1962. The Computer Journal, 5(3):183–192, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/5/3/183.full.pdf+html.

Anonymous:1962:EN

[Ano62k] Anonymous. Editorial note. The Computer Journal, 5(3):244, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/3/244.full.pdf+html.

Anonymous:1962:E

[Ano62l] Anonymous. Errata. The Computer Journal, 5(3):176, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
3/176.full.pdf+html.

Anonymous:1962:MIC

[Ano62m] Anonymous. Midsummer International Conferences. The Computer Journal, 5(1):27, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/5/1/27.full.pdf+html.

Anonymous:1963:DFS

[Ano63a] Anonymous. Discussion forum — the systems approach to data transmission. The Computer Journal, 6(3):228–231, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/6/
3/228.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/228.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/229.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/230.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/231.
tif.

Anonymous:1963:NMB

[Ano63b] Anonymous. Notice to members of the British computer society. The Computer Journal, 6(3):292, November 1963. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/6/3/292.
full.pdf+html.

Anonymous:1963:SAD

[Ano63c] Anonymous. The systems approach to data transmission. The Computer Journal, 6(3):209, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/6/3/209.full.pdf+html.

Anonymous:1964:CES

[Ano64a] Anonymous. Correction to “Eigenvectors of the successive overrelaxation process, and its combination with Chebyshev semiiteration”, by G. J. Tee. The Computer Journal, 7(1):39, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
1/39.1.full.pdf+html. See [Tee63b].

Anonymous:1964:C

[Ano64b] Anonymous. Correspondence. The Computer Journal, 6(4):335, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/4/335.full.pdf+html.

Anonymous:1964:E

[Ano64c] Anonymous. Errata. The Computer Journal, 7(1):39, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
1/39.2.full.pdf+html.

Anonymous:1964:EE

[Ano64d] Anonymous. Errata to errata. The Computer Journal, 7(2):162, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/2/162.full.pdf+html.

Anonymous:1964:O

[Ano64e] Anonymous. Obituary. The Computer Journal, 7(2):123, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
2/123.full.pdf+html.

Anonymous:1965:DTH

[Ano65a] Anonymous. Data transmission handbook. The Computer Journal, 7(4):302, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/4/302.1.full.pdf+html.

Anonymous:1965:EB

[Ano65b] Anonymous. Editorial board. The Computer Journal, 7(4):307, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/4/307.full.pdf+html.

Anonymous:1965:EBC

[Ano65c] Anonymous. Editorial board — corrigendum. The Computer Journal, 8(2):94, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/2/94.full.pdf+html.

Anonymous:1965:E

[Ano65d] Anonymous. Errata. The Computer Journal, 8(1):27, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
1/27.full.pdf+html.

Anonymous:1965:IA

[Ano65e] Anonymous. Index to advertisers. The Computer Journal, 7(4): xvi, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/4/xvi.full.pdf+html.

Anonymous:1965:CDP

[Ano65f] Anonymous. The Computing and Data Processing Society of Canada. The Computer Journal, 7(4):302, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
4/302.2.full.pdf+html.

Anonymous:1966:AS

[Ano66a] Anonymous. Algorithms supplement. The Computer Journal, 9(3):321–324, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/3/321.full.pdf+html.

Anonymous:1966:C

[Ano66b] Anonymous. Correspondence. The Computer Journal, 9(1):115a– 115, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/1/115-a.full.pdf+html.

Anonymous:1966:EB

[Ano66c] Anonymous. Editorial board. The Computer Journal, 9(3):xiv, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/xiv.full.pdf+html.

Anonymous:1966:E

[Ano66d] Anonymous. Errata. The Computer Journal, 9(1):116, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
1/116.full.pdf+html.

Anonymous:1966:ECL

[Ano66e] Anonymous. Error curves for Lanczos selected points method. The Computer Journal, 9(1):115b–115, May 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/9/1/115-b.full.pdf+
html.

Anonymous:1966:IA

[Ano66f] Anonymous. Index to advertisers. The Computer Journal, 9(1): xviii, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/1/xviii.full.pdf+html.

Anonymous:1966:NPT

[Ano66g] Anonymous. Nonlinear programming test problems. The Computer Journal, 8(4):391, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/4/391.full.pdf+html.

Anonymous:1966:NSP

[Ano66h] Anonymous. Notes on the submission of papers. The Computer Journal, 9(1):xx, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/1/xx.full.pdf+html.

Anonymous:1966:NFP

[Ano66i] Anonymous. Notice: Future publication of Algorithms Supplement. The Computer Journal, 9(1):80, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/80.full.pdf+html.

Anonymous:1966:S

[Ano66j] Anonymous. “systematic”. The Computer Journal, 9(3):229, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/229.full.pdf+html.

Anonymous:1967:ASa

[Ano67a] Anonymous. Algorithms supplement. The Computer Journal, 10(1):112–115, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/1/112.full.pdf+html.

Anonymous:1967:ASb

[Ano67b] Anonymous. Algorithms supplement. The Computer Journal, 10(2):206–210, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/2/206.full.pdf+html.

Anonymous:1967:ASc

[Ano67c] Anonymous. Algorithms supplement. The Computer Journal, 10(3):306–312, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/3/306.full.pdf+html.

Anonymous:1967:AAC

[Ano67d] Anonymous. An audit approach to computers. The Computer Journal, 10(2):134, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/2/134.full.pdf+html.

Anonymous:1967:BCJ

[Ano67e] Anonymous. The Computer Journal. The Computer Journal, 9(4):419–420, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/4/419.full.pdf+html.

Anonymous:1967:C

[Ano67f] Anonymous. Corrigendum. The Computer Journal, 10(3):314, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/314.full.pdf+html.

Anonymous:1967:DCO

[Ano67g] Anonymous. Discussion and correspondence: One-day symposium on PL/1 held at National Physical Laboratory, 18 May 1967. The Computer Journal, 10(2):211–215, August 1967.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
2/211.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/211.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/212.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/213.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/214.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_02/tiff/215.tif.

Anonymous:1967:EBa

[Ano67h] Anonymous. Editorial board. The Computer Journal, 10(2):125, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/2/125.full.pdf+html.

Anonymous:1967:Ea

[Ano67i] Anonymous. Errata. The Computer Journal, 9(4):372, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
4/372.2.full.pdf+html.

Anonymous:1967:Eb

[Ano67j] Anonymous. Erratum. The Computer Journal, 10(1):84, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/84.full.pdf+html.

Anonymous:1967:LE

[Ano67k] Anonymous. Letters to the editor. The Computer Journal, 10(2): 216–217, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/216.full.pdf+html.

Anonymous:1967:MSB

[Ano67l] Anonymous. The memory system of the brain. The Computer Journal, 10(2):187, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/2/187.full.pdf+html.

Anonymous:1967:NSP

[Ano67m] Anonymous. Notes on the submission of papers. The Computer Journal, 10(2):218, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/2/218.full.pdf+html.

Anonymous:1968:ASa

[Ano68a] Anonymous. Algorithms supplement. The Computer Journal, 10(4):414–416, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/4/414.full.pdf+html.

Anonymous:1968:ASb

[Ano68b] Anonymous. Algorithms supplement. The Computer Journal, 11(1):112–115, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/1/112.full.pdf+html.

Anonymous:1968:ASc

[Ano68c] Anonymous. Algorithms supplement. The Computer Journal, 11(2):231–235, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/2/231.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/231.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/232.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/233.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/234.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/235.tif.

Anonymous:1968:ASd

[Ano68d] Anonymous. Algorithms supplement. The Computer Journal, 11(3):347–350, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/3/347.full.pdf+html.

Anonymous:1968:EBa

[Ano68e] Anonymous. Editorial board. The Computer Journal, 11(2):127, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/2/127.full.pdf+html.

Anonymous:1968:Ea

[Ano68f] Anonymous. Errata. The Computer Journal, 10(4):405, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
4/405.full.pdf+html.

Anonymous:1968:Eb

[Ano68g] Anonymous. Errata. The Computer Journal, 11(1):25, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/25.full.pdf+html.

Anonymous:1969:ASa

[Ano69a] Anonymous. Algorithms supplement. The Computer Journal, 12(1):100–103, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/1/100.full.pdf+html.

Anonymous:1969:ASb

[Ano69b] Anonymous. Algorithms supplement. The Computer Journal, 12(2):197–199, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/2/197.full.pdf+html.

Anonymous:1969:ASc

[Ano69c] Anonymous. Algorithms supplement. The Computer Journal, 12(3):291–295, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/3/291.full.pdf+html.

Anonymous:1969:ASd

[Ano69d] Anonymous. Algorithms supplement. The Computer Journal, 12(4):405–409, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/4/405.full.pdf+html.

Anonymous:1969:AHA

[Ano69e] Anonymous. Alston S. Householder Award. The Computer Journal, 12(4):392, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/4/392.full.pdf+html.

Anonymous:1969:BCJ

[Ano69f] Anonymous. The Computer Journal — Honorary Editors. The Computer Journal, 12(4):364, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/364.full.
pdf+html.

Anonymous:1969:EBa

[Ano69g] Anonymous. Editorial board. The Computer Journal, 12(1):87, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/87.full.pdf+html.

Anonymous:1969:EBb

[Ano69h] Anonymous. Editorial board. The Computer Journal, 12(3):viii, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/3/viii.full.pdf+html.

Anonymous:1969:FPa

[Ano69i] Anonymous. Future papers. The Computer Journal, 12(3):295, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/3/295.full.pdf+html.

Anonymous:1969:FPb

[Ano69j] Anonymous. Future papers. The Computer Journal, 12(4):404, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/404.full.pdf+html.

Anonymous:1969:HEC

[Ano69k] Anonymous. Honorary Editor — The Computer Journal. The Computer Journal, 12(2):170, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/2/170.full.pdf+html.

Anonymous:1969:LEa

[Ano69l] Anonymous. Letter to the Editor. The Computer Journal, 12 (1):104–105, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/1/104.2.full.pdf+html.

Anonymous:1969:LEb

[Ano69m] Anonymous. Letters to the Editor. The Computer Journal, 12(3): 298–300, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/3/298.full.pdf+html.

Anonymous:1969:NMB

[Ano69n] Anonymous. Notice to members of the British computer society. The Computer Journal, 12(3):287, August 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/12/3/287.
full.pdf+html.

ApSimon:1965:IP

[ApS65] H. G. ApSimon. “An impossible program”. The Computer Journal, 8(1):72, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/72.full.pdf+html. Rebuttal to [Str65e].

Arscott:1965:BR

[Ars65] F. M. Arscott. Book review. The Computer Journal, 8(3):189, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
3/189.2.full.pdf+html.

Anstey:1963:URD

[AS63] T. H. Anstey and K. W. Smillie. Use of a remote digital computer on an open-shop basis in agricultural research. The Computer Journal, 6(2):118–120, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/118.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060118.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
118.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/119.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/120.tif.

Ashworth:1967:BR

[Ash67a] W. Ashworth. Book review. The Computer Journal, 10(1):28, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/28.full.pdf+html.

Ashworth:1967:LPA

[Ash67b] Wilfred Ashworth. Library planning for automation. The Computer Journal, 10(2):197, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/2/197.full.pdf+html.

Ashby:1968:MPA

[Ash68] T. Ashby. A modification to Paulson’s approximation to the variance ratio distribution. The Computer Journal, 11(2):209– 210, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/209.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/110209.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/209.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/210.tif.

Allen:1970:CWA

[AT70] R. L. Allen and The Editor. Correspondence: What is an analyst? The Computer Journal, 13(1):119–120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/119.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/120.

tif. See [Fel69, Fel70].

Ayre:1964:UER

[Ayr64] T. H. Ayre. User’s experience of rapidwrite. The Computer Journal, 7(2):101–102, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/101.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
02/070101.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/101.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/
tiff/102.tif.

Baer:1962:NEL

[Bae62] Robert M. Baer. Note on an extremum locating algorithm. The Computer Journal, 5(3):193, November 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/193.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/050193.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/
193.tif.

Baecker:1964:BRb

[Bae64a] H. D. Baecker. Book review. The Computer Journal, 7(3):184, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
3/184.full.pdf+html.

Baecker:1964:BRa

[Bae64b] H. D. Baecker. Book reviews. The Computer Journal, 6(4): 307, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/4/307.full.pdf+html.

Baecker:1964:IUR

[Bae64c] H. D. Baecker. Initial users’ reactions: what do they really want? The Computer Journal, 7(2):108–109, July 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/108.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070108.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
108.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/109.tif.

Baecker:1966:ATS

[Bae66a] H. D. Baecker. Analysis of a time-shared computer system. The Computer Journal, 9(2):128, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/128.full.
pdf+html.

Baecker:1966:BRB

[Bae66b] H. D. Baecker. Book review: Computer Augmentation of Human Resourcing. The Computer Journal, 9(1):20, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
1/20.full.pdf+html.

Baecker:1966:CAB

[Bae66c] H. D. Baecker. Correspondence: On “An analysis, both theoretical and by simulation, of a time-shared computer system”. The Computer Journal, 9(2):128, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
090128.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/128.tif. See [Pen66].

Baecker:1967:BR

[Bae67] H. D. Baecker. Book review. The Computer Journal, 9(4):352, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/352.full.pdf+html.

Bagley:1961:IPO

[Bag61a] P. R. Bagley. Improving problem-oriented language by stratifying it. The Computer Journal, 4(3):217–221, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/3/217.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/040217.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/
217.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/218.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_03/tiff/219.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/221.tif.

Bagshaw:1961:OCS

[Bag61b] A. R. Bagshaw. Organization of a computing service for industry and commerce. The Computer Journal, 4(2):181–184, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/2/181.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_02/040181.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/181.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/182.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/183.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/184.tif.

Bagley:1962:PPU

[Bag62] P. R. Bagley. Principles and problems of a Universal Computer-Oriented Language. The Computer Journal, 4(4):305–312, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/4/305.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_04/040305.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/305.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/306.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/307.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/308.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/309.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/310.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/311.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/312.tif.

Bakes:1964:AMS

[Bak64] M. D. Bakes. An alternative method of solution of certain tridiagonal systems of linear equations. The Computer Journal, 7(2): 135–136, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/2/135.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/070135.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/135.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/136.tif.

Barrett:1961:CPG

[Bar61] W. Barrett. Convergence properties of Gaussian quadrature formulae. The Computer Journal, 3(4):272–277, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/272.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030272.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
272.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/273.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/274.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/

Issue_04/tiff/275.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/276.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/277.tif.

Barron:1962:BR

[Bar62] D. W. Barron. Book review. The Computer Journal, 4(4):331, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
4/331.full.pdf+html.

Barron:1964:CRA

[Bar64] D. W. Barron. Correspondence: On the representation of AL-GOL. The Computer Journal, 6(4):335, January 1964. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/335.tif. See [Sam63c, Tay64].

Barnes:1965:ASN

[Bar65a] J. G. P. Barnes. An algorithm for solving non-linear equations based on the secant method. The Computer Journal, 8(1):66– 72, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/66.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/080066.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/66.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/67.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/68.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/70.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/71.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/72.tif.

Barnes:1965:KAL

[Bar65b] J. G. P. Barnes. A KDF9 Algol list-processing scheme. The Computer Journal, 8(2):113–119, July 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/2/113.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/080113.sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/
113.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/114.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/tiff/115.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/117.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/119.tif.

Barraclough:1965:ADC

[Bar65c] E. D. Barraclough. The application of a digital computer to the construction of timetables. The Computer Journal, 8(2):136–146, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/2/136.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/080136.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/136.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/137.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/139.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/140.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/141.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/142.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/143.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/144.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/146.tif.

Barron:1965:LEI

[Bar65d] D. W. Barron. Letter to the Editor: The ISO character code. The Computer Journal, 7(4):281, January 1965. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/7/4/281.
full.pdf+html.

Barron:1966:CES

[Bar66a] D. W. Barron. Correspondence: On “The Egdon system for the KDF9”. The Computer Journal, 9(3):247, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090247.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/

247.tif. See [BHJV66] and correspondence [Sam66a].

Barron:1966:P

[Bar66b] D. W. Barron. “progress”. The Computer Journal, 9(3):247, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/247.full.pdf+html.

Barton:1967:SMA

[Bar67] D. Barton. A scheme for manipulative algebra on a computer. The Computer Journal, 9(4):340–344, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/340.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090340.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
340.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/341.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/342.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/343.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/344.tif.

Barron:1969:DCN

[Bar69a] D. W. Barron. Discussion and correspondence: a note on program debugging in an on-line environment. The Computer Journal, 12(1):104, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_12/Issue_01/120104.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/104.tif.

Barron:1969:NPD

[Bar69b] D. W. Barron. A note on program debugging in an on-line environment. The Computer Journal, 12(1):104, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
1/104.1.full.pdf+html.

Bauldreay:1965:C

[Bau65a] J. Bauldreay. Correspondence. The Computer Journal, 8(3):207, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/3/207.full.pdf+html.

Bauldreay:1965:CCR

[Bau65b] J. Bauldreay. Correspondence: character recognition. The Computer Journal, 8(3):207, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/207.
tif.

Box:1969:CVR

[BB69] M. J. Box and R. M. Box. Computation of the variance ratio distribution. The Computer Journal, 12(3):277–278, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
3/277.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/120277.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/277.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/278.tif.

Barton:1968:SAS

[BBB68] D. Barton, S. R. Bourne, and C. J. Burgess. A simple algebra system. The Computer Journal, 11(3):293–298, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/293.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/110293.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/tiff/293.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/294.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/295.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/296.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/297.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/298.tif.

Burbridge:1966:CS

[BBdD66] D. A. Burbridge, A. Robert Brown, A. d’Agapeyeff, and F. G. Duncan. Correspondence: Systematics. The Computer Journal, 9(3):229–229, 237–237, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/229.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/237.tif. See [Gri66].

Backus:1963:RRA

[BBG^{+}63] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Revised report on the algorithmic language ALGOL 60. The Computer Journal, 5(4):349–367, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/349.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050349.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
349.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/350.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/351.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/352.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/353.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/355.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/356.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/357.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/358.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/359.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/360.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/361.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/362.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/363.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/364.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/365.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/367.tif.

Barron:1963:MFC

[BBH^{+}63] D. W. Barron, J. N. Buxton, D. F. Hartley, E. Nixon, and C. Strachey. The main features of CPL. The Computer Journal, 6(2):134–143, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/134.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060134.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
134.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/135.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/136.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_02/tiff/137.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/138.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/139.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/140.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/141.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/142.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/143.tif.

Baum:1966:UDA

[BBS66] M. M. Baum, R. G. Blake, and R. J. Smale. Use of digital analogue simulator (DAS). The Computer Journal, 9(2):175–180, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/2/175.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_02/090175.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/176.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/177.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/178.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/179.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/180.tif.

Brooker:1967:IDC

[BC67] R. A. Brooker and S. R. Clark. An index directed compiler. The Computer Journal, 10(1):60–63, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/60.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100060.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
60.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/62.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
63.tif.

Bennett:1961:CAC

[BD61] J. M. Bennett and R. J. Dakin. Computers as an aid in computer design assessment. The Computer Journal, 3(4):253–256, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/4/253.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_04/030253.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/253.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/254.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/255.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/256.tif.

Beasley:1965:NAC

[Bea65] J. D. Beasley. A note on the arrangement of Chebyshev series. The Computer Journal, 8(3):278–279, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/278.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080278.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
278.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/279.tif.

Beauchamp:1968:SWF

[Bea68] K. G. Beauchamp. The simulation of wave filters having polynomial transfer functions on an analogue computer. The Computer Journal, 10(4):352–359, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/352.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100352.sgm.abs.html.

Bell:1963:AFT

[Bel63] D. A. Bell. Approximations in Fourier transforms. The Computer Journal, 6(3):244–247, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/244.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060244.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
244.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/245.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/246.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/247.tif.

Bell:1967:BR

[Bel67a] D. A. Bell. Book review. The Computer Journal, 10(3):281, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/281.full.pdf+html.

Bell:1967:AT

[Bel67b] M. Bell. Automata theory. The Computer Journal, 10(2):171, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/2/171.full.pdf+html.

Bell:1968:BR

[Bel68] Michael Bell. Book review. The Computer Journal, 10(4):391, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/391.full.pdf+html.

Bell:1969:BRa

[Bel69a] D. A. Bell. Book review. The Computer Journal, 12(2):182, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/182.full.pdf+html.

Bell:1969:BRb

[Bel69b] D. A. Bell. Book review. The Computer Journal, 12(4):376, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/376.full.pdf+html.

Bell:1969:MTS

[Bel69c] Michael Bell. Mathematical theory of switching circuits and automata. The Computer Journal, 12(3):220, August 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/12/3/220.
full.pdf+html.

Bemer:1969:C

[Bem69] R. W. Bemer. Correspondence. The Computer Journal, 12(4): 412, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/412.1.full.pdf+html.

Berlin:1961:C

[Ber61] M. M. Berlin. Correspondence. The Computer Journal, 3(4): 210, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/4/210.full.pdf+html.

Berry:1966:GP

[Ber66] R. E. Berry. Grader programs. The Computer Journal, 9(3):252– 256, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/252.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/090252.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/252.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/253.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/255.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/256.tif.

Bergson:1969:CNC

[Ber69] A. Bergson. Correspondence: Note on the calculation of e to many significant digits. The Computer Journal, 12(1):105, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/105.tif.

Braunholtz:1961:NPL

[BFH61] T. G. H. Braunholtz, A. G. Fraser, and P. M. Hunt. NEB-ULA: a programming language for data processing. The Computer Journal, 4(3):197–211, October 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/3/197.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/040197.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/
197.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/198.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_03/tiff/199.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_03/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/201.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/203.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/204.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/205.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/206.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/207.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/208.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/209.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/210.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/211.tif.

Baker:1964:NSF

[BFMW64] C. T. H. Baker, L. Fox, D. F. Mayers, and K. Wright. Numerical solution of Fredholm integral equations of first kind. The Computer Journal, 7(2):141–148, July 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/141.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070141.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
141.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_07/Issue_02/tiff/142.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/143.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/tiff/144.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/145.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/
tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/147.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/148.
tif.

Barron:1963:TPE

[BH63] D. W. Barron and D. F. Hartley. Techniques for program error diagnosis on EDSAC 2. The Computer Journal, 6(1):44– 49, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/44.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060044.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/46.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/48.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/49.tif.

Barnard:1968:URD

[BH68] D. F. Barnard and D. F. Holman. The use of Roth’s decomposition algorithm in multi-level design of circuits. The Computer Journal, 11(3):269–276, November 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/269.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110269.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
269.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/270.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/271.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/272.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/273.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/

tiff/274.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/275.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/276.
tif.

Bruyendonck:1969:C

[BH69a] D. Bruyendonck and D. Hirschberg. Correspondence. The Computer Journal, 12(2):200, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/2/200.1.full.pdf+html.

Bruyendonck:1969:CCD

[BH69b] D. Bruyendonck and D. Hirschberg. Correspondence: Convex differentiable curves. The Computer Journal, 12(2):200, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/200.tif. See [Phi68a].

Burns:1966:ESK

[BHJV66] D. Burns, E. N. Hawkins, D. R. Judd, and J. L. Venn. The Egdon system for the KDF9. The Computer Journal, 8(4):297–302, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/297.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080297.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/297.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/298.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/299.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/300.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/301.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

08/Issue_04/tiff/302.tif. See correspondence [Bar66a, Sam66a].

Buckle:1969:CSF

[BHRS69] D. G. R. Buckle, C. H. Hunt, Maureen M. Ryder, and S. H. Storey. Computers and the small firm: 2. The Computer Journal, 12(1):10–14, February 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/10.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120010.sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
10.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/12.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
13.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/14.tif.

Beale:1965:EUD

[BHS65] E. M. L. Beale, P. A. B. Hughes, and R. E. Small. Experiences in using a decomposition program. The Computer Journal, 8(1):13–18, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/8/1/13.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_01/080013.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/13.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/14.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/15.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/16.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/17.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/18.tif.

Bickley:1968:PCI

[Bic68] W. G. Bickley. Piecewise cubic interpolation and two-point boundary problems. The Computer Journal, 11(2):206–208, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/206.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110206.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/206.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/207.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/208.
tif.

Bickley:1969:CPC

[Bic69] W. G. Bickley. Correspondence: Piecewise cubic interpolation and two-point boundary problems. The Computer Journal, 12(1): 105, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/105.tif.

Birtwistle:1968:ESE

[Bir68] G. M. Birtwistle. The explicit solution of the equation of heat conduction. The Computer Journal, 11(3):317–323, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/317.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/110317.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/tiff/317.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/318.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/319.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/320.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/321.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/322.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/323.tif.

Bishop:1966:ACM

[Bis66] A. H. J. Bishop. Analogue computing methods. The Computer Journal, 9(2):180, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/2/180.full.pdf+html.

Buxton:1962:CSL

[BL62] J. N. Buxton and J. G. Laski. Control and simulation language. The Computer Journal, 5(3):194–199, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/194.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/050194.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/
194.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/195.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_03/tiff/196.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_03/tiff/197.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/198.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/199.tif.

Brooker:1960:APP

[BM60] R. A. Brooker and D. Morris. An assembly program for a phrase structure language. The Computer Journal, 3(3):168–174, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/3/168.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_03/030168.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_03/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/169.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/
tiff/170.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_03/tiff/171.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/172.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_03/tiff/173.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/174.tif.

Brooker:1961:C

[BM61a] R. A. Brooker and D. Morris. Correspondence. The Computer Journal, 4(2):176, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/4/2/176.full.pdf+html.

Brooker:1961:SPR

[BM61b] R. A. Brooker and D. Morris. Some proposals for the realization of a certain assembly program. The Computer Journal, 3(4):220–231, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/4/220.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_04/030220.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/221.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/222.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/223.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/224.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/225.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/226.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/227.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/228.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/229.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/231.tif.

Barrett:1963:EAP

[BM63] W. Barrett and A. J. Mitchell. An extended Autocode for Pegasus. The Computer Journal, 6(3):237–240, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/237.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060237.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
237.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/238.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/239.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/240.tif.

Brooker:1962:TR

[BMR62] R. A. Brooker, D. Morris, and J. S. Rohl. Trees and routines. The Computer Journal, 5(1):33–47, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/5/1/33.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_01/050033.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/33.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
01/tiff/34.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/35.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/36.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/37.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/38.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
39.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/40.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/41.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
42.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/43.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/44.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
45.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

05/Issue_01/tiff/46.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/47.tif.

Brooker:1967:CCF

[BMR67a] R. A. Brooker, D. Morris, and J. S. Rohl. Compiler compiler facilities in Atlas Autocode. The Computer Journal, 9(4):350–352, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/350.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090350.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/350.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/351.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/352.tif.

Brooker:1967:ECC

[BMR67b] R. A. Brooker, D. Morris, and J. S. Rohl. Experience with the compiler compiler. The Computer Journal, 9(4):345–349, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/345.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090345.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/345.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/346.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/347.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/348.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/349.
tif.

Bohte:1968:NSI

[Boh68] Z. Bohte. Numerical solution of the inverse algebraic eigenvalue problem. The Computer Journal, 10(4):385–388, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
4/385.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/100385.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/tiff/385.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/386.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/387.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/388.tif.

Bonney:1969:SCC

[Bon69] M. C. Bonney. Some considerations of the cost and value of information. The Computer Journal, 12(2):118–123, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/2/118.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/120118.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/
118.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/119.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/120.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/121.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/122.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/123.tif.

Boothroyd:1967:AS

[Boo67a] J. Boothroyd. Algorithm 18: SUMFAC. The Computer Journal, 9 (4):416, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/416.tif.

Boothroyd:1967:APR

[Boo67b] J. Boothroyd. Algorithm 20: Permutations of the rows or columns of a matrix. The Computer Journal, 10(2):206–207, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/206.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/207.
tif.

Boothroyd:1967:ASPa

[Boo67c] J. Boothroyd. Algorithm 22: Shortest path between start node and end node of a network. The Computer Journal, 10(3):306– 307, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/306.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/307.tif.

Boothroyd:1967:ASPb

[Boo67d] J. Boothroyd. Algorithm 23: Shortest path between start node and all other nodes of a network. The Computer Journal, 10(3):307– 308, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/307.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/308.tif.

Boothroyd:1967:ALN

[Boo67e] J. Boothroyd. Algorithm 24: The list of nodes on the shortest path from start node to end node of a network. The Computer Journal, 10(3):308, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/308.

tif. See note [Ham69d].

Boothroyd:1967:ASS

[Boo67f] J. Boothroyd. Algorithm 25: Sort a section of the elements of an array by determining the rank of each element. The Computer Journal, 10(3):308–309, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
308.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/309.tif. See note [Hil69, Woo70].

Boothroyd:1967:AOS

[Boo67g] J. Boothroyd. Algorithm 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 10(3):309–310, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
309.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/310.tif. See note [Hil69, Woo70].

Boothroyd:1967:ARE

[Boo67h] J. Boothroyd. Algorithm 27: Rearrange the elements of an array section according to a permutation of the subscripts. The Computer Journal, 10(3):310, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/310.
tif.

Boothroyd:1967:APEa

[Boo67i] J. Boothroyd. Algorithm 28: Permutations of the elements of a vector in lexicographic order. The Computer Journal, 10(3): 311, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/311.tif.

Boothroyd:1967:APEb

[Boo67j] J. Boothroyd. Algorithm 29: Permutation of the elements of a vector. The Computer Journal, 10(3):311, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/311.tif.

Boothroyd:1967:AFP

[Boo67k] J. Boothroyd. Algorithm 30: Fast permutation of the elements of a vector. The Computer Journal, 10(3):311–312, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/311.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/312.
tif.

Boothroyd:1967:AANa

[Boo67l] J. Boothroyd. Algorithms: Author’s note on algorithms 22, 23, 24. The Computer Journal, 10(3):306, November 1967. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/306.tif.

Boothroyd:1967:AANb

[Boo67m] J. Boothroyd. Algorithms: Author’s note on algorithms 25, 26, 27. The Computer Journal, 10(3):308, November 1967. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/308.tif.

Boothroyd:1967:AANc

[Boo67n] J. Boothroyd. Algorithms: Author’s note on algorithms 28, 29, 30. The Computer Journal, 10(3):310, November 1967. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/310.tif.

Boothroyd:1968:ACF

[Boo68a] J. Boothroyd. Algorithm 31: Complex Fourier series. The Computer Journal, 10(4):414–416, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/414.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/415.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/416.
tif.

Boothroyd:1968:SNA

[Boo68b] J. Boothroyd. Special note on Algorithm 31. complex Fourier analysis. The Computer Journal, 11(1):115, May 1968. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/115.tif.

Boothroyd:1979:CAF

[Boo79] J. Boothroyd. Certification of Algorithm 2: Fibonacci search. The Computer Journal, 22(1):86, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/

tiff/86.tif. Reprinted from The Computer Bulletin. See [PP79].

Boutel:1965:C

[Bou65] B. E. Boutel. Correspondence. The Computer Journal, 8(3): 215, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/3/215.1.full.pdf+html.

Box:1965:NMC

[Box65] M. J. Box. A new method of constrained optimization and a comparison with other methods. The Computer Journal, 8(1): 42–52, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/42.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/080042.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/45.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/46.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
49.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/51.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
52.tif.

Box:1966:CSC

[Box66a] M. J. Box. A comparison of several current optimization methods, and the use of transformations in constrained problems. The Computer Journal, 9(1):67–77, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/67.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090067.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/67.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/68.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/70.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/71.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/72.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
73.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/74.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/75.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
76.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/77.tif.

Box:1966:CNP

[Box66b] M. J. Box. Correspondence: On “Nonlinear programming test problems”. The Computer Journal, 9(1):115–116, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/115.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/116.

tif. See [Myl66].

Box:1966:NPT

[Box66c] M. J. Box. Nonlinear programming test problems. The Computer Journal, 9(1):115c–116, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/1/115-c.full.pdf+html.

Box:1967:BR

[Box67] M. J. Box. Book review. The Computer Journal, 10(1):73, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/73.full.pdf+html.

Boyd:1960:SEL

[Boy60] K. T. Boyd. Simultaneous equations and linear programming. The Computer Journal, 3(1):45–46, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/1/45.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/030045.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/46.tif.

Boyell:1963:MSG

[Boy63] R. L. Boyell. The method of successive grids for reduction of function storage requirements. The Computer Journal, 5(4):320–321, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/4/320.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/050320.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/320.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/321.tif.

Boyles:1968:O

[Boy68] E. E. Boyles. Obituary. The Computer Journal, 11(1):16, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/16.full.pdf+html.

Byrne:1969:IGI

[BP69] J. L. Byrne and L. G. Proll. Initialising Geoffrion’s implicit enumeration algorithm for the zero-one linear programming problem.

The Computer Journal, 12(4):381–384, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/381.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120381.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
381.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/382.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/383.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/384.tif.

Brace:1963:BR

[Bra63a] D. A. Brace. Book reviews. The Computer Journal, 6(3):276, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/276.1.full.pdf+html.

Brace:1963:DCE

[Bra63b] D. A. Brace. Direct coding of English language names. The Computer Journal, 6(2):113–117, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/113.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060113.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
113.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/114.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/115.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_02/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/117.tif. See corre-

spondence [Ker63b].

Brooker:1966:MFA

[BRC66] R. A. Brooker, J. S. Rohl, and S. R. Clark. The main features of Atlas Autocode. The Computer Journal, 8(4):303–310, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/303.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080303.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_08/Issue_04/tiff/304.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/305.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/306.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/307.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/308.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/309.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/

tiff/310.tif. See comment [CL66a].

Bridger:1965:BR

[Bri65] M. Bridger. Book review. The Computer Journal, 8(3):245, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
3/245.full.pdf+html.

Briggs:1966:BRB

[Bri66] P. A. N. Briggs. Book review: Optimization Theory. The Computer Journal, 8(4):397, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/4/397.full.pdf+html.

Brittan:1967:NSM

[Bri67] J. N. G. Brittan. The nature and structure of man-computer communication in team control systems. The Computer Journal, 10(3):219–226, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/3/219.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
03/100219.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/219.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/220.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/221.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/222.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/223.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/224.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/226.tif.

Brooker:1960:STD

[Bro60] R. A. Brooker. Some techniques for dealing with two-level storage. The Computer Journal, 2(4):189–194, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/4/189.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/020189.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/
189.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/tiff/190.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_04/tiff/191.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/
Issue_04/tiff/192.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_04/tiff/193.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/
tiff/194.tif.

Brookes:1968:BR

[Bro68] B. C. Brookes. Book review. The Computer Journal, 11(3):263, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/3/263.full.pdf+html.

Brown:1969:STT

[Bro69a] A. R. Brown. Selling television time: an optimisation problem. The Computer Journal, 12(3):201–207, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/3/201.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/120201.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/
201.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/202.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/203.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/204.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/205.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/206.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/207.tif.

Brown:1969:UMP

[Bro69b] P. J. Brown. Using a macro processor to aid software implementation. The Computer Journal, 12(4):327–331, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/327.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/120328.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/tiff/328.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/329.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/330.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_04/tiff/331.tif.

Broyden:1969:ASN

[Bro69c] C. G. Broyden. Algorithm 44: Solution of nonlinear simultaneous equations. The Computer Journal, 12(4):406–408, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/406.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/407.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/408.tif.  See note [Fie70,

Red71].

Broyden:1969:NMS

[Bro69d] C. G. Broyden. A new method of solving nonlinear simultaneous equations. The Computer Journal, 12(1):94–99, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/94.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120094.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/94.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/95.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
96.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/97.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/98.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
99.tif.

Bennett:1961:ATC

[BS61] R. O. Bennett and J. B. Stringer. Acceptance trials of computer systems for government use. The Computer Journal, 4(3):185–196, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/3/185.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_03/040185.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/185.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/186.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/187.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/188.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/189.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/190.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/191.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/192.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/193.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/194.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/195.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/196.tif.

Barron:1960:SSL

[BSD60] D. W. Barron and H. P. F. Swinnerton-Dyer. Solution of simultaneous linear equations using a magnetic-tape store. The Computer Journal, 3(1):28–33, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/1/28.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/030028.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/29.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/31.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/32.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_01/tiff/33.tif.

Boutel:1965:CIP

[BSPI65] B. E. Boutel, C. Strachey, J. H. G. Phillips, and Michael Irish. Correspondence: An impossible program. The Computer Journal, 8 (3):215, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/215.tif. See [Str65e,

ApS65, HLOS65, AMKM66].

Brace:1963:BRB

[BT63] D. A. Brace and G. J. Tee. Book reviews: Elementary Mathematical Programming, by Robert W. Metzger, 1963; U.S.S.R. Computational Mathematics and Mathematical Physics, Number 1, 1962. The Computer Journal, 6(3):276–276, 286–286, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/276.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/286.
tif.

Buckingham:1960:OUC

[Buc60] R. A. Buckingham. The organization of a university computer centre. The Computer Journal, 3(3):131–135, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/131.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030131.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
131.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/132.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_03/tiff/133.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_03/tiff/134.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/135.tif.

Buckingham:1965:CU

[Buc65] R. A. Buckingham. The computer in the university. The Computer Journal, 8(1):1–7, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/8/1/1.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_01/080001.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/1.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/2.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/3.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/tiff/4.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/

Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
7.tif.

Burstall:1966:BRB

[Bur66a] R. M. Burstall. Book review: Computer and Information Sciences. The Computer Journal, 9(1):15, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/15.full.pdf+html.

Burstall:1966:CDE

[Bur66b] R. M. Burstall. Computer design of electricity supply networks by a heuristic method. The Computer Journal, 9(3):263–274, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/263.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090263.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/263.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/264.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/265.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/266.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/267.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/268.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/269.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/270.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/271.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/272.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/273.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/274.tif.

Burley:1967:PLL

[Bur67] H. T. Burley. A programming language for linear algebra. The Computer Journal, 10(1):69–73, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/69.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100069.sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
69.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/70.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/71.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
72.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/73.tif.

Burstall:1968:ATL

[Bur68] R. M. Burstall. Automation theory and learning systems. The Computer Journal, 11(2):168, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/2/168.full.pdf+html.

Burstall:1969:PSW

[Bur69a] R. M. Burstall. A program for solving word sum puzzles. The Computer Journal, 12(1):48–51, February 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/48.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120048.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
48.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/49.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/50.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
51.tif.

Burstall:1969:PPP

[Bur69b] Rod M. Burstall. Proving properties of programs by structural induction. The Computer Journal, 12(1):41–48, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/41.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120041.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/41.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/42.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
43.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
46.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

12/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/48.tif.

Butcher:1961:RSN

[But61] J. C. Butcher. Random sampling from the normal distribution. The Computer Journal, 3(4):251–253, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/251.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030251.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
251.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/252.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/253.tif.

Buxton:1963:C

[Bux63] J. M. R. Buxton. Correspondence. The Computer Journal, 6 (1):56, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/56.2.full.pdf+html.

Buxton:1966:WSC

[Bux66] J. N. Buxton. Writing simulations in CSL. The Computer Journal, 9(2):137–143, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/137.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090137.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
137.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/138.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/139.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/140.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/141.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/142.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/143.tif.

Beasley:1969:DTS

[BW69a] J. D. Beasley and K. Wilson. Design and testing of the System 4 random number generator. The Computer Journal, 12(4):368– 372, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/368.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/120368.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/368.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/369.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/370.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/371.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/372.tif.

Bellamy:1969:MPO

[BW69b] N. W. Bellamy and M. J. West. Methods of profile optimisation by iterative analogue computation. The Computer Journal, 12(2): 132–138, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/132.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120132.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/132.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/133.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/134.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/135.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/136.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/137.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/138.
tif.

Barrodale:1966:NNP

[BY66] Ian Barrodale and Andrew Young. A note on numerical procedures for approximation by spline functions. The Computer Journal, 9(3):318–320, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/318.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090318.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
318.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/319.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/320.tif.

Bye:1965:OSN

[Bye65] John A. T. Bye. Obtaining solutions of the Navier–Stokes equation by relaxation processes. The Computer Journal, 8(1):53– 56, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/53.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/080053.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/55.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/56.tif.

Bye:1967:NMD

[Bye67] J. A. T. Bye. A new method of determining eigenvalues and eigenfunctions. The Computer Journal, 10(1):108–111, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/108.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/100108.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
01/tiff/108.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/109.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/110.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/111.tif.

Cadwell:1961:LSS

[Cad61] J. H. Cadwell. A least squares surface fitting program. The Computer Journal, 3(4):266–269, January 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/266.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030266.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
266.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/267.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/268.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/269.tif.

Cameron:1969:LCP

[Cam69] P. T. Cameron. A language for chemical plant design and simulation programs. The Computer Journal, 12(1):24–28, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/24.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120024.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/25.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
26.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/28.tif.

Carre:1961:DOA

[Car61] B. A. Carré. The determination of the optimum accelerating factor for successive over-relaxation. The Computer Journal, 4(1): 73–78, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/73.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/040073.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/73.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/74.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/75.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/76.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/77.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/78.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/79.tif.

Carling:1963:BRB

[Car63] L. N. Carling. Book review — Analogue computers. The Computer Journal, 6(1):111–112, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/6/1/111.full.pdf+html.

Carling:1964:SSI

[Car64] L. N. Carling. A study of the solution of an initial-value problem with a hybrid computer. The Computer Journal, 7(1):40– 46, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/1/40.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_01/070040.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/40.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/41.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/42.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/43.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/44.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/45.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/46.tif.

Carre:1966:PNE

[Car66] B. A. Carré. The partitioning of network equations for block iteration. The Computer Journal, 9(1):84–97, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/1/84.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/090084.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/84.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/85.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
86.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/87.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/88.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
89.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/90.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/91.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
92.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/93.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/94.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
95.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/96.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/97.tif.

Carrick:1970:CIL

[Car70] D. A. Carrick. Correspondence: Interpretation of limited entry decision table format. The Computer Journal, 13(2):220–221, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/2/220.2.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/220.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/
Issue_02/tiff/221.tif. See [Kin69a, Kin71].

Caskey:1963:ASP

[Cas63] W. S. Caskey. Accounting for the soldier’s pay — organization of programming. The Computer Journal, 5(4):258–263, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/4/258.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/050258.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/258.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/259.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/260.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/261.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/262.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/263.tif.

Cayless:1961:SSO

[Cay61] M. A. Cayless. Solution of systems of ordinary and partial differential equations by quasi-diagonal matrices. The Computer Journal, 4(1):54–61, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/1/54.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/040054.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/55.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/57.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/59.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/

60.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/61.tif.

Cole:1969:LSP

[CD69] A. J. Cole and A. J. T. Davie. Local smoothing by polynomials in n dimensions. The Computer Journal, 12(1):72–76, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/72.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120072.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/73.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
74.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/75.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/76.tif.

Cooper:1965:CID

[CG65a] G. J. Cooper and E. Gal. Correspondence: Integration of differential equations. The Computer Journal, 8(2):176, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/176.tif.

Cooper:1965:C

[CG65b] G. L. Cooper and E. Gal. Correspondence. The Computer Journal, 8(2):176, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/2/176.2.full.pdf+html.

Coulouris:1968:LCC

[CGH^{+}68] G. F. Coulouris, T. J. Goodey, R. W. Hill, R. W. Keeling, and D. Levin. The London CPL1 compiler. The Computer Journal, 11(1):26–30, May 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/26.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110026.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
26.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/28.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
29.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/30.tif.

Ross:1961:CCC

[cGR61] H. M. c. G. Ross. Considerations in choosing a character code for computers and punched tapes. The Computer Journal, 3(4):202– 210, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/4/202.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/030202.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/202.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/203.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/204.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/205.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/206.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/207.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/208.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/209.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/210.tif.

Cockayne:1960:PNC

[CH60] A. H. Cockayne and E. Hyde. Prime number code for information retrieval. The Computer Journal, 3(1):21–22, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/1/21.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_01/030021.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/21.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/22.tif.

Chartres:1962:AJM

[Cha62] B. A. Chartres. Adaptation of the Jacobi method for a computer with magnetic-tape backing store. The Computer Journal, 5(1): 51–60, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/1/51.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_01/050051.sgm.

abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/51.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/52.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
01/tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/54.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/55.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/56.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/57.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
58.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/59.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/60.tif.

Chapman:1963:ULC

[Cha63] F. G. Chapman. Use of large computers at a distance. The Computer Journal, 6(3):214–218, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/214.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060214.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
214.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/215.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/216.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/217.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/218.tif.

Chawla:1967:NEC

[Cha67] M. M. Chawla. A note on the estimation of the coefficients in the Chebyshev series expansion of a function having a logarithmic singularity. The Computer Journal, 9(4):413, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/413.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090413.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
413.tif.

Chakravarti:1968:BR

[Cha68a] P. C. Chakravarti. Book review. The Computer Journal, 11(3): 323, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/323.full.pdf+html.

Chawla:1968:AEE

[Cha68b] M. M. Chawla. Asymptotic estimates for the error of the Gauss– Legendre quadrature formula. The Computer Journal, 11(3):339– 340, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/339.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/110339.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/339.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/340.tif.

Checksfield:1962:FCR

[Che62a] A. E. Checksfield. The first computer in Rhodesia. The Computer Journal, 5(2):79–87, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/5/2/79.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_02/050079.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/tiff/79.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
02/tiff/80.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/81.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/82.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/84.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
85.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/86.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/87.tif.

Cheesman:1962:SBC

[Che62b] D. V. Cheesman. A small business computer at work. The Computer Journal, 5(1):1–6, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/5/1/1.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_01/050001.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/1.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_

01/tiff/2.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/3.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_01/tiff/4.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/6.tif.

Cherry:1963:BR

[Che63] Colin Cherry. Book review. The Computer Journal, 5(4):262– 263, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/4/262.2.full.pdf+html.

Chew:1964:NTS

[Che64] R. L. Chew. Note on timing simulation of a large asynchronous computer. The Computer Journal, 7(2):122–123, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/122.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070122.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
122.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/123.tif.

Chedzoy:1965:PP

[Che65a] O. B. Chedzoy. Perspectives in programming. The Computer Journal, 8(1):12a–12, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/12-a.full.pdf+html.

Chezem:1965:NTD

[Che65b] C. G. Chezem. Note on “three-dimensional” plotting as a technique for finding the zeros of functions in the complex plane. The Computer Journal, 8(3):288, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/288.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080288.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
288.tif.

Churchhouse:1968:DCN

[Chu68] R. F. Churchhouse. Discussion and correspondence: note on the twenty-five most cited papers in some leading journals. The Computer Journal, 11(1):116–120, May 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/116.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110116.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
116.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/117.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_01/tiff/118.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_01/tiff/119.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/120.tif.

Churchhouse:1969:BR

[Chu69] R. F. Churchhouse. Book review. The Computer Journal, 12 (2):159, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/159.full.pdf+html.

Clark:1966:CMP

[CL66a] S. R. Clark and W. F. Lunnon. Correspondence: Multiple precision arithmetic (real and complex). The Computer Journal, 9(2): 174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/174.tif. See [BRC66].

Clark:1966:MPA

[CL66b] S. R. Clark and W. F. Lunnon. Multiple precision arithmetic in Atlas Autocode. The Computer Journal, 9(2):174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
2/174.full.pdf+html.

Clarke:1960:NSD

[Cla60] L. Clarke. Notes on the state of digital computing in the U.S.S.R. The Computer Journal, 3(3):164–167, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/164.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030164.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
164.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/165.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_03/Issue_03/tiff/166.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_03/tiff/167.tif.

Clarke:1962:BR

[Cla62] L. Clarke. Book review. The Computer Journal, 4(4):304, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
4/304.full.pdf+html.

Clarke:1966:BJC

[Cla66] L. T. G. Clarke. International Journal of Computer Mathematics. The Computer Journal, 8(4):351, January 1966. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/8/4/351.
full.pdf+html.

Clenshaw:1960:CFD

[Cle60] C. W. Clenshaw. Curve fitting with a digital computer. The Computer Journal, 2(4):170–173, January 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/4/170.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/020170.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/
170.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/tiff/171.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_04/tiff/172.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/
Issue_04/tiff/173.tif.

Clenshaw:1963:BRN

[Cle63] C. W. Clenshaw. Book reviews: Numerical analysis. The Computer Journal, 6(2):207–208, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/6/2/207.full.pdf+html.

Clenshaw:1965:BRI

[Cle65] C. W. Clenshaw. Book review: Integration of Equations of Parabolic Type by the Method of Nets, V. K. Saul’yev, 1964; 346 pages. (Oxford: Pergamon Press Ltd.). The Computer Journal, 8 (2):156, February 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/2/156.full.pdf+html.

Clementson:1966:ECS

[Cle66a] A. T. Clementson. Extended control and simulation language. The Computer Journal, 9(3):215–220, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/215.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090215.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
215.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/216.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/217.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/218.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/219.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/220.tif.

Clenshaw:1966:BRA

[Cle66b] C. W. Clenshaw. Book review: Approximation of Functions, edited by H. Garabedian, 1965; 215 pages. The Computer Journal, 9(3): 285, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/285.full.pdf+html.

Cleave:1967:BR

[Cle67a] J. P. Cleave. Book review. The Computer Journal, 10(3):264, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/264.full.pdf+html.

Clementson:1967:ALS

[Cle67b] A. T. Clementson. An assembly and loading system for computers with parallel peripheral operation. The Computer Journal, 9(4):370–372, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/370.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090370.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
370.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_09/Issue_04/tiff/371.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/372.tif.

Clenshaw:1968:BRC

[Cle68] C. W. Clenshaw. Book review: Chebyshev Polynomials in Numerical Analysis, by L. Fox and I. B. Parker, 1969; 205 pages (Oxford University Press). The Computer Journal, 11(3):310, March 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/310.full.pdf+html.

Clippinger:1962:C

[Cli62a] R. F. Clippinger. COBOL. The Computer Journal, 5(3):177–180, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/177.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_03/050177.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/177.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/180.tif.

Clippinger:1962:F

[Cli62b] R. F. Clippinger. FACT. The Computer Journal, 5(2):112–119, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/2/112.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/050112.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/112.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/113.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/
tiff/114.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/115.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/116.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/117.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/118.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/120.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/121.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/122.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/123.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/
tiff/124.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/125.tif.

Clippinger:1962:IA

[Cli62c] R. F. Clippinger. Information algebra. The Computer Journal, 5(3):180–183, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/5/3/180.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
03/050180.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/180.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/181.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/182.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/183.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/184.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/185.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/186.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/187.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/188.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/189.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/190.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/191.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/192.tif.

Cloot:1965:WUO

[Clo65] P. L. Cloot. What is the use of operating systems? The Computer Journal, 7(4):249–254, January 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/249.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070249.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
249.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/250.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_07/Issue_04/tiff/251.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/252.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/253.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/254.tif.

Cluley:1963:BRB

[Clu63a] J. C. Cluley. Book review: Analogue Computers. The Computer Journal, 6(1):112, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/6/1/112.full.pdf+html.

Cluley:1963:BR

[Clu63b] J. C. Cluley. Book reviews. The Computer Journal, 6(3):296, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/296.full.pdf+html.

Clenshaw:1963:SNO

[CN63] C. W. Clenshaw and H. J. Norton. The solution of nonlinear ordinary differential equations in Chebyshev series. The Computer Journal, 6(1):88–92, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/6/1/88.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060088.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/88.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/89.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/90.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/91.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/92.tif.

Cohen:1966:NOG

[CND66] J. Cohen and X. Nguyen-Dinh. Note on ordering of grammar rules in syntax-analyzers. The Computer Journal, 9(3):250–251, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/250.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090250.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

09/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/251.tif.

Curtis:1966:CMR

[CO66] Alan Curtis and M. R. Osborne. The construction of minimax rational approximations to functions. The Computer Journal, 9(3):286– 293, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/286.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/090286.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/286.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/287.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/288.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/289.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/290.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/291.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/292.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/293.tif.

Cobb:1969:NAB

[Cob69] S. M. Cobb. Note on Algorithm 35: Best rational approximation to a real number. The Computer Journal, 12(3):293– 295, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/293.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/294.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/295.tif. See [Pre68, PP69].

Colin:1963:NCR

[Col63a] A. J. T. Colin. Note on coding reverse Polish expressions for single-address computers with one accumulator. The Computer Journal, 6(1):67–68, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/6/1/67.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060067.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/67.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/68.tif.

Collison:1963:NMF

[Col63b] D. M. Collison. Note on a method of forming a sorting key for a partly ordered list, and an application. The Computer Journal, 6 (1):74, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/74.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060074.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/74.tif.

Cole:1964:PET

[Col64a] A. J. Cole. The preparation of examination time-tables using a small-store computer. The Computer Journal, 7(2):117–121, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/2/117.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_02/070117.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/117.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/118.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/120.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/121.
tif.

Colin:1964:MVC

[Col64b] Andrew Colin. The multiple variate counter. The Computer Journal, 6(4):339–347, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/339.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060339.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
339.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/340.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/341.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_04/tiff/342.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/343.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/

tiff/344.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/345.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/346.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/347.tif.

Collins:1964:BR

[Col64c] J. S. Collins. Book review. The Computer Journal, 7(3):207, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
3/207.full.pdf+html.

Cole:1966:PSP

[Col66a] A. J. Cole. Plane and stereographic projections of convex polyhedra from minimal information. The Computer Journal, 9(1): 27–31, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/1/27.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_01/090027.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/29.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/31.
tif.

Colin:1966:SPU

[Col66b] A. J. T. Colin. A simple program for use in the ‘conversational mode’. The Computer Journal, 9(3):238–241, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/238.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090238.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/239.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/241.tif.

Cole:1967:CCN

[Col67] A. J. Cole. Correspondence: Chromatic numbers and timetabling problems. The Computer Journal, 10(2):217, August 1967. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_02/tiff/217.tif. See [WP67].

Constantinescu:1966:CSE

[Con66] Paul Constantinescu. The classification of a set of elements with respect to a set of properties. The Computer Journal, 8(4):352–357, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/352.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080352.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/352.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/353.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/355.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/356.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/357.tif.

Cook:1960:MSS

[Coo60a] R. L. Cook. Market surveys with a small computer. The Computer Journal, 3(3):140–142, October 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/140.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030140.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
140.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/141.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_03/tiff/142.tif.

Cook:1960:TSN

[Coo60b] R. L. Cook. Time-sharing on the National-Elliott 802. The Computer Journal, 2(4):185–188, January 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/4/185.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/020185.sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/
185.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/tiff/186.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_04/tiff/187.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/
Issue_04/tiff/188.tif.

Cooper:1966:ECC

[Coo66] D. C. Cooper. The equivalence of certain computations. The Computer Journal, 9(1):45–52, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/45.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090045.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/46.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/47.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/48.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/49.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/50.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
51.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/52.tif.

Cooper:1967:ECC

[Coo67] G. J. Cooper. The evaluation of the coefficients in the Chebyshev expansion. The Computer Journal, 10(1):94–100, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/1/94.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_10/Issue_01/100094.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/100.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/94.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
95.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/96.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/97.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
98.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/99.tif.

Cooper:1968:BSI

[Coo68a] B. E. Cooper. Basic subroutine for the input of numbers, words and special characters. The Computer Journal, 11(2):157–159, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/157.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110157.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/157.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/158.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/159.
tif.

Cooper:1968:SMA

[Coo68b] D. C. Cooper. Sequential machines and automata theory. The Computer Journal, 11(2):176, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/2/176.full.pdf+html.

Cooper:1969:BRB

[Coo69] D. C. Cooper. Book review: Theory of Recursive Functions and Effective Computability. The Computer Journal, 12(1):28, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
1/28.full.pdf+html.

Cormack:1962:EOE

[Cor62] A. S. Cormack. Early operating experience with language H. The Computer Journal, 5(3):158–161, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/158.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/050158.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/
158.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/159.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_03/tiff/160.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_03/tiff/161.tif.

Cormack:1965:IBU

[Cor65] A. S. Cormack. Interaction between user’s needs and languagecompiler-computer systems. The Computer Journal, 8(1):8– 12, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/8.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/080008.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/10.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/11.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/8.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/tiff/9.tif.

Clowes:1961:NTA

[CP61] M. B. Clowes and J. R. Parks. A new technique in automatic character recognition. The Computer Journal, 4(2):121–128, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/2/121.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_02/040121.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/121.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/122.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/123.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/124.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/125.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/126.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/127.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/128.tif.

Curtis:1962:PTL

[CP62] A. R. Curtis and I. C. Pyle. A proposed target language for compilers on Atlas. The Computer Journal, 5(2):100–106, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/2/100.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/050100.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/100.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/101.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/
tiff/102.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_05/Issue_02/tiff/103.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/104.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/105.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/106.tif.

Clowes:1964:AP

[CP64] J. S. Clowes and E. S. Page. Assignment problems. The Computer Journal, 6(4):304–307, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/304.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060304.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
304.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/305.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/306.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_04/tiff/307.tif.

Curtis:1966:NCM

[CP66a] A. R. Curtis and M. J. D. Powell. Necessary conditions for a minimax approximation. The Computer Journal, 8(4):358–361, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/358.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080358.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/358.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/359.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/360.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/361.tif.

Curtis:1966:CEA

[CP66b] A. R. Curtis and M. J. D. Powell. On the convergence of exchange algorithms for calculating minimax approximations. The Computer Journal, 9(1):78–80, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/78.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/tiff/78.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/79.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
80.tif.

Cranley:1968:EMI

[CP68] R. Cranley and T. N. L. Patterson. The evaluation of multidimensional integrals. The Computer Journal, 11(1):102–110, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/102.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/110102.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
01/tiff/102.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/103.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/104.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/105.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/106.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/
tiff/107.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/108.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/109.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/110.tif.

Cowdrey:1963:AMC

[CR63] D. R. Cowdrey and C. M. Reeves. An application of the Monte Carlo method to the evaluation of some molecular integrals. The Computer Journal, 6(3):277–286, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/277.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060277.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
277.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/278.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/279.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/280.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/281.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/282.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/283.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/284.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

06/Issue_03/tiff/285.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/286.tif.

Craddock:1966:BRB

[Cra66a] J. M. Craddock. Book review: Programming Systems for use in Research Computation. The Computer Journal, 8(4):315–318, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/315.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080315.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/315.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/316.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/317.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/318.tif.

Craig:1966:CCR

[Cra66b] T. K. Craig. CRESTS — Courtauld’s rapid extract, sort and tabulate system. The Computer Journal, 9(1):3–10, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
9/1/3.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/090003.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/10.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/3.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_01/tiff/4.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
7.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/8.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/9.tif.

Craddock:1967:CAC

[Cra67] I. D. Craddock. Correspondence: On analogue computers. The Computer Journal, 10(2):217, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/217.tif.

Cuddeford:1967:CCI

[Cud67] G. C. Cuddeford. Computer control of industrial processes. The Computer Journal, 10(2):161, August 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/161.full.
pdf+html.

Curtis:1965:ETE

[Cur65] A. R. Curtis. Estimation of the truncation error in Runge– Kutta and allied processes. The Computer Journal, 8(1):52, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
1/52.full.pdf+html.

Curtis:1967:BR

[Cur67] A. R. Curtis. Book review. The Computer Journal, 10(1):52, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/52.full.pdf+html.

Currie:1968:BR

[Cur68] I. F. Currie. Book review. The Computer Journal, 11(3):331, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/331.full.pdf+html.

Cadwell:1961:SOM

[CW61] J. H. Cadwell and D. E. Williams. Some orthogonal methods of curve and surface fitting. The Computer Journal, 4(3):260–264, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/3/260.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_03/040260.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/260.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/261.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/262.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/263.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/264.
tif.

Cooper:1962:AAL

[CW62] D. C. Cooper and H. Whitfield. ALP: an Autocode listprocessing language. The Computer Journal, 5(1):28–32, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/1/28.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/050028.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/28.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/29.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
30.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/31.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/32.tif.

dAgapeyeff:1962:CDC

[d’A62] A. d’Agapeyeff. Current developments in commercial automatic programming. The Computer Journal, 5(2):107–111, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/2/107.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/050107.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/107.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/108.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/
tiff/109.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/110.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/111.
tif.

Dace:1963:EPU

[Dac63] D. J. Dace. Experience in the practical use of data transmission. The Computer Journal, 6(1):17–23, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/6/1/17.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060017.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/18.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/19.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/20.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/21.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/22.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
23.tif.

Dakin:1965:TSA

[Dak65] R. J. Dakin. A tree search algorithm for mixed integer programming problems. The Computer Journal, 8(3):250–255, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/3/250.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_03/080250.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/251.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/253.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/255.tif.

Davey:1963:DPT

[Dav63] Wing Commander L. B. Davey. Data preparation and transmission in the Royal Air Force integrated supply system. The Computer Journal, 6(3):219–222, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/219.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060219.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
219.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/220.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/221.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/222.tif.

Davies:1967:BR

[Dav67a] M. Davies. Book review. The Computer Journal, 9(4):329, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
4/329.full.pdf+html.

Davison:1967:HOC

[Dav67b] E. J. Davison. A high-order Crank–Nicolson technique for solving differential equations. The Computer Journal, 10(2):195– 197, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/195.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/100195.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/195.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/196.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/197.tif.

Davidon:1968:VAM

[Dav68] William C. Davidon. Variance algorithm for minimization. The Computer Journal, 10(4):406–410, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/406.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100406.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
406.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/407.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/408.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/409.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/410.tif.

Day:1965:OSM

[Day65] J. T. Day. A one-step method for the numerical integration of the differential equation y′ = f(x)y + g(x). The Computer Journal, 7(4):314–317, January 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/314.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070314.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
314.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/315.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/316.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/317.tif.

Day:1966:RKM

[Day66] J. T. Day. A Runge–Kutta method for the numerical solution of the Goursat problem in hyperbolic partial differential equations. The Computer Journal, 9(1):81–83, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/81.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090081.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/81.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/82.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/83.tif.

Day:1967:NNS

[Day67] James Thomas Day. Note on the numerical solution of integrodifferential equations. The Computer Journal, 9(4):394–395, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/394.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090394.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/394.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/395.tif.

Day:1969:ASA

[Day69] A. C. Day. Algorithm 38: a searching algorithm. The Computer Journal, 12(1):101–102, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/
tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/102.tif.

Dawkins:1970:CUF

[DD70] B. W. Dawkins and H. C. Dawkins. Correspondence: On “A University faculty timetable”. The Computer Journal, 13(1):121– 122, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/121.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/122.tif. See [Alm69].

Delves:1968:NEP

[Del68] L. M. Delves. The numerical evaluation of principal value integrals. The Computer Journal, 10(4):389–391, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/389.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100389.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
389.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/390.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/391.tif.

Dempsey:1960:SAS

[Dem60] J. M. Dempsey. Some aspects of simulator design. The Computer Journal, 3(3):158–161, October 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/158.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030158.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
158.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/159.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_03/tiff/160.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_03/tiff/161.tif.

Deutsch:1969:CCL

[Deu69] E. S. Deutsch. Correspondence: Comments on a line thinning algorithm. The Computer Journal, 12(4):412, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/4/412.2.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/412.tif.

See correspondence [Fra70].

deFerranti:1966:HB

[dF66] Basil de Ferranti. The human brain. The Computer Journal, 9(2): 117–123, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/2/117.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/090117.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/117.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_09/Issue_02/tiff/118.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/119.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/120.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/121.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/122.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/123.
tif.

Davison:1968:AMT

[DGM68] P. S. Davison, P. Giles, and D. A. R. Matthews. Ageing of magnetic tape: a critical bibliography and comparison of literature sources. The Computer Journal, 11(3):241–246, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/241.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110241.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
241.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/242.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/243.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/244.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/245.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/246.tif.

Duncan:1960:DAT

[DH60] F. G. Duncan and D. H. R. Huxtable. The DEUCE Alphacode translator. The Computer Journal, 3(2):98–107, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/2/98.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_02/030098.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/100.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/101.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/
tiff/102.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/103.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/104.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/105.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_02/tiff/106.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/
tiff/107.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/98.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/99.
tif.

dInverno:1969:AAL

[d’I69] R. A. d’Inverno. ALAM — Atlas Lisp Algebraic Manipulator. The Computer Journal, 12(2):124–127, May 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/2/124.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/120124.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/
124.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/125.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/126.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/127.tif.

Dijkstra:1962:OEA

[Dij62] E. W. Dijkstra. Operating experience with Algol 60. The Computer Journal, 5(2):125–127, August 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/2/125.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/050125.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
125.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/126.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/tiff/127.tif.

Dirac:1968:BR

[Dir68] G. A. Dirac. Book review. The Computer Journal, 11(3):340, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/340.full.pdf+html.

Douglas:1960:ALS

[DM60] A. S. Douglas and A. J. Mitchell. AUTOSTAT: a language for statistical data processing. The Computer Journal, 3(2):61– 66, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/2/61.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_02/030061.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/61.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_02/tiff/62.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
02/tiff/63.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/65.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/66.tif.

DSylva:1964:RIS

[DM64] E. D’Sylva and G. A. Miles. The S.S.O.R. iteration scheme for equations with σ_{1} ordering. The Computer Journal, 6(4):366–367, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/4/366.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_04/060366.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/366.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/367.tif.

Dearnley:1964:ASP

[DN64] F. H. Dearnley and G. B. Newell. Automatic segmentation of programs for a two-level store computer. The Computer Journal, 7(3):185–187, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/185.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070185.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
185.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/186.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/187.tif.

Dollimore:1966:GFS

[Dol66] Jean Dollimore. A general Fourier synthesis program for the London University Atlas computer. The Computer Journal, 8(4):347–351, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/347.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_08/Issue_04/080347.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/347.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/348.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/349.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/350.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/351.
tif.

Doran:1966:PR

[Dor66] James Doran. Pattern recognition. The Computer Journal, 9(3): 297, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/297.full.pdf+html.

Doran:1967:BR

[Dor67] J. Doran. Book review. The Computer Journal, 10(3):299, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/299.full.pdf+html.

Doran:1968:BR

[Dor68] J. Doran. Book review. The Computer Journal, 11(1):62, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/62.full.pdf+html.

Douglas:1961:ICD

[Dou61] A. S. Douglas. The impact of computers on documentation. The Computer Journal, 4(2):145–149, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/145.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040145.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
145.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/146.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/147.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/148.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/149.tif.

Douglas:1965:ULC

[Dou65] A. S. Douglas. The use of a large computer on a bureau basis. The Computer Journal, 7(4):262–263, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/262.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070262.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
262.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/263.tif.

Douglas:1966:BRB

[Dou66] A. S. Douglas. Book review: IFIP-ICC Vocabulary of Information Processing. The Computer Journal, 9(2):136, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
2/136.full.pdf+html.

Douce:1969:BR

[Dou69] J. L. Douce. Book review. The Computer Journal, 12(2):187, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/187.full.pdf+html.

Downham:1969:RT

[Dow69] D. Y. Downham. The runs up and down test. The Computer Journal, 12(4):373–376, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/373.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120373.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
373.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/374.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/375.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/376.tif.

dePaula:1960:PAC

[dP60] F. C. de Paula. Problems of auditing computer data. Section 2: The external auditor and computers. The Computer Journal, 3(1):11–14, April 1960. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/1/11.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/030011.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/11.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/14.
tif.

Doust:1964:LRV

[DP64] A. Doust and V. E. Price. The latent roots and vectors of a singular matrix. The Computer Journal, 7(3):222–227, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/3/222.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_03/070222.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/222.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/223.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/224.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_03/tiff/225.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/226.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/227.tif.

Dineley:1966:KAB

[DP66] J. L. Dineley and C. Preece. KALDAS, an algorithmically based digital simulation of analogue computation. The Computer Journal, 9(2):181–187, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/181.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090181.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
181.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/182.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/183.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/184.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/185.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/

tiff/186.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/187.tif.

Downham:1967:MCP

[DR67] D. Y. Downham and F. D. K. Roberts. Multiplicative congruential pseudo-random number generators. The Computer Journal, 10(1): 74–77, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/1/74.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_10/Issue_01/100074.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/74.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/75.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
76.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/77.tif.

Dresel:1966:ENA

[Dre66] L. A. G. Dresel. Elementary numerical analysis. The Computer Journal, 9(2):199a–199, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/2/199-a.full.pdf+html.

Drummond:1963:SAR

[Dru63] J. Drummond. Some aspects of recording graduated national insurance contributions. The Computer Journal, 6(1):1– 4, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/1.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060001.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/1.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/2.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/3.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/4.tif.

Deist:1967:SSN

[DS67] F. H. Deist and L. Sefor. Solution of systems of non-linear equations by parameter variation. The Computer Journal, 10(1):78–82, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/1/78.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_10/Issue_01/100078.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/78.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/79.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
80.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/81.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/82.tif.

Duke:1966:CRL

[Duk66a] B. J. Duke. Computation of rotational levels. The Computer Journal, 9(2):210, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/2/210.full.pdf+html.

Duke:1966:CCR

[Duk66b] B. J. Duke. Correspondence: On the computation of rotational levels of rigid asymmetric top molecules. The Computer Journal, 9(2):210, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_09/Issue_02/090210.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/210.tif. See note [Rac65] and corre-

spondence [Jon66].

Duncan:1962:EIA

[Dun62a] F. G. Duncan. Errata: “Implementation of ALGOL 60 for the English Electric KDF9”. The Computer Journal, 5(3):176, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/176.tif. See [Dun62b].

Duncan:1962:IAE

[Dun62b] F. G. Duncan. Implementation of Algol 60 for the English Electric KDF9. The Computer Journal, 5(2):130–131, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/2/130.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/050130.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/131.tif. See errata

[Dun62a].

Duncan:1963:IOA

[Dun63] F. G. Duncan. Input and output for ALGOL 60 on KDF 9. The Computer Journal, 5(4):341–344, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/341.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050341.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
341.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/342.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/343.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/344.tif.

Duncan:1967:RAL

[Dun67] F. G. Duncan. Readings in automatic lanuage processing. The Computer Journal, 9(4):364, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/364.full.
pdf+html.

deWitte:1964:EES

[dW64] L. de Witte. Error estimates for smoothing and extrapolation formulae. The Computer Journal, 7(3):220–221, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/3/220.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_03/070220.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/221.tif.

Edwards:1966:AMT

[EAL66] D. B. G. Edwards, D. Aspinall, and R. Lane. Addressing magnetic tape on the Atlas 1 computing machine. The Computer Journal, 8(4):323–329, January 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/323.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080323.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
323.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/324.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_08/Issue_04/tiff/325.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_04/tiff/326.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/327.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/328.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/329.tif.

Edelstein:1963:PLB

[Ede63] L. A. Edelstein. “picture logic” for “Bacchus” a fourthgeneration computer. The Computer Journal, 6(2):144–153, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/2/144.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_02/060144.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/144.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/145.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/147.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/150.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/151.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/152.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/153.
tif.

Edelstein:1965:NAA

[Ede65] L. A. Edelstein. Note on analytic angular integrations using series in Legendre polynomials. The Computer Journal, 8(2):157–158, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
2/157.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/157.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/158.tif.

Edmonds:1960:GPR

[Edm60] A. R. Edmonds. The generation of pseudo-random numbers on electronic digital computers. The Computer Journal, 2(4):181–185, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/2/4/181.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_04/020181.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_04/tiff/181.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_04/tiff/182.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/
tiff/183.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_04/tiff/184.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/185.
tif.

Evans:1963:IPO

[EF63a] D. J. Evans and C. V. D. Forrington. An iterative process for optimizing symmetric successive over-relaxation. The Computer Journal, 6(3):271–273, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/271.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060271.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
271.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/272.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/273.tif.

Evans:1963:NSC

[EF63b] D. J. Evans and C. V. D. Forrington. Note on the solution of certain tri-diagonal systems of linear equations. The Computer Journal, 5(4):327–328, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/327.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050327.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
327.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/328.tif.

El-gendi:1969:CSD

[Eg69] S. E. El-gendi. Chebyshev solution of differential, integral and integro differential equations. The Computer Journal, 12(3):282– 287, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/3/282.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_12/Issue_03/120282.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/282.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/283.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/284.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/285.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/286.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/287.tif.

Elcock:1965:NAL

[Elc65] E. W. Elcock. Note on the addressing of lists by their sourcelanguage names. The Computer Journal, 8(3):242–243, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/3/242.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_03/080242.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/242.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/243.tif.

Elliott:1963:CSM

[Ell63] D. Elliott. A Chebyshev series method for the numerical solution of Fredholm integral equations. The Computer Journal, 6(1):102–112, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/1/102.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/060102.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/103.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/
tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/105.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/106.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/107.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/108.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/
tiff/109.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/110.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/111.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/112.tif.

England:1969:EER

[Eng69] R. England. Error estimates for Runge–Kutta type solutions to systems of ordinary differential equations. The Computer Journal, 12 (2):166–170, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/166.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120166.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/166.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/167.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/169.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/170.tif.

Erdelyi:1965:SCI

[Erd65] I. Erdelyi. On speeding convergence of an iterative eigenvalue process. The Computer Journal, 8(2):159–165, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/2/159.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/080159.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/
159.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/160.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/tiff/161.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/165.tif.

Evans:1962:NLR

[Eva62] D. J. Evans. Note on the line over-relaxation factor for small mesh size. The Computer Journal, 5(1):48–50, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/1/48.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/050048.sgm.abs.

html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/48.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/49.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
50.tif.

Evans:1963:EMA

[Eva63] D. J. Evans. The extrapolated modified Aitken iteration method for solving elliptic difference equations. The Computer Journal, 6(2):193–201, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/193.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060193.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
193.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/194.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/195.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_02/tiff/196.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/197.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/198.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/199.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/200.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/201.tif.

Evans:1964:EMA

[Eva64] D. J. Evans. The extrapolated modified Aitken iteration method applied to σ_{1}-ordered systems of linear equations. The Computer Journal, 7(2):137–140, July 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/137.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070137.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
137.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/138.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/139.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/tiff/140.tif.

Evans:1965:ELR

[Eva65a] D. J. Evans. Estimation of the line over-relaxation factor and convergence rates of an alternating direction line overrelaxation technique. The Computer Journal, 7(4):318–321, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/4/318.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_04/070318.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_04/tiff/318.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/319.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/320.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/321.tif.

Evans:1965:SEM

[Eva65b] D. J. Evans. A stable explicit method for the finite-difference solution of a fourth-order parabolic partial differential equation. The Computer Journal, 8(3):280–287, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/280.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080280.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
280.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/281.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/282.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/284.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/285.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/286.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/287.

tif. See comments [Eva66].

Evans:1966:CSE

[Eva66] D. J. Evans. Correspondence: On “A stable explicit method for the finite-difference solution of a fourth-order parabolic partial differential equation”. The Computer Journal, 9(1):115, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/115.tif. See [Eva65b].

Eve:1963:SAI

[Eve63] J. Eve. Starting approximations for the iterative calculation of square roots. The Computer Journal, 6(3):274–276, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/3/274.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_03/060274.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_03/tiff/274.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/275.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/276.tif.

Eve:1968:BR

[Eve68] J. Eve. Book review. The Computer Journal, 11(3):316, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/316.full.pdf+html.

Eve:1969:BR

[Eve69] J. Eve. Book review. The Computer Journal, 12(2):157, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/157.full.pdf+html.

Fairthorne:1961:C

[Fai61] R. A. Fairthorne. Correspondence. The Computer Journal, 4(3): 216, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/3/216.full.pdf+html.

Fox:1969:ISD

[FE69] A. J. Fox and P. W. Edwards. Implementation of a syntax-driven interpreter for data retrieval. The Computer Journal, 12(3):225– 232, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/3/225.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/120225.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/225.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/226.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/

Issue_03/tiff/227.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/228.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/229.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/230.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/231.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/232.tif.

Fellgett:1963:C

[Fel63] Peter Fellgett. Correspondence. The Computer Journal, 5(4): 348, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/4/348.full.pdf+html.

Fellgett:1964:C

[Fel64] Peter Fellgett. Correspondence. The Computer Journal, 7(1):85, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/1/85.full.pdf+html.

Fellgett:1969:CWA

[Fel69] Peter Fellgett. Correspondence: What is an analyst? The Computer Journal, 12(1):104, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/104.

tif. See correspondence [Fel70, AT70].

Fellgett:1970:CWA

[Fel70] P. B. Fellgett. Correspondence: What is an analyst? The Computer Journal, 13(1):118–119, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/119.tif. See [Fel69, AT70].

Fettis:1965:NME

[Fet65] H. E. Fettis. Note on the matrix equation Ax = λBx. The Computer Journal, 8(3):279, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/279.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080279.sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
279.tif.

Fair:1966:NCA

[FFMR66] G. R. Fair, A. D. J. Flowerdew, W. G. Munro, and D. Rowley. Note on the computer as an aid to the architect. The Computer Journal, 9(1):16–20, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/16.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090016.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/16.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/17.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/19.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/20.tif.

Forrington:1967:AOC

[FGR67] C. V. D. Forrington, Sheila M. Grey, and M. M. Reid. An application-oriented computer approach to inventory control problems. Part 2: Implementation and operational experience. The Computer Journal, 9(4):325–329, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/325.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090325.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
325.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/326.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/327.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/329.tif.

Fieldhouse:1964:BRL

[Fie64] Martin Fieldhouse. Book review: Linear programming. The Computer Journal, 7(2):148, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/148.full.pdf+html.

Fielding:1970:NAS

[Fie70] K. Fielding. Note on Algorithm 44: Solution of nonlinear simultaneous equations. The Computer Journal, 13(2):219, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_02/tiff/219.tif. See [Bro69c, Red71].

Findler:1960:SRG

[Fin60] N. V. Findler. Some remarks on the game ‘dama’ which can be played on a digital computer. The Computer Journal, 3(1): 40–44, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/1/40.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_01/030040.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/40.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/41.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/42.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/43.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/44.
tif.

Finkelstein:1968:COT

[Fin68] M. Finkelstein. A compiler optimization technique. The Computer Journal, 11(1):22–25, May 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/22.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110022.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
22.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/23.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/24.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
25.tif.

Firth:1961:OPS

[Fir61] A. W. O. Firth. Optimization problems: solution by an analogue computer. The Computer Journal, 4(1):68–72, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/1/68.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_04/Issue_01/040068.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/68.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/69.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/
70.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/71.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/72.tif.

Fox:1966:FER

[FJ66] A. J. Fox and F. A. Johnson. On finding the eigenvalues of real symmetric tridiagonal matrices. The Computer Journal, 9(1): 98–105, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/1/98.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_01/090098.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_01/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/103.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/
tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/105.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/98.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/99.tif.

Foxley:1968:ISA

[FK68] Eric Foxley and Peter King. The implementation of syntax analysis using ALGOL, and some mathematical applications. The Computer Journal, 10(4):325–335, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/325.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100325.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
325.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/326.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/327.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/329.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/330.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/331.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/332.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/333.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/334.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/335.tif.

Foxley:1968:CET

[FL68] E. Foxley and K. Lockyer. The construction of examination timetables by computer. The Computer Journal, 11(3):264–268, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/264.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/110264.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/264.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/265.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/266.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/267.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/268.tif.

Fletcher:1965:FME

[Fle65] R. Fletcher. Function minimization without evaluating derivatives — a review. The Computer Journal, 8(1):33–41, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/1/33.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/080033.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/34.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
35.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/37.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
38.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/40.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
41.tif.

Fletcher:1968:GIM

[Fle68] R. Fletcher. Generalized inverse methods for the best least squares solution of systems of non-linear equations. The Computer Journal, 10(4):392–399, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/392.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100392.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
392.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/393.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/394.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/395.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/396.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/397.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/398.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/399.
tif.

Flores:1964:NMA

[Flo64] Ivan Flores. Note on a machine algorithm for conversion from reflected binary to natural binary. The Computer Journal, 7(2): 121, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/2/121.full.pdf+html.

Florentin:1968:BRa

[Flo68a] J. J. Florentin. Book review. The Computer Journal, 10(4): 346, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/346.full.pdf+html.

Florentin:1968:BRb

[Flo68b] J. J. Florentin. Book review. The Computer Journal, 10(4): 410, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/410.full.pdf+html.

Florentin:1968:MI

[Flo68c] J. J. Florentin. Machine intelligence 2. The Computer Journal, 11 (2):156, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/156.full.pdf+html.

Fairweather:1964:GAD

[FM64] G. Fairweather and A. R. Mitchell. A generalized alternating direction method of Douglas-Rachford type for solving the biharmonic equation. The Computer Journal, 7(3):242–245, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/3/242.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_03/070242.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/242.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/243.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/244.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_03/tiff/245.tif.

Fairweather:1966:SCR

[FM66] G. Fairweather and A. R. Mitchell. Some computational results of an improved A.D.I. method for the Dirichlet problem. The Computer Journal, 9(3):298–303, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/298.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090298.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
298.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/299.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/300.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/301.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/302.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/303.tif.

Forrington:1961:EPC

[For61] C. V. D. Forrington. Extensions of the predictor-corrector method for the solution of systems of ordinary differential equations. The Computer Journal, 4(1):80–84, April 1961. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/1/80.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/040080.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/80.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/81.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/82.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/83.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/84.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/85.tif.

Forbes:1965:ICW

[For65] J. M. Forbes. An introduction to compiler writing. The Computer Journal, 8(2):98–102, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/8/2/98.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/080098.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/tiff/100.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
02/tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/102.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/98.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/99.tif.

Forrington:1966:AOC

[For66] C. V. D. Forrington. An application-oriented computer approach to inventory control problems. Part 1: Facilities of the system. The Computer Journal, 9(2):129–136, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/129.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090129.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
129.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/130.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/131.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/132.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/133.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/

tiff/134.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/135.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/136.
tif.

Foster:1964:CTG

[Fos64] F. G. Foster. A computer technique for game-theoretic problems I: chemin-de-fer analyzed. The Computer Journal, 7(2):124–130, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/2/124.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_02/070124.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/124.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/125.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/127.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/128.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/129.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/130.tif.

Foster:1965:SLP

[Fos65a] D. M. Foster. A simple list-processing interpreter. The Computer Journal, 8(2):120–129, July 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/2/120.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/080120.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/
120.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/121.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/tiff/122.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/123.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/124.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/125.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/126.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/127.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/128.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/129.tif.

Foster:1965:MTA

[Fos65b] J. M. Foster. Mathematical theory of automata. The Computer Journal, 8(1):87, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/87.full.pdf+html.

Foster:1968:SAO

[Fos68a] Caxton C. Foster. Sorting almost ordered arrays. The Computer Journal, 11(2):134–137, August 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/2/134.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/110134.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/
134.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/135.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/136.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/137.tif.

Foster:1968:SIP

[Fos68b] J. M. Foster. A syntax improving program. The Computer Journal, 11(1):31–34, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/1/31.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_01/110031.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/31.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/tiff/32.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
01/tiff/33.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/34.tif.

Fox:1961:BR

[Fox61a] L. Fox. Book review. The Computer Journal, 4(1):78–79, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
1/78.full.pdf+html.

Fox:1961:CMT

[Fox61b] L. Fox. Computing machines for teaching and research. The Computer Journal, 4(3):212–216, October 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/3/212.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/040212.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/
212.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/213.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_03/tiff/214.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_03/tiff/215.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/216.tif.

Fox:1962:CMO

[Fox62] L. Fox. Chebyshev methods for ordinary differential equations. The Computer Journal, 4(4):318–331, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/4/318.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/040318.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/
318.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/319.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_04/tiff/320.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/321.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/322.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/323.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/324.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/325.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/326.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/327.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/328.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/329.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/330.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/331.tif.

Fox:1963:PDE

[Fox63] L. Fox. Partial differential equations. The Computer Journal, 6 (1):69–74, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/69.full.pdf+html;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_06/Issue_01/060069.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/70.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/71.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/72.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/73.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/74.tif.

Fox:1967:RIC

[Fox67] L. Fox. Romberg integration for a class of singular integrands. The Computer Journal, 10(1):87–93, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/87.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100087.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
87.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
90.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/92.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
93.tif.

Fox:1968:BR

[Fox68] L. Fox. Book review. The Computer Journal, 11(1):30, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/30.full.pdf+html.

Fox:1969:CEM

[Fox69] L. Fox. Computer evaluation of mathematical function. The Computer Journal, 12(3):267, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/3/267.full.pdf+html.

Fletcher:1963:RCD

[FP63] R. Fletcher and M. J. D. Powell. A rapidly convergent descent method for minimization. The Computer Journal, 6(2):163–168, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/2/163.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_02/060163.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/163.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/164.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/165.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/166.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/167.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/168.tif.

Fletcher:1963:MAD

[FR63] R. Fletcher and C. M. Reeves. A mechanization of algebraic differentiation and the automatic generation of formulae for molecular integrals of Gaussian orbitals. The Computer Journal, 6(3):287–292, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/287.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060287.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
287.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/288.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/289.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/290.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/291.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/292.tif.

Fletcher:1964:FMC

[FR64] R. Fletcher and C. M. Reeves. Function minimization by conjugate gradients. The Computer Journal, 7(2):149–154, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/2/149.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_02/070149.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/150.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/

tiff/151.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/152.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/153.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/154.tif.

Francis:1961:TUA

[Fra61] J. G. F. Francis. The QR transformation: a unitary analogue to the LR transformation — part 1. The Computer Journal, 4(3):265–271, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/3/265.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_03/040265.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/265.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/266.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/267.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/268.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/270.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/271.tif.  See also

[Fra62, Wat08, Wat11].

Francis:1962:TP

[Fra62] J. G. F. Francis. The QR transformation — part 2. The Computer Journal, 4(4):332–345, January 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/4/332.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/040332.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/
332.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/333.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_04/tiff/334.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/336.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/338.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/339.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/340.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/341.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/342.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/343.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/344.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

04/Issue_04/tiff/345.tif. See also [Fra61, Wat08, Wat11].

Fraser:1967:DCA

[Fra67] A. G. Fraser. Data compression and automatic programming. The Computer Journal, 10(2):165–167, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/165.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/100165.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
165.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/166.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/167.tif.

Fraser:1968:UCM

[Fra68] A. G. Fraser. User control in a multi-access system. The Computer Journal, 11(1):12–16, May 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/12.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110012.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
12.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/13.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/14.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
15.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/16.tif.

Francis:1969:BR

[Fra69a] L. Francis. Book review. The Computer Journal, 12(2):150, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/150.full.pdf+html.

Fraser:1969:IMS

[Fra69b] A. G. Fraser. Integrity of a mass storage filing system. The Computer Journal, 12(1):1–5, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/12/1/1.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_01/120001.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/tiff/1.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
01/tiff/2.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/3.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_01/tiff/4.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_01/tiff/5.tif.

Fraser:1970:CFC

[Fra70] J. G. Fraser. Correspondence: Further comments on a linethinning scheme. The Computer Journal, 13(2):221–222, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/2/221.2.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/221.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/

Issue_02/tiff/222.tif. See correspondence [Deu69].

Frielink:1965:AED

[Fri65] A. B. Frielink. The auditor and electronic data processing. The Computer Journal, 8(2):95–97, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/8/2/95.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/080095.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/tiff/95.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
02/tiff/96.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/97.tif.

Fraser:1966:CLO

[FS66] A. G. Fraser and J. D. Smart. The COMPL language and operating system. The Computer Journal, 9(2):144–156, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/2/144.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_09/Issue_02/090144.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/144.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/145.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/147.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/150.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/151.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/152.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/153.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/154.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/155.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/156.tif.

Fuller:1966:BRBa

[Ful66a] T. Fuller. Book review: Error in Digital Computation, Volume 1. The Computer Journal, 8(4):367, January 1966. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/8/4/367.
full.pdf+html.

Fuller:1966:BRBb

[Ful66b] T. Fuller. Book review: Error in Digital Computation, Volume II. The Computer Journal, 9(3):317, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
3/317.full.pdf+html.

Fuller:1967:BR

[Ful67] A. T. Fuller. Book review. The Computer Journal, 9(4):387, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/387.full.pdf+html.

Fox:1963:C

[FY63] L. Fox and Andrew Young. Correspondence. The Computer Journal, 5(4):296, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/5/4/296.full.pdf+html.

Fyfe:1969:UCS

[Fyf69] D. J. Fyfe. The use of cubic splines in the solution of twopoint boundary value problems. The Computer Journal, 12(2): 188–192, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/188.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120188.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/188.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/189.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/190.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/191.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/192.tif.

Gatehouse:1965:EDP

[Gat65] J. S. Gatehouse. Electronic data processing for the international vocabulary of terms used in information processing. The Computer Journal, 7(4):271–274, January 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/271.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070271.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
271.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/272.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/273.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/274.tif.

Gatehouse:1967:BR

[Gat67] J. S. Gatehouse. Book reviews. The Computer Journal, 9(4): 380, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/4/380.full.pdf+html.

Grimsdale:1961:CRD

[GB61] R. L. Grimsdale and J. M. Bullingham. Character recognition by digital computer using a special flying-spot scanner.

The Computer Journal, 4(2):129–136, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/129.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040129.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
129.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/130.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/131.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/132.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/133.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/134.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/135.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/136.
tif.

Gould:1969:CUT

[GB69] I. H. Gould and R. W. Bemer. Correspondence: On the use of the term ‘processor’. The Computer Journal, 12(4):411–412, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/4/411.2.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/411.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/

Issue_04/tiff/412.tif. See [LWS68, PB69, LWS69c].

Gough:1969:GPE

[GE69] N. E. Gough and J. B. A. Epton. A general procedure for evaluating the controllability of time delay feedback control systems. The Computer Journal, 12(3):273–276, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/3/273.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/120273.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/
273.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/274.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/275.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/276.tif.

Gearing:1961:BR

[Gea61a] H. W. Gearing. Book review. The Computer Journal, 3(4):236, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/3/
4/236.full.pdf+html.

Gearing:1961:UPA

[Gea61b] H. W. Gearing. The use of Pegasus Autocode in some experimental business applications of computers. The Computer Journal, 4(1):30–34, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/1/30.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/040030.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/30.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/31.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/32.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/33.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/34.tif.

Gear:1964:OAF

[Gea64a] C. W. Gear. Optimization of the address field compilation in the IL-LIAC 2 assembler. The Computer Journal, 6(4):332–335, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/6/
4/332.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/332.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/
tiff/333.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/334.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/335.
tif.

Gearing:1964:BRB

[Gea64b] H. W. Gearing. Book review: Data processing. The Computer Journal, 7(1):22–23, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/1/22.full.pdf+html.

Gearing:1965:EMA

[Gea65] H. W. Gearing. Extended Mercury Autocode. The Computer Journal, 8(1):23, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/23.full.pdf+html.

Gearing:1967:BR

[Gea67] H. W. Gearing. Book review. The Computer Journal, 10(3): 277, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/3/277.full.pdf+html.

Gee:1966:BRB

[Gee66] R. D. Gee. Book review: On Retrieval System Theory. The Computer Journal, 9(1):97, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/1/97.full.pdf+html.

Gee:1967:BR

[Gee67] R. D. Gee. Book review. The Computer Journal, 10(1):21, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/21.full.pdf+html.

Gentleman:1969:EAG

[Gen69] W. M. Gentleman. An error analysis of Goertzel’s (Watt’s) method for computing Fourier coefficients. The Computer Journal, 12(2): 160–164, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/160.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120160.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/160.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/161.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/164.tif.

George:1963:SAJ

[Geo63] A. F. George. SAS aids for the jet age: telecommunication for electronic reservations. The Computer Journal, 6(1):14– 17, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/14.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060014.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/14.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/15.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/16.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/17.tif.

George:1966:CAI

[Geo66] F. H. George. Computer assisted instruction. The Computer Journal, 9(1):32–34, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/32.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090032.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/32.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/33.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/34.tif.

George:1968:CHM

[Geo68] F. H. George. Computers and the human mind. The Computer Journal, 11(2):212, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/2/212.full.pdf+html.

Gibbons:1960:PAI

[Gib60] A. Gibbons. A program for the automatic integration of differential equations using the method of Taylor series. The Computer Journal, 3(2):108–111, July 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/2/108.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/030108.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/
108.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/109.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_02/tiff/110.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_02/tiff/111.tif.

Gibbons:1961:RPA

[Gib61] A. Gibbons. Running Pegasus Autocode programs on Mercury. The Computer Journal, 3(4):232–236, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/232.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030232.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
232.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/233.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/234.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/236.tif.

Gibbons:1968:KA

[Gib68] A. Gibbons. K autocode. The Computer Journal, 11(2):160– 168, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/160.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/110160.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/160.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/161.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/163.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/165.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/166.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/167.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/168.tif.

Gill:1962:BR

[Gil62] S. Gill. Book reviews. The Computer Journal, 5(3):237, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
3/237.full.pdf+html.

Gill:1963:BRM

[Gil63a] S. Gill. Book review: Management. The Computer Journal, 6(2): 162, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/2/162.full.pdf+html.

Gill:1963:TCM

[Gil63b] S. Gill. A technique for the composition of music in a computer. The Computer Journal, 6(2):129–133, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/129.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060129.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
129.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/130.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/131.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_02/tiff/132.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/133.tif.

Giles:1964:PLA

[Gil64] P. Giles. Problems of life assurance data processing. The Computer Journal, 7(2):95–98, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/2/95.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/070095.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_02/tiff/95.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
02/tiff/96.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/97.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/98.
tif.

Gill:1965:ACP

[Gil65] S. Gill. Automatic computing: its problems and prizes. The Computer Journal, 8(3):177–189, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/177.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080177.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/

177.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/178.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/179.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/180.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/181.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/182.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/183.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/184.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/185.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/186.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/187.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/188.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/189.
tif.

Giles:1966:SCR

[Gil66] P. Giles. Some considerations regarding arrangement of records on magnetic tape. The Computer Journal, 8(4):319–322, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/319.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080319.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/319.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/320.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/321.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/322.tif.

Gilles:1968:BR

[Gil68] D. C. Gilles. Book review. The Computer Journal, 11(3):346, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/346.full.pdf+html.

Giles:1969:MC

[Gil69a] P. Giles. Mini-COBOL. The Computer Journal, 12(3):208– 214, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/3/208.full.pdf+html;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_12/Issue_03/120208.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/208.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/209.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/210.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/211.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/212.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/213.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/214.
tif.

Gill:1969:BR

[Gil69b] S. Gill. Book review. The Computer Journal, 12(4):341, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/341.full.pdf+html.

Garside:1968:NMS

[GJM68] G. R. Garside, P. Jarratt, and C. Mack. A new method for solving polynomial equations. The Computer Journal, 11(1):87–90, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/87.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110087.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/87.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/88.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
89.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

11/Issue_01/tiff/90.tif. See correspondence [O’B68b].

Grant:1964:NIP

[GL64] J. A. Grant and Oliver G. Ludwig. Note on the integrals of products of associated Legendre functions. The Computer Journal, 6(4):356–357, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/356.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060356.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
356.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/357.tif.

Garwick:1970:CEI

[GL70] Jan V. Garwick and J. N. Lyness. Correspondence: On “The effect of inadequate convergence criteria in automatic routines”. The Computer Journal, 13(1):122, February 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/122.tif. See [Lyn69].

Gladwin:1969:UGD

[Gla69] B. J. Gladwin. The utilisation of graphic display units as the main form of computer input. The Computer Journal, 12(2): 114–117, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/114.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120114.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/114.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/115.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/116.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/117.tif.

Glennie:1962:OEF

[Gle62] A. E. Glennie. Operating experience with Fortran. The Computer Journal, 5(2):132–134, August 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/2/132.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/050132.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
132.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/133.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/tiff/134.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_02/tiff/135.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/136.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/
tiff/137.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/138.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/139.
tif.

Gourlay:1968:DAL

[GlM68] A. R. Gourlay and J. L. l. Morris. Deferred approach to the limit in non-linear hyperbolic systems. The Computer Journal, 11(1):95– 101, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/95.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110095.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/100.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/101.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/
tiff/95.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/96.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/97.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/98.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/99.tif.

Green:1965:SSC

[GM65] D. J. Green and S. Michaelson. Series solution of certain Sturm– Liouville eigenvalue problems. The Computer Journal, 7(4):322– 336, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/4/322.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/070322.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/322.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/323.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/324.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/325.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/326.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/327.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/328.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_04/tiff/329.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/330.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/331.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/332.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/333.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

07/Issue_04/tiff/334.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/335.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/336.tif.

Golds:1964:UEC

[Gol64] I. M. Golds. User’s experience of Cobol. The Computer Journal, 7 (2):99–101, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/2/99.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/070099.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/tiff/99.tif.

Goldman:1969:CAT

[Gol69] M. S. Goldman. Computer-aided typesetting with a CRT. The Computer Journal, 12(4):357–364, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/357.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120357.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
357.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/358.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/359.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/360.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/361.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/362.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/363.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/364.
tif.

Goodman:1960:BR

[Goo60] R. Goodman. Book review. The Computer Journal, 2(4):194, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/2/
4/194.full.pdf+html.

Good:1961:CSM

[Goo61] I. J. Good. A comparison of some methods of calculating covariance functions on an electronic computer. The Computer Journal, 3(4):262–265, January 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/262.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030262.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
262.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/263.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/264.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/265.tif.

Goodstein:1963:BR

[Goo63] R. L. Goodstein. Book reviews. The Computer Journal, 6(3):297– 298, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/297.2.full.pdf+html.

Goodman:1964:AMT

[Goo64] J. V. Goodman. Auditing magnetic-tape systems. The Computer Journal, 7(1):4–7, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/7/1/4.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070004.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/4.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/5.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/6.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_01/tiff/7.tif.

Good:1965:BMM

[Goo65] I. J. Good. Brains, machines and mathematics. The Computer Journal, 8(1):88, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/88.full.pdf+html.

Good:1968:CGT

[Goo68a] I. J. Good. Correspondence: Gödel’s theorem. The Computer Journal, 10(4):417, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_10/Issue_04/tiff/417.tif.

Good:1968:GT

[Goo68b] I. J. Good. Gödel’s Theorem. The Computer Journal, 10(4): 417, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/417.1.full.pdf+html.

Gosden:1960:MRA

[Gos60] J. A. Gosden. Market research applications on LEO. The Computer Journal, 3(3):142–143, October 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/142.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030142.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
142.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/143.tif.

Gosden:1963:ECP

[Gos63] J. A. Gosden. Estimating computer performance. The Computer Journal, 5(4):276–283, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/276.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050276.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
276.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/277.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/278.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/279.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/280.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/281.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/282.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/283.
tif.

Gourlay:1968:APR

[Gou68] A. R. Gourlay. The acceleration of the Peaceman–Rachford method by Chebyshev polynomials. The Computer Journal, 10(4):378– 382, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/378.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/100378.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/378.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/379.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/380.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/381.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/382.tif.

Gower:1962:HMT

[Gow62] J. C. Gower. The handling of multiway tables on computers. The Computer Journal, 4(4):280–286, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/4/280.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/040280.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/
280.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/281.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_04/tiff/282.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/284.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/285.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/286.tif.

Grady:1967:NG

[Gra67] J. W. Grady. The National Giro. The Computer Journal, 10(2): 121–125, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/121.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/100121.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/121.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/122.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/123.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/124.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/125.tif.

Grad:1968:ACE

[GRB68] J. Grad, K. A. Redish, and M. A. Brebner. Algorithm 32: Calculation of eigenvalues of real matrices by the QR method using double QR step. The Computer Journal, 11(1):112– 114, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/112.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/
tiff/113.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/114.tif.

Greenwood:1965:BR

[Gre65] H. H. Greenwood. Book review. The Computer Journal, 8(3): 241, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/3/241.full.pdf+html.

Greenspan:1967:NAB

[Gre67] Donald Greenspan. A numerical approach to biharmonic problems. The Computer Journal, 10(2):198–201, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/198.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/100198.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
198.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/199.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/200.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/201.tif.

Greenspan:1969:NSP

[Gre69] Donald Greenspan. Numerical studies of prototype cavity flow problems. The Computer Journal, 12(1):88–93, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/88.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120088.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
90.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

12/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/92.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
93.tif.

Grimmond:1962:ARS

[Gri62] R. Grimmond. An analysis of real and simulated statistics for system design purposes. The Computer Journal, 5(2):94– 99, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/2/94.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/050094.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/94.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/tiff/95.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
02/tiff/96.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/97.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/98.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/99.tif.

Grindley:1966:SNP

[Gri66] C. B. B. Grindley. Systematics — a non-programming language for designing and specifying commercial systems for computers. The Computer Journal, 9(2):124–128, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/124.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090124.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
124.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/125.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/126.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/127.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/128.tif. See com-

ments [BBdD66].

Grindley:1968:UDT

[Gri68] C. B. B. Grindley. The use of decision tables within systematics. The Computer Journal, 11(2):128–133, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/2/128.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/110128.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/
128.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/129.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/130.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/131.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/132.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/133.tif.

Groves:1968:MEC

[Gro68] P. D. Groves. Marking and evaluating class tests and examinations by computer. The Computer Journal, 10(4):365–367, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
4/365.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/100365.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/tiff/365.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/366.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/367.
tif.

Gerard:1963:HRA

[GS63] J. M. Gerard and A. Sambles. A hardware representation for ALGOL 60 using Creed teleprinter equipment. The Computer Journal, 5(4):338–340, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/338.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050338.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
338.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/339.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/340.tif.

Goyer:1963:BR

[GT63a] M. Goyer and G. C. Tootill. Book review. The Computer Journal, 6(3):273, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/6/3/273.full.pdf+html.

Goyer:1963:BRB

[GT63b] M. Goyer and G. C. Tootill. Boor review: Russian Supplement to Elsevier’s Dictionary of Automation, Computers, Control and Measuring, by W. E. Clason, 1962. The Computer Journal, 6(3): 273, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/273.tif.

Gould:1965:TWI

[GT65] I. H. Gould and G. C. Tootill. The terminology work of IFIP and ICC. The Computer Journal, 7(4):264–270, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/264.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070264.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
264.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/265.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/266.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/268.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/269.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/270.tif.

Guin:1968:DCM

[Gui68] J. A. Guin. Discussion and correspondence: modification of the complex method of constrained optimization. The Computer Journal, 10(4):416–417, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/416.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100416.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
416.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/417.tif.

HWG:1968:MIS

[H. 68] H. W. G. Management information systems. The Computer Journal, 10(4):418, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/10/4/418.full.pdf+

html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/418.tif.

Hadjidimos:1968:GAD

[Had68] A. Hadjidimos. On a generalised alternating direction implicit method for solving Laplace’s equation. The Computer Journal, 11(3):324–328, November 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/324.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110324.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
324.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/325.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/326.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/327.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/328.tif.

Hagan:1969:BRB

[Hag69] T. Hagan. Book review: Mathematical Linguistics in Eastern Europe. The Computer Journal, 12(1):32, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
1/32.full.pdf+html.

Haines:1969:IIP

[Hai69] C. F. Haines. Implicit integration processes with error estimate for the numerical solution of differential equations. The Computer Journal, 12(2):183–187, May 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/2/183.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/120183.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/
183.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/184.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/185.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/186.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/187.tif.

Haley:1966:PGT

[Hal66] K. B. Haley. Programming, games and transportation networks. The Computer Journal, 9(2):166, August 1966. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/9/2/166.
full.pdf+html.

Hall:1967:SPC

[Hal67] G. Hall. The stability of predictor-corrector methods. The Computer Journal, 9(4):410–412, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/410.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090410.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
410.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/411.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/412.tif.

Hall:1968:PCI

[Hal68] P. A. V. Hall. Pattern classification as interpolation in N dimensions. The Computer Journal, 11(3):287–292, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/287.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/110287.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/tiff/287.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/288.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/289.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/290.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/291.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/292.tif.

Hamblin:1962:TPN

[Ham62] C. L. Hamblin. Translation to and from Polish notation. The Computer Journal, 5(3):210–213, November 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/210.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/050210.sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/
210.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/211.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_03/tiff/212.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_03/tiff/213.tif.

Hammersley:1966:AS

[Ham66a] P. Hammersley. Algorithms supplement. The Computer Journal, 9(2):211–212, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/2/211.full.pdf+html.

Hammersley:1966:NIL

[Ham66b] P. Hammersley. A note on the implementation of LITHP on the I.C.T. 1905. The Computer Journal, 9(2):173–174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/2/173.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_02/090173.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/173.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/174.tif.

Hammersley:1966:NFP

[Ham66c] P. Hammersley. Notice: Future publication of Algorithms Supplement. The Computer Journal, 9(1):80, May 1966. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/80.tif.

Hammond:1966:BRB

[Ham66d] P. H. Hammond. Book review: Modern Control Theory. The Computer Journal, 8(4):322, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/4/322.full.pdf+html.

Hamblin:1967:APO

[Ham67a] C. L. Hamblin. An algorithm for polynomial operations. The Computer Journal, 10(2):168–171, August 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/168.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_10/Issue_02/100168.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
168.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/169.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/170.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/171.tif.

Hammersley:1967:ASPa

[Ham67b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 9(4):414, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/414.tif.

Hammersley:1967:ASPb

[Ham67c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(1):112, May 1967. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/112.tif.

Hammersley:1967:ASPc

[Ham67d] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(2):206, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/206.tif.

Hammersley:1967:ASPd

[Ham67e] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(3):306, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/306.tif.

Hammersley:1967:SP

[Ham67f] P. Hammersley. Statement of policy. The Computer Journal, 10(2): 209–210, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/209.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/210.tif.

Hammersley:1968:ASPa

[Ham68a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 10(4):414, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/414.tif.

Hammersley:1968:ASPb

[Ham68b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 11(1):112, May 1968. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/112.tif.

Hammersley:1968:ASPc

[Ham68c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 11(3):347, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/347.tif.

Hammersley:1969:ASPa

[Ham69a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 12(2):197, May 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_02/tiff/197.tif.

Hammersley:1969:ASPb

[Ham69b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 12(3):291, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/291.tif.

Hammersley:1969:ASPc

[Ham69c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 12(4):405–406, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/405.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/406.
tif.

Hammersley:1969:NAP

[Ham69d] P. Hammersley. Note on Algorithm 34: Procedures for the basic arithmetical operations in multi-length working. The Computer Journal, 12(1):102–103, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/
tiff/102.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/103.tif. See [Boo67e].

Handscomb:1962:CLR

[Han62] D. C. Handscomb. Computation of the latent roots of a Hessenberg matrix by Bairstow’s method. The Computer Journal, 5(2):139– 141, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/2/139.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/050139.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/139.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/tiff/140.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_02/tiff/141.tif.

Harrison:1963:BR

[Har63] John D. Harrison. Book review. The Computer Journal, 5(4): 293, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/4/293.full.pdf+html.

Harding:1965:BRB

[Har65a] C. E. Harding. Book review: Management Standards for Data Processing. The Computer Journal, 7(4):254, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
4/254.full.pdf+html.

Harding:1965:MSD

[Har65b] C. E. Harding. Management standards for data processing. The Computer Journal, 7(4):254, January 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/254.full.
pdf+html.

Haselgrove:1961:SNL

[Has61] C. B. Haselgrove. The solution of non-linear equations and of differential equations with two-point boundary conditions. The Computer Journal, 4(3):255–259, October 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/3/255.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/040255.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/
255.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/256.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_03/tiff/257.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_03/tiff/258.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/259.tif.

Hayes:1969:MSO

[Hay69] D. G. Hayes. A method of storing the orthogonal polynomials used for curve and surface fitting. The Computer Journal, 12(2): 148–150, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/148.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120148.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/148.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/149.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/150.tif.

Herzenstein:1961:C

[HB61] Z. Herzenstein and F. E. A. Briggs. Correspondence. The Computer Journal, 4(2):128, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/4/2/128.full.pdf+html.

Hughes:1966:DSA

[HB66] F. M. Hughes and A. Brameller. Digital simulation of analogue methods. The Computer Journal, 9(1):35–44, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/1/35.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/090035.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

09/Issue_01/tiff/35.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/36.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
37.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/38.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/39.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
40.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/41.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/42.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
43.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/44.tif.

Hopgood:1967:AAP

[HB67] F. R. A. Hopgood and A. G. Bell. The Atlas ALGOL preprocessor for non-standard dialects. The Computer Journal, 9(4):360–364, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/360.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090360.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/360.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/361.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/362.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/363.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/364.
tif.

Healy:1961:BR

[Hea61] M. J. R. Healy. Book review. The Computer Journal, 4(1):24, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
1/24.1.full.pdf+html.

Healy:1963:PMR

[Hea63a] M. J. R. Healy. Programming multiple regression. The Computer Journal, 6(1):57–61, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/6/1/57.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060057.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/57.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/58.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/59.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/60.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/61.tif.

Heap:1963:PI

[Hea63b] B. R. Heap. Permutations by interchanges. The Computer Journal, 6(3):293–294, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/293.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060293.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
293.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/294.tif.

Healy:1968:TFV

[Hea68] M. J. R. Healy. Towards Fortran vi? The Computer Journal, 11(2): 169–172, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/169.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/110169.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/169.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/170.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/171.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/172.tif.

Herring:1961:BR

[Her61a] G. J. Herring. Book review. The Computer Journal, 4(2):149, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
2/149.full.pdf+html.

Herring:1961:BRB

[Her61b] G. J. Herring. Book review: Analogue Computation in Engineering Design, by A. E. Rogers and T. W. Connolly, 1960. The Computer Journal, 4(1):41, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_04/Issue_01/tiff/41.tif.

Hext:1967:CTT

[Hex67] J. B. Hext. Compile-time type-matching. The Computer Journal, 9(4):365–369, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/4/365.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
04/090365.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/365.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/366.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/367.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/368.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/369.tif.

Huesmann:1967:ECS

[HG67] L. Rowell Huesmann and Robert P. Goldberg. Evaluating computer systems through simulation. The Computer Journal, 10(2): 150–156, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/150.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/100150.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/150.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/151.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/152.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/153.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/154.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/155.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/156.
tif.

Hickman:1960:EED

[Hic60] T. C. Hickman. Early experiences with an E.D.P. system. The Computer Journal, 2(4):152–163, January 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/4/152.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/020152.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/
152.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_02/Issue_04/tiff/153.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_04/tiff/154.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/
Issue_04/tiff/155.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_04/tiff/156.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/
tiff/157.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_04/tiff/158.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/159.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_04/tiff/160.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_04/tiff/161.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/
tiff/162.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_04/tiff/163.tif.

Higman:1963:WES

[Hig63] B. Higman. What EVERYBODY should know about ALGOL. The Computer Journal, 6(1):50–56, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/6/1/50.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060050.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/50.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/51.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/52.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/53.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/55.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
56.tif.

Higman:1965:C

[Hig65] B. Higman. Correspondence. The Computer Journal, 8(2):175, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
2/175.1.full.pdf+html.

Higman:1967:CPS

[Hig67] Bryan Higman. Correspondence: Paging and segmentation. The Computer Journal, 10(2):216, August 1967. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/216.tif.

Hill:1967:CEC

[Hil67] I. D. Hill. Correspondence: On “Efficiency of computer use should not be a consideration”. The Computer Journal, 10(2):209, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/209.tif.

Hill:1969:NAS

[Hil69] I. D. Hill. Note on Algorithms 25, 26. 25: Sort a section of the elements of an array by determining the rank of each element. 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 12(4):409, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/409.tif. See [Boo67f, Boo67g,

Woo70].

Hindle:1960:BAA

[Hin60] R. Hindle. A bank adopts automatic data processing. The Computer Journal, 3(3):127–130, October 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/127.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030127.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
127.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/128.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_03/tiff/129.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_03/tiff/130.tif.

Hindle:1961:CRD

[Hin61] R. Hindle. Character recognition and document handling in banks. The Computer Journal, 4(2):157–160, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/157.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040157.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/

157.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/158.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/159.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/160.tif.

Hiorns:1967:SAF

[Hio67] R. W. Hiorns. Seasonal adjustment and forecasting in the presence of trend. The Computer Journal, 10(2):143–149, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
2/143.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/100143.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
02/tiff/143.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/144.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_02/tiff/146.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/147.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/148.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/149.tif.

Hiorns:1968:BR

[Hio68a] R. W. Hiorns. Book review. The Computer Journal, 11(1):71, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/71.full.pdf+html.

Hiorns:1968:MFA

[Hio68b] R. W. Hiorns. Modern factor analysis. The Computer Journal, 11 (2):219, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/219.full.pdf+html.

Hill:1967:NAN

[HJBH67] I. D. Hill, S. A. Joyce, A. Bergson, and P. Hammersley. Note on Algorithm 13: normalarea. The Computer Journal, 10(1): 113–115, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/113.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/

tiff/114.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/115.tif.

Howarth:1962:ASS

[HJW62] D. J. Howarth, P. D. Jones, and M. T. Wyld. The Atlas scheduling system. The Computer Journal, 5(3):238–244, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/238.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_03/050238.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/239.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/241.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/242.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/243.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/244.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/246.tif.

Harris:1967:CAL

[HK67] D. J. Harris and A. K. Kent. The computer as an aid to lawyers. The Computer Journal, 10(1):22–28, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/22.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100022.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
22.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/23.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/24.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
25.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/26.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/27.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
28.tif.

Hammersley:1970:CTF

[HL70] P. Hammersley and J. Larmouth. Correspondence: Towards Fortran VI? The Computer Journal, 13(2):220, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/220.tif. See [Red71, Wel70].

Hartley:1968:SMS

[HLN68] D. F. Hartley, B. Landy, and R. M. Needham. The structure of a multiprogramming supervisor. The Computer Journal, 11(3):247– 255, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/247.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/110247.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/247.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/248.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/249.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/250.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/251.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/252.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/253.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/255.tif.

Higman:1965:CIP

[HLOS65] Bryan Higman, W. F. Lunnon, C. F. J. Outred, and C. Strachey. Correspondence: An impossible program. The Computer Journal, 8(2):175–176, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/175.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/176.tif. See [Str65e, ApS65,

BSPI65, AMKM66].

Hoare:1962:Q

[Hoa62a] C. A. R. Hoare. Quicksort. The Computer Journal, 5(1):10– 16, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/1/10.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_01/050010.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_05/Issue_01/tiff/10.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/11.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/14.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/15.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/16.tif.

Hoare:1962:REA

[Hoa62b] C. A. R. Hoare. Report on the Elliott Algol translator. The Computer Journal, 5(2):127–129, August 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/2/127.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/050127.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
127.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/128.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/tiff/129.tif.

Hoare:1963:BR

[Hoa63a] C. A. R. Hoare. Book reviews. The Computer Journal, 6(2): 143, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/2/143.full.pdf+html.

Hoare:1963:EAI

[Hoa63b] C. A. R. Hoare. The Elliott ALGOL input/output system. The Computer Journal, 5(4):345–348, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/345.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050345.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
345.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/346.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/347.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/348.tif.

Hoare:1964:BRB

[Hoa64] C. A. R. Hoare. Book review: Algol on the KDF9. The Computer Journal, 7(2):154, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/154.full.pdf+html.

Hockney:1962:AAE

[Hoc62] R. W. Hockney. ABS12 ALGOL: an extension to Algol 60 for industrial use. The Computer Journal, 4(4):292–300, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/4/292.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_04/040292.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/292.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/293.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/294.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/295.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/296.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/297.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/298.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/299.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/300.tif.

Hope:1961:SRT

[Hop61] K. S. Hope. SABER: a real time problem in tele-processing. The Computer Journal, 4(2):109–113, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/109.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040109.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
109.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/110.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/111.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/112.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/113.tif.

Hopkin:1962:BR

[Hop62] H. R. Hopkin. Book reviews. The Computer Journal, 5(2):154, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/2/154.full.pdf+html.

Hopkinson:1963:IAU

[Hop63] J. R. Hopkinson. Integrated accounting using a variety of equipment. The Computer Journal, 6(1):5–6, April 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/6/1/5.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060005.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/5.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/6.tif.

Hornsby:1960:FIS

[Hor60] J. S. Hornsby. A function interpretive scheme for Pegasus. The Computer Journal, 2(4):174–180, January 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/4/174.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/020174.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/
174.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/tiff/175.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_04/tiff/176.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/
Issue_04/tiff/177.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_04/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/
tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_04/tiff/180.tif.

Houstoun:1963:BR

[Hou63a] J. Houstoun. Book review. The Computer Journal, 6(3):247, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/247.full.pdf+html.

Houstoun:1963:BRB

[Hou63b] J. Houstoun. Book review: Mathematical Methods for the Study of Automatic Control Systems, by V. I. Zubov, 1962. The Computer Journal, 6(3):247, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/247.
tif.

Housden:1969:DIL

[Hou69a] R. J. W. Housden. The definition and implementation of Lsix in BCL. The Computer Journal, 12(1):15–23, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/15.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120015.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
15.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/16.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/17.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
18.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
21.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/23.tif.

Housden:1969:NLP

[Hou69b] R. J. W. Housden. A note on list-processing in BCL. The Computer Journal, 12(4):332–341, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/332.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120332.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
332.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/333.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/334.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/336.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/

tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/338.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/339.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_04/tiff/340.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/341.tif.

Howlett:1962:BR

[How62] J. Howlett. Book reviews. The Computer Journal, 5(2):151, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
2/151.full.pdf+html.

Howlett:1965:BR

[How65] J. Howlett. Book review. The Computer Journal, 8(3):287, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
3/287.full.pdf+html.

Howden:1967:PCP

[How67a] W. E. Howden. A program for the construction of PERT flow charts. The Computer Journal, 10(3):278–281, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/278.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/100278.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
03/tiff/278.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/279.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/280.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/281.tif.

Howell:1967:MPA

[How67b] Keith M. Howell. Multiple precision arithmetic techniques. The Computer Journal, 9(4):383–387, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/383.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090383.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
383.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/384.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_09/Issue_04/tiff/385.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/386.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/387.tif.

Howden:1968:SP

[How68a] W. E. Howden. The sofa problem. The Computer Journal, 11(3):299–301, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/3/299.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/110299.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/299.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/301.tif.

Howlett:1968:BR

[How68b] J. Howlett. Book review. The Computer Journal, 11(1):34, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/34.full.pdf+html.

Hill:1967:NAT

[HP67] I. D. Hill and M. C. Pike. Note on algorithm 4: TWO BY TWO. The Computer Journal, 9(4):416, February 1967. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/416.tif. See [HP79].

Hill:1979:AT

[HP79] I. D. Hill and M. C. Pike. Algorithm 4: TWOBYTWO. The Computer Journal, 22(1):87–88, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/
tiff/87.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/88.tif. Reprinted from The

Computer Bulletin, September 1964, p. 67. See note [HP67].

Howarth:1961:MUA

[HPS61] D. J. Howarth, R. B. Payne, and F. H. Sumner. The Manchester University Atlas operating system part II: users’ description. The Computer Journal, 4(3):226–229, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/3/226.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/040226.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/
226.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/227.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_03/tiff/228.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_03/tiff/229.tif.

Hunt:1968:CSF

[HRS68] C. H. Hunt, M. M. Ryder, and S. H. Storey. Computers and the small firm: 1. The Computer Journal, 11(1):1– 4, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/1/1.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_01/110001.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/1.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/tiff/2.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
01/tiff/3.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/4.tif.

Hendry:1969:TFV

[HS69] D. F. Hendry and P. A. Samet. Towards FORTRAN VI? Part 2. FORTRAN in the modern world. The Computer Journal, 12(3): 218–220, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/3/218.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/120218.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/218.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/219.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/220.tif.

Humby:1966:CCE

[HSP66] E. Humby, P. Shackleton, and R. M. Paine. Correspondence: Commercial English languages. The Computer Journal, 8(4):310–311, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/310.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_08/Issue_04/tiff/310.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
04/tiff/311.tif.

Hughes:1961:CPC

[Hug61] D. J. L. Hughes. Computer production control — the second year. The Computer Journal, 3(4):198–201, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/198.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030198.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
198.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/199.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/200.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/201.tif.

Hughes:1967:BR

[Hug67] P. A. B. Hughes. Book review. The Computer Journal, 9(4): 382, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/4/382.full.pdf+html.

Humby:1962:RNA

[Hum62] E. Humby. RAPIDWRITE — a new approach to COBOL readability. The Computer Journal, 4(4):301–304, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/4/301.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/040301.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/
301.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/302.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_04/tiff/303.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/304.tif.

Humpidge:1967:BR

[Hum67] H. B. Humpidge. Book review. The Computer Journal, 9(4): 372, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/4/372.1.full.pdf+html.

Hunter:1960:NTR

[Hun60] D. G. N. Hunter. Note on a test for repeating cycles in a pseudo-random number generator. The Computer Journal, 3 (1):9, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/1/9.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_01/030009.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/9.tif.

Hunter:1961:IMN

[Hun61] D. B. Hunter. An iterative method of numerical differentiation. The Computer Journal, 3(4):270–271, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/270.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030270.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
270.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/271.tif.

Hunter:1968:NMT

[Hun68] D. B. Hunter. Neville’s method for trigonometric interpolation. The Computer Journal, 11(3):311–313, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/311.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110311.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
311.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/312.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/313.tif.

Hurt:1965:LEP

[Hur65a] Robert Hurt. Letter to the Editor: Program development with FACT. The Computer Journal, 7(4):270, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
4/270.full.pdf+html.

Hurt:1965:PDF

[Hur65b] Robert Hurt. Program development with fact. The Computer Journal, 7(4):270, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/4/270.full.pdf+html.

Hurst:1968:EBM

[Hur68] S. L. Hurst. An extension of binary minimisation techniques to ternary equations. The Computer Journal, 11(3):277–286, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/277.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/110277.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/277.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/278.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/279.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/280.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/281.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/282.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/283.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/284.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/285.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/286.tif.

Huskey:1961:CTA

[Hus61] H. D. Huskey. Compiling techniques for algebraic expressions. The Computer Journal, 4(1):10–19, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/1/10.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/040010.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/13.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/14.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/15.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/
16.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/17.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_04/Issue_01/tiff/18.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/
19.tif.

Haddon:1967:CPV

[HW67] B. K. Haddon and W. M. Waite. A compaction procedure for variable-length storage elements. The Computer Journal, 10(2): 162–165, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/162.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/100162.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/162.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/163.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/164.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/165.tif.

Iliffe:1962:DSA

[IJ62] J. K. Iliffe and Jane G. Jodeit. A dynamic storage allocation scheme. The Computer Journal, 5(3):200–209, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/200.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_03/050200.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/201.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/203.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/204.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/205.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/206.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/207.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/208.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/209.
tif.

Iliffe:1964:BRC

[Ili64] J. K. Iliffe. Book review: Computer organization. The Computer Journal, 7(2):107, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/107.full.pdf+html.

Iliffe:1969:EBB

[Ili69] J. K. Iliffe. Elements of BLM (Basic Language Machine). The Computer Journal, 12(3):251–258, August 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/3/251.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/120251.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/
251.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/252.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/253.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/254.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/255.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/256.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/257.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/258.
tif.

James:1966:GGI

[Jam66] E. B. James. God and Golem, Inc. The Computer Journal, 9(2):180a–180, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/2/180-a.full.pdf+html.

Janes:1963:MPC

[Jan63] J. D. W. Janes. Measuring the profitability of a computer system. The Computer Journal, 5(4):284–293, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/284.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050284.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
284.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/285.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/286.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/287.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/288.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/289.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/290.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/291.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/292.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/293.tif.

Jarratt:1966:MIM

[Jar66a] P. Jarratt. Multipoint iterative methods for solving certain equations. The Computer Journal, 8(4):398–400, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/398.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080398.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
398.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/399.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/400.tif.

Jarratt:1966:RIF

[Jar66b] P. Jarratt. A rational iteration function for solving equations. The Computer Journal, 9(3):304–307, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/304.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090304.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
304.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/305.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/306.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/307.tif.

Jarratt:1967:IML

[Jar67a] P. Jarratt. An iterative method for locating turning points. The Computer Journal, 10(1):82–84, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/82.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100082.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
82.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

10/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/84.tif.

Jarratt:1967:NAE

[Jar67b] P. Jarratt. A note on the asymptotic error constant of a certain method for solving equations. The Computer Journal, 9(4):408–409, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/408.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090408.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/408.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/409.tif.

Jarratt:1968:UCS

[Jar68] P. Jarratt. The use of comparison series in analysing iteration functions. The Computer Journal, 11(3):314–316, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/314.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/110314.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/tiff/314.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/315.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/316.
tif.

Jennings:1961:DOR

[Jen61] G. Jennings. A direct ordering, recording and invoicing system. The Computer Journal, 4(2):150–156, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/150.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040150.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
150.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/151.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/152.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/153.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/154.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/155.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/156.tif.

Jennings:1966:CSS

[Jen66] Alan Jennings. A compact storage scheme for the solution of symmetric linear simultaneous equations. The Computer Journal, 9(3):281–285, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/281.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090281.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
281.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/282.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/283.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/284.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/285.tif.

Jones:1967:CAC

[JJ67] K. Sparck Jones and D. Jackson. Current approaches to classification and clump-finding at the Cambridge Language Research Unit. The Computer Journal, 10(1):29–37, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/29.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100029.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
29.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/30.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/31.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
32.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/34.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
35.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/37.tif.

Jarratt:1968:LSM

[JM68] P. Jarratt and C. Mack. A least squares method for Laplace’s equation with Dirichlet boundary conditions. The Computer Journal, 11 (1):83–87, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/1/83.full.pdf+html;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_11/Issue_01/110083.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/tiff/84.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
01/tiff/85.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/86.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/87.
tif.

Jarratt:1965:URF

[JN65] P. Jarratt and D. Nudds. The use of rational functions in the iterative solution of equations on a digital computer. The Computer Journal, 8(1):62–65, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/8/1/62.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_01/080062.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/62.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/63.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/65.
tif.

Johnson:1961:OTM

[Joh61] H. H. Johnson. Optimum time for multiplication on a digital computer. The Computer Journal, 3(4):256–261, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/4/256.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_04/030256.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/256.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/257.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/258.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/259.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/260.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/261.tif.

Johnston:1966:NSP

[Joh66] H. C. Johnston. Numerical solution of partial differential equations. The Computer Journal, 9(2):204, August 1966. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/9/2/204.
full.pdf+html.

Jones:1964:BR

[Jon64] Karen S. Jones. Book review. The Computer Journal, 7(3):172, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
3/172.full.pdf+html.

Jones:1965:SPC

[Jon65a] C. B. Jones. A special-purpose compiler. The Computer Journal, 8 (2):109–112, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/2/109.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/080109.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/109.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/tiff/110.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/111.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/112.tif.

Jones:1965:BR

[Jon65b] K. S. Jones. Book review. The Computer Journal, 8(3):255, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
3/255.full.pdf+html.

Jones:1966:NAM

[Jon66] Eric L. Jones. Note on an alternate method for the computation of rotational energy levels of rigid asymmetric top molecules. The Computer Journal, 9(1):65–66, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/65.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090065.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/65.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_

01/tiff/66.tif. See [Rac65] and correspondence [Duk66b].

Jones:1969:BRB

[Jon69] D. S. Jones. Book review: Field Computation by Moment Methods. The Computer Journal, 12(1):37, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
1/37.full.pdf+html.

Jackson:1969:DAC

[JP69] K. Jackson and J. R. Prior. Debugging and assessment of control programs for an automatic radar. The Computer Journal, 12(4):303–306, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/303.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120303.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
303.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/304.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/305.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/306.tif.

Jain:1967:NSL

[JS67] M. K. Jain and K. D. Sharma. Numerical solution of linear differential equations and Volterra’s integral equation using Lobatto quadrature formula. The Computer Journal, 10(1):101–107, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/101.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/100101.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
01/tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/102.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/105.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/
tiff/106.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/107.tif.

Jardine:1968:CHN

[JS68] N. Jardine and R. Sibson. The construction of hierarchic and non-hierarchic classifications. The Computer Journal, 11(2):177– 184, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/177.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/110177.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/177.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/178.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/179.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/180.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/181.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/182.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/183.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/184.tif.

Karim:1966:SFO

[Kar66] Abbas I. Abdel Karim. Stability of the fourth order Runge–Kutta method for the solution of systems of differential equations. The Computer Journal, 9(3):308–311, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/308.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090308.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
308.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/309.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/310.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/311.tif.

Kelley:1962:MIP

[KB62] D. H. Kelley and J. N. Buxton. Montecode — an interpretive program for Monte Carlo simulations. The Computer Journal, 5(2):88–93, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/5/2/88.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_02/050088.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_02/tiff/88.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
02/tiff/89.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_05/Issue_02/tiff/90.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/91.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_02/tiff/92.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/93.tif.

Kendall:1965:BR

[Ken65] M. G. Kendall. Book review. The Computer Journal, 8(3):189, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
3/189.1.full.pdf+html.

Kermode:1963:C

[Ker63a] E. R. Kermode. Correspondence. The Computer Journal, 6(3): 236, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/236.full.pdf+html.

Kermode:1963:CDC

[Ker63b] E. R. Kermode. Correspondence: On “Direct coding of English language names”. The Computer Journal, 6(3):236, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/236.tif. See [Bra63b].

Khabaza:1963:ILS

[Kha63] I. M. Khabaza. An iterative least-square method suitable for solving large sparse matrices. The Computer Journal, 6(2):202–206, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/2/202.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_02/060202.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/202.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/203.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/205.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/206.
tif.

Khabaza:1966:BDE

[Kha66] I. M. Khabaza. Journal of Differential Equations. The Computer Journal, 8(4):343, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/4/343.full.pdf+html.

Kilburn:1961:MUA

[KHPS61] T. Kilburn, D. J. Howarth, R. B. Payne, and F. H. Sumner. The Manchester University Atlas operating system part I: internal organization. The Computer Journal, 4(3):222–225, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/3/222.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_03/040222.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/222.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/223.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/224.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/225.tif.

Kingsbury:1962:OEC

[Kin62] M. A. Kingsbury. Operating experience with COBOL in a service bureau. The Computer Journal, 5(3):157–158, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/157.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_03/050157.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/158.tif.

King:1967:DT

[Kin67a] P. J. H. King. Decision tables. The Computer Journal, 10(2):135– 142, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/135.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/100135.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/135.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/136.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/137.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/138.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/139.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_10/Issue_02/tiff/140.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/141.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_02/tiff/142.tif.

King:1967:DCS

[Kin67b] P. J. H. King. Discussion and correspondence: some comments on systematics. The Computer Journal, 10(1):116–119, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/116.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/100116.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
01/tiff/116.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/117.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/118.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/119.tif.

King:1969:ILE

[Kin69a] P. J. H. King. The interpretation of limited entry decision table format and relationships among conditions. The Computer Journal, 12(4):320–326, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/320.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120320.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
320.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/321.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/322.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/323.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/324.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/325.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/326.tif. See correspondence

[Car70, Kin71].

King:1969:SAD

[Kin69b] P. J. H. King. Systems analysis documentation: Computeraided data dictionary definition. The Computer Journal, 12(1): 6–9, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/1/6.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_01/120006.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/6.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/tiff/7.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
01/tiff/8.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/9.tif.

King:1971:CIL

[Kin71] P. J. H. King. Correspondence: Interpretation of limited entry decision table format. The Computer Journal, 14(1):54, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/140054.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/

54.tif. See [Car70, Kin69a].

Kirkby:1963:BR

[Kir63] S. Kirkby. Book review. The Computer Journal, 6(1):23, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/6/
1/23.full.pdf+html.

Kizner:1966:CEC

[Kiz66a] W. Kizner. Correspondence: “Error curves for Lanczos’ ‘selected points’ method”. The Computer Journal, 9(3):324, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/

3/324.full.pdf+html. See [Kiz66b, Wri66a].

Kizner:1966:ECL

[Kiz66b] W. Kizner. Error curves for Lanczos’ “Selected Points” method. The Computer Journal, 8(4):372–382, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/372.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080372.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
372.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/373.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/374.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_04/tiff/375.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/376.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/378.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/379.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/380.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/381.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/

tiff/382.tif. See comments [Wri66a, Kiz66a].

King:1968:C

[KK68] P. J. H. King and P. R. King. Correspondence. The Computer Journal, 11(3):352, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/3/352.2.full.pdf+html.

Klerer:1965:UOP

[KM65] M. Klerer and J. May. A user oriented programming language. The Computer Journal, 8(2):103–109, July 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/2/103.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/080103.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/
103.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/104.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/tiff/105.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/106.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/107.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/108.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/109.tif.

Keast:1966:ICN

[KM66] P. Keast and A. R. Mitchell. On the instability of the Crank Nicolson formula under derivative boundary conditions. The Computer Journal, 9(1):110–114, May 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/1/110.full.

pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_01/090110.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/
110.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/111.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_01/tiff/112.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/tiff/113.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/114.tif. See [PC64]

and correspondence [Ros66a].

Knight:1968:ASS

[Kni68] D. C. Knight. An algorithm for scheduling storage on a non-paged computer. The Computer Journal, 11(1):17–21, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/17.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110017.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
17.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/18.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/19.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
20.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/21.tif.

King:1960:AHE

[KP60] P. F. King and D. A. Peel. An analysis of a hydro-electric system. The Computer Journal, 3(3):161–163, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/161.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030161.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
161.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/162.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_03/tiff/163.tif.

Kershaw:1962:EDE

[KV62] D. Kershaw and S. Vajda. The economics of dumping from electronic computers. The Computer Journal, 4(4):346–348, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/

content/4/4/346.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_04/040346.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_04/tiff/346.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/347.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/348.tif.

Kitz:1963:PSE

[KV63] Beryl Kitz and S. Vajda. A program to study the effect of random delays on the ability of trains to run to a schedule. The Computer Journal, 6(2):121–128, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/121.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060121.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
121.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/122.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/123.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_02/tiff/124.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/125.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/127.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/128.
tif.

Lambourn:1963:RAM

[Lam63] S. Lambourn. Resource allocation and multi-project scheduling (RAMPS) — a new tool in planning and control. The Computer Journal, 5(4):300–304, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/300.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050300.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
300.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/301.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/302.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/303.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/304.tif.

Landin:1964:MEE

[Lan64] Peter J. Landin. The mechanical evaluation of expressions. The Computer Journal, 6(4):308–320, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/308.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060308.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
308.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/309.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/310.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_04/tiff/311.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/312.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/
tiff/313.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/314.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/315.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/316.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/317.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/
tiff/318.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/319.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/320.
tif.

Langham:1965:USC

[Lan65] E. J. Langham. The underground storage of compressed air for gas turbines: a dynamic study on an analogue computer. The Computer Journal, 8(3):216–224, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/216.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080216.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
216.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/217.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/218.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/219.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/220.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/

tiff/221.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/222.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/223.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/224.tif.

Landin:1968:BR

[Lan68] P. J. Landin. Book review. The Computer Journal, 11(3):286, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/286.full.pdf+html.

Larkin:1964:CGI

[Lar64] F. M. Larkin. A combined graphical and iterative approach to the problem of finding zeros of functions in the complex plane. The Computer Journal, 7(3):212–219, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/212.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070212.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
212.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/213.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/214.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/215.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/216.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_03/tiff/218.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/219.
tif.

Larkin:1967:STR

[Lar67] F. M. Larkin. Some techniques for rational interpolation. The Computer Journal, 10(2):178–187, August 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/178.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/100178.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
178.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/179.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/180.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/181.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/182.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/183.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/184.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/185.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_02/tiff/186.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/187.tif.

Laski:1968:SVA

[Las68] J. G. Laski. Segmentation and virtual address topology — an essay in virtual research. The Computer Journal, 11(1):35–40, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/35.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110035.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/35.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/36.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
37.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/38.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/39.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
40.tif.

Laver:1964:UVD

[Lav64] F. J. M. Laver. A user’s view of A.D.P. systems design. The Computer Journal, 7(2):93–94, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/2/93.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/070093.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_02/tiff/93.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
02/tiff/94.tif.

Laver:1968:BR

[Lav68] F. J. M. Laver. Book review. The Computer Journal, 10(4): 335, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/335.full.pdf+html.

Lawrie:1969:ILP

[Law69] N. L. Lawrie. An integer linear programming model of a school timetabling problem. The Computer Journal, 12(4):307–316, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/307.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/120307.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/307.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/308.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/309.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/310.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/311.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/312.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/313.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_04/tiff/314.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/315.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/316.tif.

Lay:1965:BR

[Lay65] E. C. Lay. Book review. The Computer Journal, 8(2):119, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
2/119.full.pdf+html.

LeRiche:1969:ACP

[Le 69] P. J. Le Riche. Algorithm 41: a curve plotting procedure. The Computer Journal, 12(3):291–292, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/291.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/292.tif.

Leighton:1962:CGS

[Lei62] C. C. Leighton. Control gear simulation for an automatic car park. The Computer Journal, 4(4):313–316, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/4/313.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_04/Issue_04/040313.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/
313.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/314.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_04/tiff/315.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/316.tif.

Lehman:1963:CCL

[LEN63] M. Lehman, Rayna Eshed, and Z. Netter. The checking of computer logic by simulation on a computer. The Computer Journal, 6(2):154–162, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/154.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060154.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
154.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/155.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/156.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_02/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/158.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/160.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/161.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/162.tif.

Lerner:1969:FMD

[Ler69] A. J. Lerner. Formal methods of diagnosis in engineering and medicine. The Computer Journal, 12(1):29–32, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/29.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120029.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/29.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/30.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
31.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/32.tif.

Leveson:1963:BR

[Lev63] J. H. Leveson. Book review. The Computer Journal, 6(1):61, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/6/
1/61.full.pdf+html.

Levison:1965:SF

[Lev65] Michael Levison. The siting of fragments. The Computer Journal, 7(4):275–277, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/4/275.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
04/070275.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/275.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/276.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/277.tif.

Lewis:1963:TSL

[Lew63] J. W. Lewis. Time sharing on Leo III. The Computer Journal, 6 (1):24–28, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/24.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060024.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/26.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/27.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/28.
tif.

Lewis:1965:IAC

[Lew65a] C. D. Lewis. Iterative analogue computation applied to inventory policy simulation. The Computer Journal, 8(2):130–135, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/2/130.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/080130.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/131.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/132.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/133.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/134.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/135.tif.

Lewis:1965:MLC

[Lew65b] J. W. Lewis. The management of a large commercial computer bureau. The Computer Journal, 7(4):255–261, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/255.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070255.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
255.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/256.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/257.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/258.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/259.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/260.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/261.tif.

Lindley:1968:PA

[Lin68] D. V. Lindley. Prediction analysis. The Computer Journal, 11(2): 133, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/133.full.pdf+html.

Linz:1969:NMV

[Lin69] P. Linz. Numerical methods for Volterra integral equations of the first kind. The Computer Journal, 12(4):393–397, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/393.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/120393.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/tiff/393.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/394.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/395.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

12/Issue_04/tiff/396.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/397.tif.

Lions:1967:OSS

[Lio67] J. Lions. The Ontario school scheduling program. The Computer Journal, 10(1):14–21, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/14.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100014.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
14.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/15.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/16.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
17.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/18.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/19.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
20.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/21.tif.

Livesley:1960:ALS

[Liv60] R. K. Livesley. The analysis of large structural systems. The Computer Journal, 3(1):34–39, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/1/34.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/030034.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/34.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/35.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/36.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/37.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/38.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_01/tiff/39.tif.

Lloyd:1964:DR

[Llo64] D. B. Lloyd. Data retrieval. The Computer Journal, 7(2):110–113, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/2/110.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_07/Issue_02/070110.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/110.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/111.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/
tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/113.tif.

Lloyd:1968:BR

[Llo68] E. H. Lloyd. Book review. The Computer Journal, 10(4):364, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/364.full.pdf+html.

Lloyd:1969:BR

[Llo69] E. Lloyd. Book review. The Computer Journal, 12(2):153, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/153.full.pdf+html.

Lambert:1963:UHD

[LM63a] J. D. Lambert and A. R. Mitchell. The use of higher derivatives in quadrature formulae. The Computer Journal, 5(4):322–327, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/4/322.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/050322.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/322.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/323.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/324.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/325.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/326.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/327.tif.

Lyness:1963:IMH

[LM63b] J. N. Lyness and B. J. J. McHugh. Integration over multidimensional hypercubes I. A progressive procedure. The Computer Journal, 6(3):264–270, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/264.full.

pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060264.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
264.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/265.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/266.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/268.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/269.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/270.tif.

Lloyd:1968:BOR

[LM68] T. Lloyd and H. McCallion. Bounds for the optimum overrelaxation factor for the S.O.R. solution of Laplace type equations over irregular regions. The Computer Journal, 11(3):329– 331, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/329.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/110329.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/329.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/330.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/331.tif.

Lunnon:1965:C

[LO65] W. F. Lunnon and C. F. J. Outred. Correspondence. The Computer Journal, 8(2):175, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/2/175.2.full.pdf+html.

Loizou:1968:EER

[Loi68] G. Loizou. An empirical estimate of the relative error of the computed solution x of Ax = b. The Computer Journal, 11(1):91–94, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/91.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110091.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/92.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
93.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/94.tif.

Long:1961:DTS

[Lon61] P. A. Long. A data transmission survey. The Computer Journal, 4 (2):95–102, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/2/95.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/040095.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/95.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
96.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/97.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/98.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
99.tif.

Lavington:1967:SFS

[LR67] S. H. Lavington and L. E. Rosenthal. Some facilities for speech processing by computer. The Computer Journal, 9(4):330–339, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/330.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090330.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/330.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/331.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/332.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/333.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/334.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/336.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/338.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/339.
tif.

Long:1961:DTP

[LT61] P. A. Long and E. H. Truslove. Data transmission — problems and prospects. The Computer Journal, 4(1):34–37, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/34.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/040034.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/34.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/36.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/37.tif.

Lukaszewicz:1967:ESM

[Luk67] L. Lukaszewicz. EOL — a symbol manipulation language. The Computer Journal, 10(1):53–59, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/53.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100053.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
53.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/55.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
56.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/57.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/58.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
59.tif.

Lumb:1964:EPD

[Lum64] O. S. Lumb. Experience of program development with FACT. The Computer Journal, 7(2):104–107, July 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/104.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070104.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
104.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_07/Issue_02/tiff/105.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/106.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/tiff/107.tif.

Lunnon:1969:PSP

[Lun69] W. F. Lunnon. A postage stamp problem. The Computer Journal, 12(4):377–380, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/377.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120377.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
377.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/378.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/379.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/

Issue_04/tiff/380.tif. See correspondence [Sel72].

Lance:1962:BR

[LW62] G. N. Lance and J. C. West. Book reviews. The Computer Journal, 5(3):245–246, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/5/3/245.full.pdf+html.

Lance:1965:CPM

[LW65] G. N. Lance and W. T. Williams. Computer programs for monothetic classification (‘association analysis’). The Computer Journal, 8(3):246–249, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/246.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080246.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
246.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/247.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/248.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/249.tif.

Lance:1966:CPH

[LW66a] G. N. Lance and W. T. Williams. Computer programs for hierarchical polythetic classification (‘similarity analyses’). The Computer Journal, 9(1):60–64, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/60.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090060.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/60.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/61.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/62.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/63.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/64.tif.

Lindley:1966:NSS

[LW66b] J. Lindley and A. Wragg. The numerical solution of sequential decision problems involving parabolic equations with moving boundaries. The Computer Journal, 9(3):294–297, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/294.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090294.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/294.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/295.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/296.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/297.tif.

Lance:1967:GTCa

[LW67a] G. N. Lance and W. T. Williams. A general theory of classificatory sorting strategies 1. Hierarchical systems. The Computer Journal, 9(4):373–380, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/373.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090373.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
373.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/374.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/375.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/376.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/377.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/

tiff/378.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/379.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/380.
tif.

Lance:1967:GTCb

[LW67b] G. N. Lance and W. T. Williams. A general theory of classificatory sorting strategies II. Clustering systems. The Computer Journal, 10(3):271–277, November 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/3/271.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/100271.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
271.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/272.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/273.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_03/tiff/274.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/275.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/276.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/277.tif.

Lance:1967:NCM

[LW67c] G. N. Lance and W. T. Williams. Note on the classification of multi-level data. The Computer Journal, 9(4):381–382, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/4/381.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_04/090381.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_04/tiff/381.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/382.tif.

Lance:1968:NNI

[LW68] G. N. Lance and W. T. Williams. Note on a new informationstatistic classificatory program. The Computer Journal, 11(2): 195, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/195.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/110195.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/195.tif.

Larmouth:1968:MTM

[LWS68] J. Larmouth and C. Whitby-Strevens. MLS — the Titan mixed language system. The Computer Journal, 11(3):256–259, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/256.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/110256.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/tiff/256.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/257.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/258.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

11/Issue_03/tiff/259.tif. See correspondence [PB69, LWS69c, GB69].

Landy:1969:TTS

[LWS69a] B. Landy and C. Whitby-Strevens. TSAS — the time-shared supervisor assembly system. The Computer Journal, 12(2):128–131, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/128.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/120128.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
02/tiff/128.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/129.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/130.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_02/tiff/131.tif.

Larmouth:1969:C

[LWS69b] J. Larmouth and C. Whitby-Strevens. Correspondence. The Computer Journal, 12(2):200, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/2/200.2.full.pdf+html.

Larmouth:1969:CP

[LWS69c] J. Larmouth and C. Whitby-Strevens. Correspondence: Paging. The Computer Journal, 12(2):200, May 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_02/tiff/200.tif. See [LWS68, PB69, GB69].

Lynn:1964:ESS

[Lyn64] M. S. Lynn. On the equivalence of SOR, SSOR and US-SOR as applied to σ_{1}-ordered systems of linear equations. The Computer Journal, 7(1):72–75, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/1/72.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070072.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/72.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/73.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/74.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/75.
tif.

Lynch:1965:MCP

[Lyn65] William C. Lynch. More combinatorial properties of certain trees. The Computer Journal, 7(4):299–302, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/299.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070299.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
299.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/300.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/301.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/302.tif.

Lyness:1969:EIC

[Lyn69] J. N. Lyness. The effect of inadequate convergence criteria in automatic routines. The Computer Journal, 12(3):279–281, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
3/279.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/120279.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/279.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/280.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/281.

tif. See correspondence [GL70].

Maguire:1964:BRP

[Mag64] T. A. Maguire. Book review: Parity-check codes. The Computer Journal, 7(2):113, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/113.full.pdf+html.

Maguire:1965:BR

[Mag65] T. A. Maguire. Book review. The Computer Journal, 8(2):135, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
2/135.full.pdf+html.

Makinson:1968:AEC

[Mak68a] G. J. Makinson. Algorithm 36: Evaluation of the complex arcsin. The Computer Journal, 11(3):348–350, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/348.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/349.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/350.tif.

Makinson:1968:SHO

[Mak68b] G. J. Makinson. Stable high order implicit methods for the numerical solution of systems of differential equations. The Computer Journal, 11(3):305–310, November 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/305.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110305.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
305.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/306.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/307.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/308.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/309.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/310.tif.

Maley:1961:EPR

[Mal61] C. E. Maley. The effect of parameters on the roots of an equation system. The Computer Journal, 4(1):62–63, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/62.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/040062.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/62.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/63.tif.

Martin:1962:BR

[Mar62] D. W. Martin. Book reviews. The Computer Journal, 5(3):227, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/3/227.1.full.pdf+html.

Marin:1969:STN

[Mar69] Miguel A. Marin. Synthesis of TANT networks using a Boolean analyser. The Computer Journal, 12(3):259–267, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
3/259.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/120259.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/259.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/260.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/261.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/262.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/263.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/264.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/265.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/266.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/267.tif.

Massey:1963:CNS

[Mas63] R. G. Massey. Computers in a new steelworks. The Computer Journal, 5(4):271–275, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/271.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050271.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/

271.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/272.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/273.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/274.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/275.tif.

Matthewman:1963:NSS

[Mat63] J. H. Matthewman. Note on the selective summation of Fourier series. The Computer Journal, 6(3):248–249, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/3/248.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_03/060248.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_03/tiff/248.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/249.tif.

Matthewman:1967:CF

[Mat67] J. C. Matthewman. The crystallographers’ friend. The Computer Journal, 10(1):38–45, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/38.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100038.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
38.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/40.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
41.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/43.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
44.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/45.tif.

Mayers:1964:DAL

[May64] D. F. Mayers. The deferred approach to the limit in ordinary differential equations. The Computer Journal, 7(1):54– 57, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/1/54.full.pdf+html;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_07/Issue_01/070054.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/56.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/57.tif.

Mayne:1966:ACP

[May66] David Mayne. An algorithm for the calculation of the pseudoinverse of a singular matrix. The Computer Journal, 9(3):312–317, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/312.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090312.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/312.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/313.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/314.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/315.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/316.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/317.tif.

Mayne:1969:BRB

[May69] D. Q. Mayne. Book review: Conditional Markov Processes and their application to the Theory of Optimal Control. The Computer Journal, 12(1):5, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/1/5.full.pdf+html.

Miller:1966:AER

[MB66] J. C. P. Miller and D. J. Spencer Brown. An algorithm for evaluation of remote terms in a linear recurrence sequence. The Computer Journal, 9(2):188–190, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/188.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090188.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
188.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/189.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/190.tif.

McEwan:1967:AAA

[McE67] A. T. McEwan. An Atlas Autocode to ALGOL 60 translator. The Computer Journal, 9(4):353–359, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/353.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090353.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
353.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/354.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/355.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/356.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/357.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/358.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/359.tif.

McManus:1968:BR

[McM68] Peter McManus. Book review. The Computer Journal, 10(4): 401, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/401.full.pdf+html.

McOustra:1966:LPC

[McO66] C. J. C. McOustra. Legal protection for computer programs. The Computer Journal, 8(4):289–296, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/289.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080289.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
289.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/290.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/291.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_04/tiff/292.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/293.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/294.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/295.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/296.
tif.

Meggitt:1962:BR

[Meg62] J. E. Meggitt. Book review. The Computer Journal, 5(1):6, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
1/6.full.pdf+html.

Meister:1966:FCF

[Mei66] Bernd Meister. On a family of cubature formulae. The Computer Journal, 8(4):368–371, January 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/368.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080368.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
368.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/369.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/370.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_04/tiff/371.tif.

Meltzer:1966:TPC

[Mel66] B. Meltzer. Theorem-proving for computers: some results on resolution and renaming. The Computer Journal, 8(4):341–343, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/341.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080341.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/343.tif.

Merton:1961:C

[Mer61] C. R. Merton. Correspondence. The Computer Journal, 4(2): 120, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/2/120.full.pdf+html.

Maloney:1961:C

[MF61] C. J. Maloney and R. A. Fairthorne. Correspondence. The Computer Journal, 4(1):85, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/4/1/85.full.pdf+html.

Milburn:1964:CI

[MFM64] J. F. Milburn, D. J. Fearnley, and C. G. Myers. A computer in insurance. The Computer Journal, 7(3):173–176, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/173.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070173.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
173.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/174.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/175.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/176.tif.

Michie:1963:EMG

[Mic63] Donald Michie. Experiments on the mechanization of game-learning Part I. Characterization of the model and its parameters. The Computer Journal, 6(3):232–236, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/232.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060232.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
232.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/233.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/234.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/236.tif.

Michaelson:1964:PAS

[Mic64]  R. L. Michaelson.  Presidential address. Science and the

non-scientist. The Computer Journal, 6(4):299–303, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/4/299.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_04/060299.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/299.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/300.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/
tiff/301.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/302.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/303.
tif.

Miller:1960:NNE

[Mil60] J. C. P. Miller. Note on the numerical evaluation of a first derivative from a table of a function satisfying a second order differential equation. The Computer Journal, 3(2):112–113, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/2/112.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/030112.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/
112.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/113.tif.

Miller:1962:BR

[Mil62] J. C. P. Miller. Book review. The Computer Journal, 5(3):199, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/3/199.2.full.pdf+html.

Mills:1963:OET

[Mil63] M. R. Mills. Operational experience of time sharing and parallel processing. The Computer Journal, 6(1):28–36, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/1/28.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/060028.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/28.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/29.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
30.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/31.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/32.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
33.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/34.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/35.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
36.tif.

Miller:1964:BRM

[Mil64] J. C. P. Miller. Book reviews: Mathematical tables. The Computer Journal, 7(1):53, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/1/53.full.pdf+html.

Miller:1968:CMT

[Mil68a] H. D. Miller. Combinatorial methods in the theory of stochastic processes. The Computer Journal, 11(2):208, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/208.full.pdf+html.

Milner:1968:SHA

[Mil68b] R. Milner. String handling in Algol. The Computer Journal, 10(4):321–324, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/4/321.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/100321.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/321.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/322.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/323.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/324.
tif.

Miller:1969:CA

[Mil69] J. C. P. Miller. Computer approximations. The Computer Journal, 12(3):272, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/3/272.full.pdf+html.

Mitchell:1969:BR

[Mit69] A. J. Mitchell. Book review. The Computer Journal, 12(4):326, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/326.full.pdf+html.

Murray:1965:COD

[ML65] W. A. Murray and M. S. Lynn. A computer-oriented description of the Peaceman–Rachford ADI method. The Computer Journal, 8(2):

166–175, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/2/166.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/080166.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/166.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/tiff/167.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/169.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/170.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/171.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/172.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/173.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/174.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/175.tif.

Mustard:1963:NQD

[MLB63] D. Mustard, J. N. Lyness, and J. M. Blatt. Numerical quadrature in n dimensions. The Computer Journal, 6(1):75– 87, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/75.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060075.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/75.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/77.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/79.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
82.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/84.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
85.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

06/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/87.tif.

Milledge:1960:FER

[MM60] D. Milledge and M. J. Mills. Forecasting election results. The Computer Journal, 2(4):195–198, January 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/4/195.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/020195.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/
195.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/tiff/196.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_04/tiff/197.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/
Issue_04/tiff/198.tif.

Marliss:1969:WGB

[MM69] G. S. Marliss and W. A. Murray. William G. Bickley — an appreciation. The Computer Journal, 12(4):301–302, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/301.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/120301.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/tiff/301.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/302.tif.

Merry:1961:CQS

[MN61] I. W. Merry and G. O. Norrie. Character quality and scanner organization. The Computer Journal, 4(2):137–144, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/137.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040137.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
137.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/138.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/139.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/140.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/141.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/

tiff/142.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/143.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/144.
tif.

Moore:1963:ASP

[Moo63] D. W. Moore. Accounting for the soldier’s pay. The Computer Journal, 5(4):249–257, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/249.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050249.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
249.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/250.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/251.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/252.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/253.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/254.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/255.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/256.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/257.tif.

Moore:1965:C

[Moo65] R. H. Moore. Correspondence. The Computer Journal, 8(3): 224, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/3/224.full.pdf+html.

Morris:1969:ALP

[Mor69] G. P. D. Morris. Advanced linear-programming computing techniques. The Computer Journal, 12(3):232, August 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/12/3/232.
full.pdf+html.

Mitchell:1962:HAD

[MP62] A. R. Mitchell and R. P. Pearce. High accuracy difference formulae for the numerical solution of the heat conduction equation. The Computer Journal, 5(2):142–146, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/2/142.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/050142.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
142.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/143.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/tiff/144.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_02/tiff/145.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/146.tif.

Martin:1965:ECD

[MP65a] E. J. Martin, Jr. and P. C. Patton. Evaluation of certain definite integrals frequently encountered in radiational and diffraction problems involving circular geometry. The Computer Journal, 8(3):256–263, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/256.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080256.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
256.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/257.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/258.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/259.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/260.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/261.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/262.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/263.

tif. See erratum [MP66].

Melbourne:1965:SCD

[MP65b] Alan J. Melbourne and John M. Pugmire. A small computer for the direct processing of FORTRAN statements. The Computer Journal, 8(1):24–27, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/8/1/24.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_01/080024.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/24.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/25.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/27.
tif.

Martin:1966:EEC

[MP66] E. J. Martin, Jr. and P. C. Patton. Erratum: “Evaluation of certain definite integrals frequently encountered in radiational and diffraction problems involving circular geometry”. The Computer Journal, 9(1):116, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/116.tif. See [MP65a].

Morgenstern:1967:NMS

[MP67] N. R. Morgenstern and V. E. Price. A numerical method for solving the equations of stability of general slip surfaces. The Computer Journal, 9(4):388–393, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/388.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090388.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
388.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/389.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/390.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/391.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/392.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/393.tif.

Miller:1968:ACP

[MP68] J. C. P. Miller and M. J. Prentice. Additive congruential pseudorandom number generators. The Computer Journal, 11(3):341– 346, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/341.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/110341.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/341.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/342.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/343.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_11/Issue_03/tiff/344.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/
tiff/345.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/346.tif.

Moore:1965:CRF

[MR65] Roger H. Moore and R. H. Rosenbrock. Correspondence: On the Rosenbrock function. The Computer Journal, 8(3):224, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/224.tif.

Morris:1967:ACS

[MR67] D. Morris and J. S. Rohl. The Atlas compiler system. The Computer Journal, 10(3):227–230, November 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/3/227.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/100227.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
227.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/228.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/229.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_03/tiff/230.tif.

Miles:1964:EDL

[MST64] G. A. Miles, K. L. Stewart, and G. J. Tee. Elementary divisors of the Liebmann process. The Computer Journal, 6(4):352–355, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/4/352.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_04/060352.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/352.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/353.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/
tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/355.tif.

Martin:1961:IML

[MT61] D. W. Martin and G. J. Tee. Iterative methods for linear equations with symmetric positive definite matrix. The Computer Journal, 4(3):242–254, October 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/3/242.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/040242.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/
242.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/243.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_03/tiff/244.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_03/tiff/245.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/246.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/247.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/248.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/249.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/251.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/253.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/254.
tif.

Mulholland:1969:STT

[Mul69] K. A. Mulholland. Software to translate TELCOMP programs into KDF9ALGOL. The Computer Journal, 12(3):221–224, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
3/221.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/120221.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/221.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/222.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/223.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/224.tif.

Murphy:1968:BR

[Mur68] J. W. Murphy. Book review. The Computer Journal, 11(3):246, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/3/246.full.pdf+html.

Musk:1966:OMMb

[Mus66a] F. I. Musk. One man’s meat: part 2 — Now let’s pretend. The Computer Journal, 9(3):213–214, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/213.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090213.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
213.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/214.tif.

Musk:1966:OMMa

[Mus66b] F. I. Musk. One man’s meat: part I — The uses of adversity. The Computer Journal, 9(1):1–2, May 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/9/1/1.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090001.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/1.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/2.tif.

Musk:1967:OMM

[Mus67] F. I. Musk. One man’s meat: part 3 — It’s all good grist that comes to our mill. The Computer Journal, 10(2):126–127, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
2/126.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/100126.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
02/tiff/126.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/127.tif.

Mutch:1967:EBC

[Mut67] E. N. Mutch. Editorship of The Computer Journal. The Computer Journal, 9(4):420, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/4/420.full.pdf+html.

Mitra:1968:CTO

[MW68] G. Mitra and K. Wolfenden. A computer technique for optimizing the sites and heights of transmission line towers — a dynamic programming approach. The Computer Journal, 10(4):347– 351, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/347.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/100347.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/347.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/348.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/350.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/351.tif.

Mylander:1966:CNP

[Myl66] W. Charles Mylander. Correspondence: Nonlinear programming test problems. The Computer Journal, 8(4):391, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/391.tif. See comments [Box66b].

Napper:1967:SPS

[Nap67] R. B. E. Napper. Some proposals for SNAP, a language with formal macro facilities. The Computer Journal, 10(3):231–243, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/231.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/100231.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
03/tiff/231.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/232.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/233.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/234.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/235.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/236.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/237.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/238.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/239.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/240.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/241.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/242.tif;   http://www3.oup.

co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/243.
tif.

Nathan:1967:ISM

[NE67] A. Nathan and R. K. Even. The inversion of sparse matrices by a strategy derived from their graphs. The Computer Journal, 10(2):190–194, August 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/190.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/100190.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
190.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/191.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/192.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/193.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/194.tif.

Needham:1961:BRa

[Nee61a] R. M. Needham. Book review. The Computer Journal, 3(4): 265, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/4/265.1.full.pdf+html.

Needham:1961:BRb

[Nee61b] R. M. Needham. Book review. The Computer Journal, 3(4): 265, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/4/265.2.full.pdf+html.

Needham:1963:BR

[Nee63a] R. M. Needham. Book review. The Computer Journal, 6(3): 249, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/249.full.pdf+html.

Needham:1963:BRB

[Nee63b] R. M. Needham. Book review: Textbook on Mechanized Information Retrieval, by Allen Kent, 1962. The Computer Journal, 6(3): 249, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/249.tif.

Neff:1965:DIY

[Nef65] B. L. Neff. Do-it-yourself software experience. The Computer Journal, 8(3):201–207, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/201.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080201.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
201.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/202.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/203.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/204.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/205.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/206.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/207.tif.

Nelder:1968:BR

[Nel68] J. A. Nelder. Book review. The Computer Journal, 10(4):367, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/367.full.pdf+html.

Newman:1961:SCC

[New61] E. A. Newman. Some comments on character recognition. The Computer Journal, 4(2):114–120, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/114.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040114.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
114.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/115.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/116.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/117.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/118.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/120.tif.

Newman:1965:BRB

[New65a] E. A. Newman. Book review: An Introduction to Cybernetics. The Computer Journal, 7(4):277a–277, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
4/277-a.full.pdf+html.

Newman:1965:IC

[New65b] E. A. Newman. An introduction to cybernetics. The Computer Journal, 7(4):277a–277, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/4/277-a.full.pdf+html.

Newman:1966:EPA

[New66] W. M. Newman. An experimental program for architectural design. The Computer Journal, 9(1):21–26, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/21.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090021.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/21.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/22.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/24.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/25.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/26.tif.

Newman:1968:GTN

[New68] W. M. Newman. A graphical technique for numerical input. The Computer Journal, 11(1):63–64, May 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/63.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110063.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
63.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/64.tif.

Nicholson:1966:FSR

[Nic66] T. A. J. Nicholson. Finding the shortest route between two points in a network. The Computer Journal, 9(3):275–280, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/275.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090275.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/275.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/276.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/277.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/278.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/279.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/280.tif.

Nievergelt:1965:FVS

[Nie65] J. Nievergelt. Fixed versus selfmodifying programs. The Computer Journal, 8(3):244–245, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/244.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080244.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
244.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/245.tif.

Nightingale:1968:PCS

[Nig68] J. M. Nightingale. Process control systems. The Computer Journal, 11(2):184, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/2/184.full.pdf+html.

Nixon:1964:BRC

[Nix64] W. L. B. Nixon. Book review: Computers in research. The Computer Journal, 7(1):27, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/1/27.full.pdf+html.

Nelder:1965:SMF

[NM65] J. A. Nelder and R. Mead. A simplex method for function minimization. The Computer Journal, 7(4):308–313, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
4/308.full.pdf+html.

Norton:1964:ISN

[Nor64] H. J. Norton. The iterative solution of non-linear ordinary differential equations in Chebyshev series. The Computer Journal, 7(1):76–85, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/1/76.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070076.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/77.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/79.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/
82.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/84.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/
85.tif.

Nicholson:1968:PPJ

[NP68] T. A. J. Nicholson and R. D. Pullen. A permutation procedure for job-shop scheduling. The Computer Journal, 11(1):48–56, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/48.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110048.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/48.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/49.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
50.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/51.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/52.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
53.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/54.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/55.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
56.tif.

Noronha:1968:HCD

[NPW68] L. G. Noronha, C. Y. Po, and J. W. Womack. Hybrid computation of the dynamics of a distributed system. The Computer Journal, 11(2):196–205, August 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/2/196.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/110196.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/
196.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/197.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/198.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/199.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/200.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/201.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/202.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/203.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/204.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/205.tif.

Nudds:1969:MCE

[Nud69] D. Nudds. Methods of computing event times in project networks. The Computer Journal, 12(1):38–40, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/38.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120038.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
38.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/40.tif.

Oatley:1963:BR

[Oat63] C. W. Oatley. Book review. The Computer Journal, 6(2):117, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/2/117.full.pdf+html.

OBrien:1967:CCO

[O’B67] W. O’Brien. COSMOS — Courtauld’s own system for matrix operations and statistics. The Computer Journal, 10(2):128– 134, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/2/128.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/100128.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/128.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/129.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/131.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/132.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/133.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/134.
tif.

OBrien:1968:TCF

[OB68a] F. O’Brien and R. C. Beckwith. A technique for computer flow chart generation. The Computer Journal, 11(2):138–140, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/138.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110138.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/139.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/140.
tif.

OBrien:1968:CNM

[O’B68b] J. P. O’Brien. Correspondence: On “A new method for solving polynomial equations”. The Computer Journal, 11(2):194, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/194.tif. See [GJM68].

OBrien:1968:LEN

[O’B68c] J. P. O’Brien. Letter to the Editor: a new method for solving polynomial equations. The Computer Journal, 11(2):194, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/194.full.pdf+html.

Obruca:1967:CBT

[Obr67] A. K. Obruca. Correspondence between two types of trees. The Computer Journal, 10(3):290–292, November 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/3/290.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/100290.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
290.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/291.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/292.tif.

Obruca:1968:STM

[Obr68] Alex K. Obruca. Spanning tree manipulation and the travelling salesman problem. The Computer Journal, 10(4):374–377, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/374.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/100374.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/374.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/375.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/376.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/377.tif.

OCarroll:1969:DCR

[O’C69] M. J. O’Carroll. Discussion and correspondence: S.O.R. and membranes. The Computer Journal, 12(3):296–297, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
3/296.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/120296.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/296.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/297.tif.

Ortega:1963:MST

[OK63] J. M. Ortega and H. F. Kaiser. The LL^{T} and QR methods for symmetric tridiagonal matrices. The Computer Journal, 6(1):99– 101, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/99.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060099.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/tiff/101.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/tiff/99.tif.

Owen:1968:CAA

[OL68] David P. Owen and John Laski. Correspondence: An argument against paging hardware. The Computer Journal, 11(3): 351, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/tiff/351.tif.

Oliver:1969:EET

[Oli69] J. Oliver. An error estimation technique for the solution of ordinary differential equations in Chebyshev series. The Computer Journal, 12(1):57–62, February 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/57.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120057.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
57.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/59.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
60.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/62.tif.

Osborne:1964:NSE

[OM64] M. R. Osborne and S. Michaelson. The numerical solution of eigenvalue problems in which the eigenvalue problems in which the eigenvalue parameter appears nonlinearly, with an application to differential equations. The Computer Journal, 7(1):66–71, April 1964. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/1/66.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070066.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/67.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/68.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/69.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/70.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/71.tif.

Ongley:1969:EC

[Ong69] P. A. Ongley. Examinations by computer. The Computer Journal, 12(3):297–298, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/12/3/297.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/120297.sgm.abs.html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/297.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/298.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/299.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/300.tif.

Orden:1964:SPQ

[Ord64] A. Orden. Stationary points of quadratic functions under linear constraints. The Computer Journal, 7(3):238–241, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/3/238.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_03/070238.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/239.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_03/tiff/241.tif.

Ord-Smith:1965:BRB

[OS65a] R. J. Ord-Smith. Book review: Irascible Genius (Charles Babbage). The Computer Journal, 7(4):277, January 1965.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/

4/277.full.pdf+html. See comments and corrections [Phi65].

Ord-Smith:1965:EBD

[OS65b] R. J. Ord-Smith. An extension of block design methods and an application in the construction of redundant fault reducing circuits for computers. The Computer Journal, 8(1):28– 32, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/28.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/tiff/28.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_01/tiff/29.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/30.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
31.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/32.tif.

Ord-Smith:1967:BRa

[OS67a] R. J. Ord-Smith. Book review. The Computer Journal, 10(1): 45, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/1/45.1.full.pdf+html.

Ord-Smith:1967:BRb

[OS67b] R. J. Ord-Smith. Book review. The Computer Journal, 10(1): 45, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/1/45.2.full.pdf+html.

OHara:1968:EEC

[OS68a] H. O’Hara and Francis J. Smith. Error estimation in the Clenshaw– Curtis quadrature formula. The Computer Journal, 11(2):213– 219, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/213.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/110213.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/213.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/214.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/215.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_11/Issue_02/tiff/216.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/218.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/219.
tif.

Ord-Smith:1968:BR

[OS68b] R. J. Ord-Smith. Book review. The Computer Journal, 10(4): 377, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/377.full.pdf+html.

OHara:1969:EDI

[OS69] H. O’Hara and Francis J. Smith. The evaluation of definite integrals by interval subdivision. The Computer Journal, 12(2):179–182, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
2/179.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/120179.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
02/tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/180.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/181.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_02/tiff/182.tif.

Osborne:1963:IPS

[Osb63] M. R. Osborne. Iterative procedures for solving finite-difference approximations to separable partial differential equations. The Computer Journal, 6(1):93–99, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/6/1/93.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060093.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/93.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/94.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/95.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/96.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/97.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/98.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
99.tif.

Osborne:1964:EAF

[Osb64a] M. R. Osborne. An error analysis of finite-difference methods for the numerical solution of ordinary differential equations. The Computer Journal, 7(3):232–237, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/232.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070232.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
232.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/233.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/234.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/236.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/237.tif.

Osborne:1964:MFD

[Osb64b] M. R. Osborne. A method for finite-difference approximation to ordinary differential equations. The Computer Journal, 7(1): 58–65, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/1/58.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_01/070058.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/59.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/60.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/61.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/62.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/63.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/64.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/
65.tif.

Osborne:1964:NMS

[Osb64c] M. R. Osborne. A new method for the solution of eigenvalue problems. The Computer Journal, 7(3):228–232, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/228.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070228.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
228.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/229.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/230.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/231.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/232.tif.

Osborne:1965:DMS

[Osb65] M. R. Osborne. Direct methods for the solution of finitedifference approximations to separable partial differential equations. The Computer Journal, 8(2):150–156, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/2/150.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/080150.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/
150.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/151.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/tiff/152.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_02/tiff/153.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/154.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/
tiff/155.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/156.tif.

Osborne:1966:BRB

[Osb66] M. R. Osborne. Book review: Introductory Numerical Analysis of Elliptic Boundary Value Problems. The Computer Journal, 9 (1):44, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/1/44.full.pdf+html.

Osborne:1969:NSH

[Osb69] M. R. Osborne. The numerical solution of the heat conduction equation subject to separated boundary conditions. The Computer Journal, 12(1):82–87, February 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/82.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120082.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
82.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/84.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
85.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/87.tif.

ONeill:1969:PCF

[OSS69] M. O’Neill, I. G. Sinclair, and Francis J. Smith. Polynomial curve fitting when abscissas and ordinates are both subject to error. The Computer Journal, 12(1):52–56, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/52.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120052.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
52.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/54.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
55.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/56.tif.

Overholt:1967:AGa

[Ove67a] K. J. Overholt. Algorithm 16: GOLD. The Computer Journal, 9(4): 415, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/415.tif.

Overholt:1967:AGb

[Ove67b] K. J. Overholt. Algorithm 17: GOLDSEC. The Computer Journal, 9(4):415, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_09/Issue_04/tiff/415.tif.

Overholt:1967:NAF

[Ove67c] K. J. Overholt. Note on Algorithm 2: Fibonacci search. The Computer Journal, 9(4):414, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/414.

tif. See [PP79].

OHalloran:1966:NRI

[OW66] M. O’Halloran and W. M. Waite. Note on rapid instruction analysis by table lookup. The Computer Journal, 9(3):248, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/248.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090248.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/248.tif.

Osborne:1967:BLC

[OW67] M. R. Osborne and G. A. Watson. On the best linear Chebyshev approximation. The Computer Journal, 10(2):172–177, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
2/172.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/100172.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
02/tiff/172.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/173.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/174.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/176.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/177.tif.

Osborne:1968:NTM

[OW68] M. R. Osborne and G. A. Watson. Note on two methods of solving ordinary linear differential equations. The Computer Journal, 10(4):383–384, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/383.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100383.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
383.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/384.tif.

Osborne:1969:AMA

[OW69] M. R. Osborne and G. A. Watson. An algorithm for minimax approximation in the nonlinear case. The Computer Journal, 12(1): 63–68, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/1/63.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_01/120063.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/tiff/63.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
01/tiff/65.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_01/tiff/66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/67.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

12/Issue_01/tiff/68.tif. See correspondence [Tem69b].

Owen:1965:POR

[Owe65] D. G. Owen. Progress in operations research. The Computer Journal, 8(1):12, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/12.full.pdf+html.

Owen:1968:C

[Owe68] D. P. Owen. Correspondence. The Computer Journal, 11(3): 351, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/351.full.pdf+html.

Page:1960:TCR

[Pag60] E. S. Page. Theoretical considerations of routine maintenance. The Computer Journal, 2(4):199–204, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/2/4/199.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/020199.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/
199.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_02/Issue_04/tiff/200.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_02/Issue_04/tiff/201.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_02/
Issue_04/tiff/202.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_02/Issue_04/tiff/203.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/
tiff/204.tif.

Page:1962:SJC

[Pag62] E. S. Page. On the scheduling of jobs by computer. The Computer Journal, 5(3):214–220, November 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/214.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/050214.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/
214.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/215.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_03/tiff/216.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_03/tiff/217.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/218.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/219.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/220.tif.

Page:1963:BR

[Pag63a] E. S. Page. Book reviews. The Computer Journal, 6(3):270, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/270.full.pdf+html.

Page:1963:NAP

[Pag63b] E. S. Page. A note on assignment problems. The Computer Journal, 6(3):241–243, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/241.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060241.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
241.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/242.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/243.tif.

Paine:1960:ACB

[Pai60] R. M. Paine. Automatic coding for business applications. The Computer Journal, 3(3):144–149, October 1960. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/3/144.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/030144.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/
144.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_03/tiff/145.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_03/tiff/146.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_03/tiff/147.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/148.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/
tiff/149.tif.

Paine:1965:GAV

[Pai65] R. M. Paine. The gradual acceptance of a variety of commercial English languages. The Computer Journal, 8(3):208–215, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/3/208.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_03/080208.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/208.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/209.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/210.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/211.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/212.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/213.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/214.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/215.tif.

Paine:1966:POC

[Pai66] R. M. Paine. Preparation for optical character recognition. The Computer Journal, 9(3):221–229, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/221.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090221.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
221.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/222.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/223.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/224.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/225.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/226.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/227.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/228.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/229.tif.

Paine:1969:CRB

[Pai69] R. M. Paine. O.C.R. — benefits and pitfalls. The Computer Journal, 12(2):108–113, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/12/2/108.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_02/120108.sgm.abs.html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/108.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/109.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/110.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/111.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/112.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/113.

tif. See correspondence [WP69].

Pakes:1967:C

[Pak67a] H. W. Pakes. Correspondence. The Computer Journal, 9(4): 403, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/4/403.full.pdf+html.

Pakes:1967:CID

[Pak67b] H. W. Pakes. Correspondence: On “An investigation into direct numerical methods for solving some calculus of variations problems”. The Computer Journal, 9(4):403, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090403.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/

403.tif. See [All66a].

Palmer:1962:SCI

[Pal62] D. R. Palmer. A stock-control and invoicing system using a Gamma 3 computer. The Computer Journal, 5(1):7– 9, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/1/7.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_01/050007.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/7.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/8.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
01/tiff/9.tif.

Palmer:1963:CCI

[Pal63] R. Palmer. Computer calculations on the initiation of highexplosive detonation. The Computer Journal, 6(1):39–43, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/1/39.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_01/060039.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/40.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/41.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/42.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/43.
tif.

Palmer:1969:IPO

[Pal69] J. R. Palmer. An improved procedure for orthogonalising the search vectors in Rosenbrock’s and Swann’s direct search optimisation methods. The Computer Journal, 12(1):69–71, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/69.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120069.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/70.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
71.tif.

Papworth:1960:CCR

[Pap60] D. G. Papworth. Computers and change-ringing. The Computer Journal, 3(1):47–50, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/1/47.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/030047.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/48.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/50.
tif.

Pape:1968:CSC

[Pap68a] Uwe Pape. Correspondence: On “Some computational notes on the shortest route problems”. The Computer Journal, 11(2):240, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/240.tif. See [Per65].

Pape:1968:LES

[Pap68b] Uwe Pape. Letter to the Editor: Some computational notes on the shortest route problem. The Computer Journal, 11(2):240, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/240.full.pdf+html.

Parlett:1967:LEF

[Par67] Beresford Parlett. Letter to the Editor: On finding the eigenvalues of real symmetric tridiagonal matrices. The Computer Journal, 9(4):344–345, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/4/344.full.pdf.

Payne:1963:C

[Pay63] J. C. F. Payne. Correspondence. The Computer Journal, 5(4): 312, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/4/312.full.pdf+html.

Payne:1966:BRB

[Pay66] L. C. Payne. Book review: Mathematics and Computer Science in Biology and Medicine. The Computer Journal, 9(1):10, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
1/10.full.pdf+html.

Pinkerton:1969:CP

[PB69] Tad Pinkerton and D. W. Barron. Correspondence: Paging. The Computer Journal, 12(1):105, February 1969. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/105.tif. See [LWS68, LWS69c, GB69].

Parker:1964:PDE

[PC64] I. B. Parker and J. Crank. Persistent discretization errors in partial differential equations of parabolic type. The Computer Journal, 7(2):163–167, July 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/163.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070163.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
163.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/164.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/165.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/tiff/166.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/167.tif. See comment

[KM66] and correspondence [Ros66a].

Perry:1963:BRB

[PCSG63] G. H. Perry, J. C. Cluley, K. L. Smith, and R. L. Goodstein. Book reviews: Large-Capacity Memory Techniques for Computing Systems, edited by Marshall C. Yovits, 1962; Digital Techniques, by D. W. Davies, 1963; Theory and Design of Digital Machines, by Thomas C. Bartee, Irwin Lebow, [and] Irving S. Reed, 1962; A Survey of Mathematical Logic, by Hao Wang, 1963. The Computer Journal, 6(3):295–298, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/295.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/296.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/297.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/298.
tif.

Pearson:1963:SC

[Pea63] K. W. Pearson. Satellite communications. The Computer Journal, 5(4):308–312, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/5/4/308.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
04/050308.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/308.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/309.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/310.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/311.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/312.tif.

Pearson:1969:VMM

[Pea69] J. D. Pearson. Variable metric methods of minimisation. The Computer Journal, 12(2):171–178, May 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/2/171.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/120171.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/
171.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/172.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/173.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/174.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/175.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/176.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/177.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/178.
tif.

Pedoe:1966:SMS

[Ped66] J. Pedoe. Sequential methods in statistics. The Computer Journal, 9(2):180b–180, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/2/180-b.full.pdf+html.

Penny:1966:ABT

[Pen66] J. P. Penny. An analysis, both theoretical and by simulation, of a time-shared computer system. The Computer Journal, 9(1): 53–59, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/1/53.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_01/090053.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/55.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/56.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/57.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/59.tif. See correspon-

dence [Bae66c].

Perry:1963:BR

[Per63] G. H. Perry. Book reviews. The Computer Journal, 6(3):295, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/295.full.pdf+html.

Perko:1965:SCN

[Per65] Aarni Perko. Some computational notes on the shortest route problem. The Computer Journal, 8(1):19–20, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/1/19.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/080019.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/20.tif. See correction

[Pap68a].

Phillips:1965:IG

[Phi65] William Phillips. “irascible genius”. The Computer Journal, 8 (1):56, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/56.full.pdf+html. See [OS65a].

Phillips:1966:PRT

[Phi66] D. C. Phillips. Programming real-time computer systems. The Computer Journal, 9(3):249, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/249.full.
pdf+html.

Phillips:1967:NRT

[Phi67a] C. S. E. Phillips. Networks for real-time programming. The Computer Journal, 10(1):46–52, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/46.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100046.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
46.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
49.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/51.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
52.tif.

Phillips:1967:NIT

[Phi67b] G. M. Phillips. Numerical integration in two and three dimensions. The Computer Journal, 10(2):202–204, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/202.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/100202.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
202.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/203.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/204.tif.

Phillips:1967:NID

[Phi67c] G. M. Phillips. Numerical integration over an N-dimensional rectangular region. The Computer Journal, 10(3):297–299, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/297.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/100297.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
03/tiff/297.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/298.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/299.
tif.

Phillips:1968:APS

[Phi68a] G. M. Phillips. Algorithms for piecewise straight line approximations. The Computer Journal, 11(2):211–212, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/211.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110211.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/211.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/212.tif. See correspondence

[BH69b].

Phillips:1968:EME

[Phi68b] G. M. Phillips. Estimate of the maximum error in best polynomial approximations. The Computer Journal, 11(1):110–111, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/110.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/110110.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
01/tiff/110.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/111.tif.

Pike:1967:NAF

[PHJ67] M. C. Pike, I. D. Hill, and F. D. James. Note on Algorithm 2: Fibonacci search, and on Algorithm 7: MINX. The Computer Journal, 9(4):416–417, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/416.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/417.tif. See [PP79].

Phillips:1965:C

[PI65] J. H. G. Phillips and M. Irish. Correspondence. The Computer Journal, 8(3):215, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/3/215.3.full.pdf+html.

Piggott:1964:BR

[Pig64] R. S. Piggott. Book review. The Computer Journal, 7(3):176, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
3/176.full.pdf+html.

Pinkerton:1961:EDS

[Pin61] J. M. M. Pinkerton. The evolution of design in a series of computers, LEO I–III. The Computer Journal, 4(1):42– 46, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/42.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/040042.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/45.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/46.
tif.

Pitteway:1965:ODL

[Pit65] M. L. V. Pitteway. The organization of a data link computing service. The Computer Journal, 8(3):190–200, October 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/190.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080190.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
190.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/191.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/192.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/193.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/194.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/

tiff/195.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/196.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/197.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/198.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/199.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/200.tif.

Pitteway:1967:ADE

[Pit67] M. L. V. Pitteway. Algorithm for drawing ellipses or hyperbolae with a digital plotter. The Computer Journal, 10(3):282–289, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/3/282.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/100282.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/282.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/283.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_03/tiff/284.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/285.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/286.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/287.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/288.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/289.tif.

Poole:1968:DLC

[PL68] P. C. Poole and T. Lang. The development of on-line computing facilities for the KDF9 part 1; COSEC — a single on-line console. The Computer Journal, 11(1):5–11, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/11/1/5.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_01/110005.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/5.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_01/tiff/6.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_01/tiff/7.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_11/Issue_01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
9.tif.

Platt:1961:EAC

[Pla61] A. J. Platt. The experience of applying a commercial computer in a British organization. The Computer Journal, 3(4):185–197, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/4/185.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_04/030185.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/185.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/186.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/187.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/188.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/189.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/190.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/191.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/192.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/193.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/194.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/195.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/196.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/197.tif.

Playfair:1964:CP

[Pla64] E. Playfair. Computers and psychology. The Computer Journal, 7 (1):1–3, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/1/1.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_01/070001.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/1.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/2.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/3.tif.

Proctor:1962:PSR

[PM62] W. G. Proctor and M. F. Mitchell. The PACE scaling routine for Mercury. The Computer Journal, 5(1):24–27, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/1/24.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_01/050024.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
01/tiff/26.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/27.tif.

Polley:1960:PRI

[Pol60] D. W. Polley. A progress report on the introduction of A.D.P. for recording contributions paid under the new graduated pensions scheme. The Computer Journal, 3(3):117–120, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/3/
3/117.full.pdf+html.

Powell:1962:IMF

[Pow62] M. J. D. Powell. An iterative method for finding stationary values of a function of several variables. The Computer Journal, 5(2):147–151, August 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/2/147.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/050147.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/tiff/
147.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_02/tiff/148.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_02/tiff/149.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_02/tiff/150.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_02/tiff/151.tif.

Powell:1964:EMF

[Pow64] M. J. D. Powell. An efficient method for finding the minimum of a function of several variables without calculating derivatives. The Computer Journal, 7(2):155–162, July 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/155.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070155.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
155.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/156.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/157.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/tiff/158.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/159.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/
tiff/160.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/161.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/162.
tif.

Powell:1965:MMS

[Pow65a] M. J. D. Powell. A method for minimizing a sum of squares of non-linear functions without calculating derivatives. The Computer Journal, 7(4):303–307, January 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/303.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070303.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
303.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/304.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/305.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/306.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/307.tif.

Power:1965:BRB

[Pow65b] A. Power. Book review: Automation in Bankwesen. The Computer Journal, 7(4):274, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/4/274.full.pdf+html.

Powell:1967:MEP

[Pow67] M. J. D. Powell. On the maximum errors of polynomial approximations defined by interpolation and by least squares criteria. The Computer Journal, 9(4):404–407, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/404.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090404.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
404.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/405.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/406.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/407.tif.

Powell:1968:COV

[Pow68] M. J. D. Powell. On the calculation of orthogonal vectors. The Computer Journal, 11(3):302–304, November 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/302.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110302.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
302.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/303.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/304.tif.

See correspondence [And69b].

Powell:1969:NPS

[Pow69a] M. J. D. Powell. Nonlinear programming-sequential unconstrained minimization techniques. The Computer Journal, 12(3):207a–207, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/3/207-a.full.pdf+html.

Powell:1969:TRO

[Pow69b] M. J. D. Powell. A theorem on rank one modifications to a matrix and its inverse. The Computer Journal, 12(3):288–290, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
3/288.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/120288.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/288.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/289.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/290.
tif.

Pitteway:1969:NAB

[PP69] M. L. V. Pitteway and R. W. Parry. Note on Algorithm 35: Best rational approximation to a real number. The Computer Journal, 12 (3):293, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/293.tif. See [Pre68,

Cob69].

Pike:1979:AFS

[PP79] M. C. Pike and J. Pixner. Algorithm 2: Fibonacci search. The Computer Journal, 22(1):86, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/

86.tif. Reprinted from The Computer Bulletin, vol. 8, p. 147, 1965. See [Boo79, Ove67c, PHJ67].

Parker-Rhodes:1964:CA

[PR64] A. F. Parker-Rhodes. The communication of algorithms. The Computer Journal, 7(1):28–35, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/1/28.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070028.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/29.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/30.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/31.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/32.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/33.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/
34.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/35.tif.

Prescott:1968:ABR

[Pre68] R. J. Prescott. Algorithm 35: Best rational approximation to a real number. The Computer Journal, 11(3):347–348, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/347.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/348.

tif. See note [PP69, Cob69].

Prinz:1960:C

[Pri60] D. G. Prinz. Correspondence. The Computer Journal, 3(2):66, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/3/
2/66.full.pdf+html.

Prinz:1961:BR

[Pri61] D. G. Prinz. Book review. The Computer Journal, 4(1):24, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
1/24.2.full.pdf+html.

Prinz:1963:BRD

[Pri63] D. G. Prinz. Book review: Dynamic programming. The Computer Journal, 6(2):201, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/6/2/201.full.pdf+html.

Price:1964:DCF

[Pri64] V. E. Price. Desirable computing facilities for a College of Advanced Technology. The Computer Journal, 7(3):177–179, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/3/177.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_03/070177.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/177.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/179.tif.

Price:1966:NMC

[Pri66] V. E. Price. Numerical methods and computers. The Computer Journal, 9(2):199, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/2/199.full.pdf+html.

Price:1967:BR

[Pri67] V. E. Price. Book review. The Computer Journal, 10(1):59, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/59.full.pdf+html.

Price:1968:OVI

[Pri68a] C. F. Price. An offset vector iteration method for solving twopoint boundary-value problems. The Computer Journal, 11(2): 220–228, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/2/220.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/110220.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/220.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/221.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/222.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/223.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/224.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/225.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/226.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/227.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/228.tif.

Price:1968:BR

[Pri68b] V. E. Price. Book review. The Computer Journal, 11(3):304, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/304.full.pdf+html.

Pullin:1964:FAT

[Pul64] D. Pullin. A Fortran to Algol translator. The Computer Journal, 7 (1):24–27, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/1/24.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_01/070024.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/26.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/27.tif.

Pullen:1967:CAT

[PW67] H. G. M. Pullen and M. H. J. Webb. A computer application to a transport scheduling problem. The Computer Journal, 10(1):10– 13, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/1/10.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_10/Issue_01/100010.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/10.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/11.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
12.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/13.tif.

Peters:1969:EBS

[PW69] G. Peters and J. H. Wilkinson. Eigenvalues of Ax = λBx with band symmetric A and B. The Computer Journal, 12(4):398– 404, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/398.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/120397.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/398.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/399.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/400.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/401.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/402.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/403.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/404.
tif.

Rachman:1965:CRE

[Rac65] A. Rachman. Computation of rotational energy levels of rigid asymmetric top molecules. The Computer Journal, 8(2):147–149, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/2/147.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/080147.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/147.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_02/tiff/148.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_02/

tiff/149.tif. See [Duk66b, Jon66].

Ralston:1961:STC

[Ral61] A. Ralston. Some theoretical and computational matters relating to predictor-corrector methods of numerical integration. The Computer Journal, 4(1):64–67, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/1/64.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/040064.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/64.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/65.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/66.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/67.
tif.

Randall:1968:NGF

[Ran68a] T. J. Randall. Note on a general finite-difference formula for the solution of axially symmetric fields. The Computer Journal, 11(3):332–333, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/11/3/332.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/110332.sgm.abs.html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/332.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_03/tiff/333.tif.

Randall:1968:NEO

[Ran68b] T. J. Randall. A note on the estimation of the optimum successive overrelaxation parameter for Laplace’s equation. The Computer Journal, 10(4):400–401, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/400.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100400.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
400.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/401.tif.

Rashba:1966:AMT

[Ras66] J. B. Rashba. The application of matrix theory to electrical engineering. The Computer Journal, 9(3):293, November 1966.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
3/293.full.pdf+html.

Redfern:1960:EUD

[Red60] P. Redfern. Experience in using a Deuce computer for the family expenditure survey. The Computer Journal, 2(4):164–169, January 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/2/4/164.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_02/Issue_04/020164.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_04/tiff/164.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_02/Issue_04/tiff/165.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_02/Issue_04/
tiff/166.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_02/Issue_04/tiff/167.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_02/Issue_04/tiff/168.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
02/Issue_04/tiff/169.tif.

Redish:1971:CNA

[Red71] K. A. Redish. Correspondence: (note on) Algorithm 44. The Computer Journal, 14(1):64, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
140064.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/64.tif. See [Bro69c,

Fie70, HL70].

Reeves:1967:DSD

[Ree67] C. M. Reeves. Description of a syntax-directed translator. The Computer Journal, 10(3):244–255, November 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/3/244.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/100244.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
244.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/245.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/246.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_03/tiff/247.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_10/Issue_03/tiff/248.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/249.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/250.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/251.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/252.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/253.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/254.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/255.tif.

Reid:1966:MFO

[Rei66] J. K. Reid. A method for finding the optimum successive overrelaxation parameter. The Computer Journal, 9(2):200–204, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/2/200.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_02/090200.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/201.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/203.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/204.
tif.

Reid:1967:NLS

[Rei67] J. K. Reid. A note on the least squares solution of a band system of linear equations by Householder reductions. The Computer Journal, 10(2):188–189, August 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/188.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/100188.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
188.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/189.tif.

Reid:1969:BRB

[Rei69] J. K. Reid. Book review: Semi-Groups of Operators and Approximation. The Computer Journal, 12(1):71, February 1969.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
1/71.full.pdf+html.

Richards:1963:BR

[Ric63] B. Richards. Book reviews. The Computer Journal, 6(3):263, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/263.full.pdf+html.

Richardson:1964:UEC

[Ric64] M. Richardson. User’s experience of Cleo. The Computer Journal, 7 (2):102–103, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/2/102.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/070102.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/102.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/103.tif.

Rigg:1964:OLC

[Rig64] F. A. Rigg. The operation of large computer systems. The Computer Journal, 7(3):169–172, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/169.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070169.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
169.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/170.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/171.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/172.tif.

Rivlin:1968:PAM

[Riv68] T. J. Rivlin. Polynomial approximation and the τ-method. The Computer Journal, 11(3):337–338, November 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/337.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110337.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/

337.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/338.tif.

Robertson:1969:NNI

[RM69] H. H. Robertson and M. J. McCann. A note on the numerical integration of conservative systems of first-order ordinary differential equations. The Computer Journal, 12(1):81, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/1/81.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_12/Issue_01/120081.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/81.tif.

Robertson:1963:BR

[Rob63a] H. H. Robertson. Book review. The Computer Journal, 5(4):330– 331, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/4/330.full.pdf+html.

Robinson:1963:BPA

[Rob63b] F. D. Robinson. The background of the PERT algorithm. The Computer Journal, 5(4):297–300, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/297.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050297.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
297.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/298.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/299.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/300.tif.

Roberts:1967:BR

[Rob67] J. D. Roberts. Book review. The Computer Journal, 10(1):107, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/107.full.pdf+html.

Rogers:1965:BR

[Rog65] M. H. Rogers. Book review. The Computer Journal, 8(3):277, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
3/277.full.pdf+html.

Rohl:1968:NBN

[Roh68] J. S. Rohl. A note on Backus Naur Form. The Computer Journal, 10(4):336–337, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/336.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100336.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
336.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/337.tif.

Rosenbrock:1960:AMF

[Ros60] H. H. Rosenbrock. An automatic method for finding the greatest or least value of a function. The Computer Journal, 3(3):175–184, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/3/175.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_03/030175.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_03/tiff/175.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/176.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/
tiff/177.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_03/tiff/178.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/179.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_03/tiff/180.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/181.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/
tiff/182.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_03/tiff/183.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/184.
tif.

Rosenbrock:1962:BR

[Ros62] H. H. Rosenbrock. Book review. The Computer Journal, 5(1): 70, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/1/70.full.pdf+html.

Rosenbrock:1963:SGI

[Ros63a] H. H. Rosenbrock. Some general implicit processes for the numerical solution of differential equations. The Computer Journal, 5(4): 329–330, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/4/329.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/050329.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/329.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/330.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/331.tif.

Ross:1963:CDL

[Ros63b] H. McG. Ross. A convention to distinguish letter O from numeral zero. The Computer Journal, 6(1):49, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/
6/1/49.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/060049.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/49.tif.

Rose:1964:CSE

[Ros64a] M. J. Rose. Classification of a set of elements. The Computer Journal, 7(3):208–211, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/208.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070208.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
208.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/209.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/210.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/211.tif.

Ross:1964:CC

[Ros64b] H. McG. Ross. The I.S.O. character code. The Computer Journal, 7(3):197–202, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/197.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_07/Issue_03/070197.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
197.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/198.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/199.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/201.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/202.tif.

Rosenbrock:1966:CSC

[Ros66a] H. H. Rosenbrock. Correspondence: On the stability of Crank and Nicolson’s procedure. The Computer Journal, 9(3):320, 324, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/320.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090320.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/320.tif. See [PC64, KM66].

Rosin:1966:AES

[Ros66b] R. F. Rosin. An approach to executive system maintenance in disk-based systems. The Computer Journal, 9(3):242–247, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/242.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090242.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/242.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/243.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/244.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/245.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/246.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/247.tif.

Rosenthal:1967:AAM

[Ros67] L. E. Rosenthal. Analytic algebraic manipulation. The Computer Journal, 10(3):265–270, November 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/3/265.full.

pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/100265.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
265.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/266.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/267.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_03/tiff/268.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/269.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/270.tif.

Ross:1968:BR

[Ros68] G. J. S. Ross. Book review. The Computer Journal, 11(1):11, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/11.full.pdf+html.

Rousell:1962:PRN

[Rou62] A. R. Rousell. A progress report on NEBULA. The Computer Journal, 5(3):162–163, November 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/162.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/050162.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/
162.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/163.tif.

Richards:1963:BRB

[RP63] B. Richards and E. S. Page. Book reviews: Annual Review in Automatic Programming, Volume 3, edited by R. Goodman; Proceedings of a Harvard Symposium on Digital Computers and Their Applications, edited by A. G. Oettinger, 1962. The Computer Journal, 6(3):263–263, 270–270, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/263.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/270.tif.

Rechard:1969:ETA

[RS69] Ottis W. Rechard and Richard H. Stark. Equivalence of two algorithms for Cooper’s generalised factorial function. The Computer Journal, 12(1):33–37, February 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/33.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120033.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
33.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/34.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/35.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
36.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/37.tif.

Rudall:1967:AC

[Rud67a] B. H. Rudall. Algorithm 19: Complex. The Computer Journal, 10 (1):112–113, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/112.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/
tiff/113.tif.

Rudall:1967:NAC

[Rud67b] B. H. Rudall. Note on Algorithm 19: Complex. The Computer Journal, 10(2):208–209, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/208.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/209.tif.

Russell:1963:PDT

[Rus63] D. G. Russell. The problems of data transmission systems in a general manufacturing data processing installation. The Computer Journal, 6(3):210–213, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/210.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060210.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
210.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/211.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/212.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/213.tif.

Rutishauser:1963:C

[Rut63] H. Rutishauser. Correspondence. The Computer Journal, 6(2): 133, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/2/133.full.pdf+html.

Rollett:1961:ESC

[RW61] J. S. Rollett and J. H. Wilkinson. An efficient scheme for the co-diagonalization of a symmetric matrix by Givens’ method in a computer with a two-level store. The Computer Journal, 4(2):177– 180, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/2/177.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_02/040177.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/177.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/180.tif.

Rollett:1962:LEG

[RW62] J. S. Rollett and J. H. Wilkinson. Letter to the Editor: [Givens and Householder processes]. The Computer Journal, 4(4):279, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/cgi/
reprint/4/4/279;   http://comjnl.oxfordjournals.org/content/
4/4/279.full.pdf+html.

Ranyard:1967:OAT

[RW67] J. C. Ranyard and A. Wren. The optimum arrangement of towers in an electric power transmission line. The Computer Journal, 10(2):157–161, August 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/2/157.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/100157.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/tiff/
157.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_02/tiff/158.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_02/tiff/159.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/

Issue_02/tiff/160.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/161.tif.

Riley:1968:GTD

[RW68] J. L. Riley and P. H. Walker. Generation of time delays on analogue computers. The Computer Journal, 11(1):72–76, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/72.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110072.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/73.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
74.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/75.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/76.tif.

Ridgion:1964:ACS

[RWT64] J. M. Ridgion, A. J. Willmott, and J. H. Thewlis. An analogue computer simulation of a Cowper Stove. The Computer Journal, 7(3):188–196, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/188.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070188.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
188.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/189.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/190.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/191.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/192.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/193.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_03/tiff/194.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/195.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/196.tif.

Ryan:1963:LFT

[Rya63] W. S. Ryan. LEAPS — the first three years. The Computer Journal, 6(1):6–13, April 1963. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/6/1/6.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060006.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/13.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/6.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/
8.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/9.tif.

Ryan:1966:DIA

[Rya66] Joseph T. Ryan. A direction-independent algorithm for determining the forward and backward compute point for a term or subscript during compilation. The Computer Journal, 9(2):157–160, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/2/157.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_02/090157.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/157.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/158.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/160.tif.

Ryder:1964:NAC

[Ryd64] K. L. Ryder. Note on an Algol 60 compiler for Pegasus I. The Computer Journal, 6(4):336–338, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/336.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060336.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
336.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/337.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/338.tif.

Sack:1962:NCT

[Sac62] R. A. Sack. Newton–Cotes type quadrature formulas with terminal corrections. The Computer Journal, 5(3):230–237, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/230.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_03/050230.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/231.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/232.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/233.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/234.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/236.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/237.tif.

Sale:1968:CMS

[Sal68] A. H. J. Sale. The calculation of e to many significant digits. The Computer Journal, 11(2):229–230, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/2/229.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/110229.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/
229.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/230.tif.

Sale:1970:NAI

[Sal70] A. H. J. Sale. Note on Algorithm 42: Interpolation by certain quintic splines. The Computer Journal, 13(1):115, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/115.tif. See [Spä69b].

Samet:1963:BR

[Sam63a] P. A. Samet. Book reviews. The Computer Journal, 6(2):208, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/2/208.full.pdf+html.

Samet:1963:C

[Sam63b] Paul A. Samet. Correspondence. The Computer Journal, 6(3): 286, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/286.full.pdf+html.

Samet:1963:CHR

[Sam63c] Paul A. Samet. Correspondence: a hardware representation for ALGOL 60 using Creed Teleprinter equipment. The Computer Journal, 6(3):286, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/286.

tif. See correspondence [Bar64, Tay64].

Samet:1965:EAB

[Sam65] P. A. Samet. The efficient administration of blocks in Algol. The Computer Journal, 8(1):21–23, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/8/1/21.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_01/080021.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/21.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/22.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/23.tif.

Samet:1966:CES

[Sam66a] Paul A. Samet. Correspondence: On “The Egdon system for the KDF9”. The Computer Journal, 9(2):174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/174.tif. See [BHJV66] and corre-

spondence [Bar66a].

Samet:1966:P

[Sam66b] Paul A. Samet. Progress? The Computer Journal, 9(2):174a– 174, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/2/174-a.full.pdf+html.

Samet:1968:ABR

[Sam68a] P. A. Samet. Appeal from the Book Review Editor. The Computer Journal, 11(3):301, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/11/3/301.full.pdf+html.

Samet:1968:BR

[Sam68b] P. A. Samet. Book review. The Computer Journal, 10(4):399, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/399.full.pdf+html.

Samet:1968:DCS

[Sam68c] P. A. Samet. Discussion and correspondence: software requirements of universities. The Computer Journal, 11(2):236–240, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/236.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110236.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/236.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/237.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/238.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/239.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/240.tif.

Savage:1969:JOT

[Sav69] M. J. Savage. Junction optimisation technique. The Computer Journal, 12(3):268–272, August 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/3/268.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/120268.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/
268.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/269.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/270.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/271.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/272.tif. See erratum

[Sav70].

Savage:1970:EJO

[Sav70] M. J. Savage. Erratum: “Junction optimisation technique”. The Computer Journal, 13(1):120, February 1970. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/120.tif. See [Sav69].

Schorr:1965:ADU

[Sch65] Herbert Schorr. Analytic differentiation using a syntax-directed compiler. The Computer Journal, 7(4):290–298, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
4/290.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/070290.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
04/tiff/290.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/291.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/292.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_04/tiff/293.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/294.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/295.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/296.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/297.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_04/tiff/298.tif.

Scowen:1969:NAS

[Sco69] R. S. Scowen. Note on Algorithm 40. spline interpolation of degree three. The Computer Journal, 12(4):409, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/408.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/409.

tif. See [Spä69a].

Scraton:1964:C

[Scr64a] R. E. Scraton. Correspondence. The Computer Journal, 7(2): 136, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/2/136.full.pdf+html.

Scraton:1964:ETE

[Scr64b] R. E. Scraton. Estimation of the truncation error in Runge–Kutta and allied processes. The Computer Journal, 7(3):246–248, October 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/3/246.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_03/070246.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_03/tiff/246.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/247.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/
tiff/248.tif.

Scraton:1964:NSS

[Scr64c] R. E. Scraton. The numerical solution of second-order differential equations not containing the first derivative explicitly. The Computer Journal, 6(4):368–370, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/368.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060368.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
368.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/369.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/370.tif.

Scraton:1965:SLD

[Scr65]  R. E. Scraton.  The solution of linear differential equa-

tions in Chebyshev series. The Computer Journal, 8(1):57– 61, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/57.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/080057.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/57.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/58.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/59.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/60.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/61.
tif.

Swinnerton-Dyer:1962:CPS

[SD62] H. P. F. Swinnerton-Dyer. The calculation of power spectra. The Computer Journal, 5(1):16–23, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/5/1/16.full.pdf+html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_01/050016.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/16.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
01/tiff/17.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/19.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/20.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/21.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
22.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/23.tif.

Seaton:1969:DCS

[Sea69] M. J. Seaton. Diagonalisation of complex symmetric matrices using a modified Jacobi method. The Computer Journal, 12(2): 156–157, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/156.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120156.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/156.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/157.tif.

Seldon:1972:CPS

[Sel72] J. L. Seldon. Correspondence: The postage stamp problem. The Computer Journal, 15(4):361, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/361.tif. See [Lun69].

Strachey:1961:RMC

[SF61] C. Strachey and J. G. F. Francis. The reduction of a matrix to codiagonal form by eliminations. The Computer Journal, 4(2):168– 176, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/4/2/168.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_02/040168.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/169.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/

tiff/170.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/171.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/172.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/173.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/174.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/175.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/176.tif.

Sharman:1965:CR

[Sha65] F. H. Sharman. Character recognition. The Computer Journal, 8 (2):89–94, July 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/2/89.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_02/080089.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_02/tiff/89.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_02/tiff/90.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
02/tiff/91.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_02/tiff/92.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_02/tiff/93.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_02/tiff/94.tif.

Sharp:1967:BR

[Sha67] J. Sharp. Book review. The Computer Journal, 10(1):13, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
1/13.full.pdf+html.

Simpson:1969:CFO

[Sim69] H. R. Simpson. A compact form of one-track syntax analyser. The Computer Journal, 12(3):233–243, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/3/233.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/120233.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/
233.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/234.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/235.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_03/tiff/236.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/237.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/238.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/239.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/240.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/241.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/242.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/243.tif.

Sipala:1969:FDA

[Sip69] P. Sipala. Formatting and display of Algol expressions. The Computer Journal, 12(4):365–367, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/365.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120365.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
365.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/366.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/367.tif.

Slater:1960:PIL

[Sla60] L. D. Slater. Problems of the introduction of large scale data processing into the Royal Army Pay Corps. The Computer Journal, 3(3):120–123, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/3/3/120.full.pdf+html.

Slater:1962:BR

[Sla62a] L. J. Slater. Book review. The Computer Journal, 4(4):312, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
4/312.full.pdf+html.

Slater:1962:RA

[Sla62b] L. J. Slater. Regression analysis. The Computer Journal, 4(4):287– 291, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/4/287.full.pdf+html;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_04/Issue_04/040287.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/287.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_04/tiff/288.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/289.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/290.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/291.tif.

Slater:1964:DPP

[Sla64a] L. J. Slater. A dynamic programming process. The Computer Journal, 7(1):36–39, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/1/36.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070036.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/36.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/37.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/38.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/39.
tif.

Slater:1964:SEP

[Sla64b] Lucy Joan Slater. Some experiences in price mapping. The Computer Journal, 6(4):348–351, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/348.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060348.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
348.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/349.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/350.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_04/tiff/351.tif.

Slater:1966:CSE

[Sla66] Lucy Joan Slater. Computing the state of the economy. The Computer Journal, 9(1):11–15, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/9/1/11.full.pdf+html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_01/090011.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/tiff/11.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/13.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_01/tiff/14.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/15.tif.

Smith:1963:BR

[Smi63a] K. L. Smith. Book reviews. The Computer Journal, 6(3):297, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/6/3/297.1.full.pdf+html.

Smith:1963:VDT

[Smi63b] K. L. Smith. The views of the data transmission committee. The Computer Journal, 6(3):222–227, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/222.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060222.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
222.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/223.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/224.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/225.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/226.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/227.tif.

Smith:1964:CDU

[Smi64] C. R. Smith. Computer development from a user’s point of view. The Computer Journal, 7(2):89–92, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/2/89.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_02/070089.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_02/tiff/89.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
02/tiff/90.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_07/Issue_02/tiff/91.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/92.
tif.

Smith:1966:LEC

[Smi66a] A. C. Smith. A lower estimate of the cumulative truncation error in Milne’s method. The Computer Journal, 8(4):395–397, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/395.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_04/080395.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/395.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/396.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/397.tif.

Smith:1966:SNC

[Smi66b] K. L. Smith. Signals and noise in communication systems. The Computer Journal, 9(3):274, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/274.full.
pdf+html.

Smith:1966:SAP

[Smi66c] Norman Smith. Systems analysis problems encountered during a large computer application. The Computer Journal, 9(3):230–237, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/3/230.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_03/090230.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/231.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/232.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_03/tiff/233.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/234.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_03/tiff/235.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/236.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/237.tif.

Smith:1968:SCE

[Smi68] Bernard Smith. SPECOL — a computer enquiry language for the non-programmer. The Computer Journal, 11(2):121–127, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/121.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110121.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/121.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/122.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/123.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/124.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/125.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/127.tif.

Sneath:1966:MCS

[Sne66] P. H. A. Sneath. A method for curve seeking from scattered points. The Computer Journal, 8(4):383–391, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/383.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080383.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
383.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/384.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/385.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_04/tiff/386.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/387.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/388.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_04/tiff/389.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/390.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_04/tiff/391.tif.

Solomon:1960:CPN

[Sol60] E. W. Solomon. A comprehensive program for network problems. The Computer Journal, 3(2):89–97, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/2/89.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_02/030089.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_02/tiff/89.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
02/tiff/90.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/91.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/92.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/93.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/94.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/
95.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/96.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/97.tif.

Solomon:1968:C

[Sol68a] E. W. Solomon. Correspondence. The Computer Journal, 11(3): 352, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/11/3/352.1.full.pdf+html.

Solomon:1968:CBD

[Sol68b] Eric W. Solomon. Correspondence: Basic data input routines. The Computer Journal, 11(3):352, November 1968. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/352.tif.

Spath:1968:AFD

[Spä68] H. Späth. Algorithm 33: Fitting data to an exponential with a straight line as background. The Computer Journal, 11(1):114–115, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_01/tiff/114.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/115.
tif.

Spath:1969:ASI

[Spä69a] H. Späth. Algorithm 40: Spline interpolation of degree three. The Computer Journal, 12(2):198–199, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/

tiff/198.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/199.tif. See note [Sco69].

Spath:1969:AIC

[Spä69b] H. Späth. Algorithm 42: Interpolation by certain quintic splines. The Computer Journal, 12(3):292–293, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/292.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/293.

tif. See note [Sal70].

Spencer:1965:BR

[Spe65] D. C. Spencer. Book review. The Computer Journal, 8(2):158, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
2/158.full.pdf+html.

Scraton:1963:NNS

[SS63a] R. E. Scraton and J. W. Searle. Note on the numerical solution of linear differential equations with constant coefficients. The Computer Journal, 6(2):206–207, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/206.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060206.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
206.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/207.tif.

Stroud:1963:MPF

[SS63b] A. H. Stroud and D. Secrest. A multiple-precision floatingpoint interpretive program for the Control Data 1604. The Computer Journal, 6(1):62–66, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/6/1/62.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_01/060062.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/tiff/62.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_
01/tiff/63.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_01/tiff/64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_01/tiff/65.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/66.tif.

Stearman:1966:ST

[Ste66] G. H. Stearman. Switching theory. The Computer Journal, 9(2): 143, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/2/143.full.pdf+html.

Stern:1967:OQF

[Ste67] M. D. Stern. Optimal quadrature formulae. The Computer Journal, 9(4):396–403, February 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/4/396.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090396.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
396.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/397.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_04/tiff/398.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_04/tiff/399.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_04/tiff/400.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/401.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/402.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/403.
tif.

Storey:1968:CFB

[Sto68] S. H. Storey. The convergence of Fourier–Bessel expansions. The Computer Journal, 10(4):402–405, February 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/4/402.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/100402.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/
402.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/403.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/404.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/405.tif.

Strachey:1960:TCT

[Str60] C. Strachey. Two contributions to the techniques of queuing problems. The Computer Journal, 3(2):114–116, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/2/114.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/030114.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/
114.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_02/tiff/115.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_02/tiff/116.tif.

Stringer:1961:PCR

[Str61] J. B. Stringer. The place of character recognition, data transmission and document handling in A.D.P. systems. The Computer Journal, 4(2):161–167, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/161.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040161.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
161.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/162.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/163.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/164.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/165.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/166.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/167.tif.

Strachey:1962:BR

[Str62] Christopher Strachey. Book reviews. The Computer Journal, 5(2): 152–153, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/2/152.full.pdf+html.

Stringer:1964:BR

[Str64] J. B. Stringer. Book review. The Computer Journal, 6(4):303, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/4/303.full.pdf+html.

Strachey:1965:Ca

[Str65a] C. Strachey. Correspondence. The Computer Journal, 8(2):176, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/2/176.1.full.pdf+html.

Strachey:1965:Cb

[Str65b] C. Strachey. Correspondence. The Computer Journal, 8(3):215, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/3/215.2.full.pdf+html.

Strachey:1965:GPM

[Str65c] C. Strachey. A general purpose macrogenerator. The Computer Journal, 8(3):225–241, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/225.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080225.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
225.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/tiff/226.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/227.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/228.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/229.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/230.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/231.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/232.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/233.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/234.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/235.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/236.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/237.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/239.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/
tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_03/tiff/241.tif.

Strachey:1965:IP

[Str65d] C. Strachey. An impossible program. The Computer Journal, 7(4): 313, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/4/313.full.pdf+html.

Strachey:1965:LEI

[Str65e] C. Strachey. Letter to the Editor: An impossible program. The Computer Journal, 7(4):313, January 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Sumner:1965:RCP

[Sum65] F. H. Sumner. Reliable computation in the presence of noise. The Computer Journal, 8(1):76, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/76.full.pdf+html.

Sutton:1960:FYE

[Sut60] R. L. Sutton. The first year’s experience with a large computer in a life assurance office. The Computer Journal, 3(1): 2–9, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/1/2.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_01/030002.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/2.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/3.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/4.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/5.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_01/tiff/6.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/tiff/7.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/
9.tif.

Shepherd:1968:CAA

[SW68] M. J. Shepherd and A. J. Willmott. Cluster analysis on the Atlas computer. The Computer Journal, 11(1):57–62, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/1/57.full.

pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_01/110057.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
57.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/59.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
60.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/62.tif.

Taylor:1961:EDC

[Tay61a] B. R. Taylor. Experience with a digital computer in an aeroplane testing establishment. The Computer Journal, 4(1):25– 29, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/25.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/040025.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/25.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/26.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/27.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/28.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/29.
tif.

Taylor:1961:BR

[Tay61b] P. Taylor. Book review. The Computer Journal, 3(4):219, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/3/
4/219.1.full.pdf+html.

Taylor:1964:CRA

[Tay64] P. Taylor. Correspondence: On the representation of AL-GOL. The Computer Journal, 6(4):335–335, 367–367, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/335.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/367.

tif. See [Sam63c, Bar64].

Taylor:1968:C

[Tay68a] R. Taylor. Correspondence. The Computer Journal, 10(4):388, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/388.full.pdf+html.

Taylor:1968:COF

[Tay68b] R. Taylor. Correspondence: Orion FORTRAN compiler. The Computer Journal, 10(4):388, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
100388.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/388.tif. See [TH64,

Wil67d].

Thorby:1963:AEC

[TB63] R. P. Thorby and B. Benjamin. The application of the electronic computer to the 1961 population census of Great Britain. The Computer Journal, 5(4):264–270, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/264.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050264.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
264.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/265.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/266.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/268.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/269.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/270.tif.

Tee:1963:BR

[Tee63a] G. J. Tee. Book reviews. The Computer Journal, 6(3):276, November 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/3/276.2.full.pdf+html.

Tee:1963:ESR

[Tee63b] G. J. Tee. Eigenvectors of the successive over-relaxation process, and its combination with Chebyshev semi-iteration. The Computer Journal, 6(3):250–263, November 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/3/250.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/060250.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/
250.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_03/tiff/251.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_03/tiff/252.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_03/tiff/253.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/254.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/255.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/256.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/257.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_03/tiff/258.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_03/tiff/259.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_03/
tiff/260.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_03/tiff/261.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_03/tiff/262.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_03/tiff/263.tif. See correction [Ano64a].

Tee:1963:NFD

[Tee63c] G. J. Tee. A novel finite-difference approximation to the biharmonic operator. The Computer Journal, 6(2):177–192, August 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/2/177.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_02/060177.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/177.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/178.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/180.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/181.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/182.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/183.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/184.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/185.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/186.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/187.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/188.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/189.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/190.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/191.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_02/tiff/192.tif.

Tee:1967:CCV

[Tee67a] G. H. Tee. Correspondence: On coding values of mathematical constants. The Computer Journal, 9(4):393, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/090393.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/tiff/
393.tif.

Tee:1967:BR

[Tee67b] G. J. Tee. Book review. The Computer Journal, 9(4):395, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
4/395.full.pdf+html.

Tee:1967:CP

[Tee67c] G. J. Tee. Correspondence: π and pi. The Computer Journal, 9(4): 393, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/4/393.full.pdf+html.

Temes:1969:C

[Tem69a] G. C. Temes. Correspondence. The Computer Journal, 12(4): 410, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/410.1.full.pdf+html.

Temes:1969:CNA

[Tem69b] G. C. Temes. Correspondence: Note on “An algorithm for minimax approximation in the nonlinear case”. The Computer Journal, 12(4):410, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/410.

tif. See [OW69].

Tewarson:1967:RCP

[Tew67] R. P. Tewarson. Row-column permutation of sparse matrices. The Computer Journal, 10(3):300–305, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/3/300.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/100300.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
300.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/301.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/302.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_03/tiff/303.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/304.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/305.tif.

Tewarson:1968:CME

[Tew68] R. P. Tewarson. A computational method for evaluating generalized inverses. The Computer Journal, 10(4):411–413, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
4/411.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/100411.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/tiff/411.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/412.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/413.
tif.

Tewarson:1969:CRS

[Tew69a] R. P. Tewarson. The Crout reduction for sparse matrices. The Computer Journal, 12(2):158–159, May 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/2/158.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/120158.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/

158.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/159.tif.

Tewarson:1969:LSI

[Tew69b] R. P. Tewarson. A least squares iterative method for singular equations. The Computer Journal, 12(4):388–392, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/388.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/120388.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/tiff/388.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/389.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/390.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_04/tiff/391.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/392.tif.

Tewarson:1969:PMS

[Tew69c] R. P. Tewarson. Projection methods for solving sparse linear systems. The Computer Journal, 12(1):77–80, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/1/77.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_01/120077.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
77.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_01/tiff/78.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/79.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/tiff/
80.tif.

Taylor:1964:FSO

[TH64] R. Taylor and D. A. Harragan. The FORTRAN system for Orion. The Computer Journal, 7(2):114–116, July 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/2/114.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/070114.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/tiff/
114.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_02/tiff/115.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_02/tiff/116.tif.

See correspondence [Wil67d, Tay68b].

Thacher:1962:IMQ

[Tha62] Henry C. Thacher, Jr. An iterative method for quadratures. The Computer Journal, 5(3):228–229, November 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/3/228.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/050228.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/
228.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_03/tiff/229.tif.

Thacher:1966:CRI

[Tha66] Henry C. Thacher, Jr. Closed rational integration formulas. The Computer Journal, 8(4):362–367, January 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/4/362.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/080362.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/tiff/
362.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_04/tiff/363.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_04/tiff/364.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_04/tiff/365.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_04/tiff/366.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_04/
tiff/367.tif.

Thompson:1960:PAC

[Tho60] T. R. Thompson. Problems of auditing computing data: Internal audit practice and external audit theory: Section 1: Internal audit. The Computer Journal, 3(1):10–11, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/1/10.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/030010.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/11.tif.

Thomason:1961:IAC

[Tho61] J. G. Thomason. An introduction to analogue computer methods. The Computer Journal, 3(4):211–219, January 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/211.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030211.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
211.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/212.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/213.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/214.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/215.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/
tiff/216.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_04/tiff/217.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/218.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_04/tiff/219.tif.

Thompson:1962:FPE

[Tho62] T. R. Thompson. Fundamental principles of expressing a procedure for a computer application. The Computer Journal, 5(3):164–169, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/164.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_03/050164.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/164.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/165.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/166.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/167.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/168.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/169.tif.

Thomas:1964:RIU

[Tho64a] D. G. W. Thomas. Record identification using variable alphanumeric names. The Computer Journal, 7(1):19–21, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/1/19.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/070019.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_07/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/
21.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/23.tif.

Thompson:1964:C

[Tho64b] J. R. Thompson. Correspondence. The Computer Journal, 7(3): 227, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/7/3/227.full.pdf+html.

Thornton:1966:CTS

[Tho66] B. S. Thornton. Correspondence: Timetabling and scheduling problems. The Computer Journal, 9(1):66, 77, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
1/66.full.pdf+html.

Thomson:1968:CGT

[Tho68a] W. E. Thomson. Correspondence: Generation of time delays on analogue computers. The Computer Journal, 11(2):172, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/172.tif.

Thomson:1968:CSL

[Tho68b] W. E. Thomson. Correspondence: Solution of linear differential equation. The Computer Journal, 10(4):417–418, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/417.2.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/417.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/418.tif.

Thomson:1968:LEG

[Tho68c] W. E. Thomson. Letters to the editor: Generation of time delays on analogue computers. The Computer Journal, 11(2):172, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/172.full.pdf+html.

Throsby:1969:FEA

[Thr69] P. W. Throsby. A finite element approach to surface definition. The Computer Journal, 12(4):385–387, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/385.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120385.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
385.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/386.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/387.tif.

Tily:1969:AFD

[Til69] C. N. J. Tily. Algorithm 37: Fitting data to an exponentially damped linear function. The Computer Journal, 12(1):100–101, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_01/tiff/100.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_01/
tiff/101.tif.

Tizard:1960:CBA

[Tiz60] R. H. Tizard. Conversion between analogue and digital measures. The Computer Journal, 3(1):51–59, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/1/51.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_01/030051.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/51.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/52.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/53.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/54.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/55.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_01/tiff/56.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/
57.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_01/tiff/59.tif.

Tizard:1962:BR

[Tiz62] R. H. Tizard. Book review. The Computer Journal, 5(1):23, ????

1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
1/23.full.pdf+html.

Tizard:1964:BR

[Tiz64] R. H. Tizard. Book review. The Computer Journal, 7(3):196, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
3/196.full.pdf+html.

Tizard:1965:BR

[Tiz65] R. H. Tizard. Book review. The Computer Journal, 8(2):102, ???? 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/8/
2/102.full.pdf+html.

Tocher:1964:BR

[Toc64] K. D. Tocher. Book review. The Computer Journal, 7(3):211, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/7/
3/211.full.pdf+html.

Tompa:1967:GNI

[Tom67] H. Tompa. Gaussian numerical integration of a function depending on a parameter. The Computer Journal, 10(2):204–205, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
2/204.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/100204.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
02/tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_02/tiff/205.tif.

Traub:1966:NNN

[Tra66] J. F. Traub. Notice: Newsletter for Numerical Analysts. The Computer Journal, 8(4):340, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/4/340.full.pdf+html.

Trundle:1966:LAL

[Tru66] R. W. L. Trundle. LITHP — an ALGOL list processor. The Computer Journal, 9(2):167–172, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/167.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090167.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
167.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/168.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/169.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/170.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/171.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/172.tif.

Taylor-Smith:1964:ACI

[TS64] A. Taylor-Smith. Authorization and control of input in the Royal Army Pay Corps Computer Centre application. The Computer Journal, 7(1):14–18, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/7/1/14.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070014.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/16.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/17.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/18.tif.

Temperly:1968:GPP

[TS68] J. F. Temperly and Barry W. Smith. A grading procedure for PL/1 student exercises. The Computer Journal, 10(4):368–373, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
4/368.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/100368.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/tiff/368.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/369.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/370.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/371.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/372.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/373.tif.

Turnbull:1966:BRB

[Tur66] G. F. Turnbull. Book review: Sampling Systems Theory. The Computer Journal, 8(4):394, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/4/394.full.pdf+html.

Turski:1968:SDA

[Tur68] Wladyslaw M. Turski. SODA — a dual activity operating system. The Computer Journal, 11(2):148–156, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/2/148.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/110148.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/
148.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/149.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/150.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/151.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/152.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/153.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/154.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/155.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/156.tif.

Treweek:1961:BR

[TW61] K. H. Treweek and D. E. Williams. Book review. The Computer Journal, 4(2):156, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/4/2/156.full.pdf+html.

Thorp:1964:PAG

[TW64] E. O. Thorp and W. E. Walden. A partial analysis of Go. The Computer Journal, 7(3):203–207, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/203.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070203.sgm.abs.html;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
203.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/204.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/205.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/206.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/207.tif.

Ullmann:1965:BAP

[Ull65] J. R. Ullmann. A basic approach to pattern recognition. The Computer Journal, 7(4):282–289, January 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/282.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070282.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
282.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/283.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/284.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/285.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_04/tiff/286.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/
tiff/287.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_04/tiff/288.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/289.
tif.

Ullmann:1967:AIP

[Ull67] J. R. Ullmann. Algebraic inference of pattern similarity. The Computer Journal, 10(3):256–264, November 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/3/256.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/100256.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/
256.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/257.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/258.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_03/tiff/259.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_03/tiff/260.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_03/
tiff/261.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_10/Issue_03/tiff/262.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/263.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/264.tif.

Underhill:1963:GCG

[Und63] L. H. Underhill. The growth of complexity of a generalpurpose program. The Computer Journal, 6(1):37–38, April 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/1/37.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_01/060037.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_01/tiff/37.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_01/tiff/38.tif.

Vajda:1960:BR

[Vaj60] S. Vajda. Book review. The Computer Journal, 3(2):107, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/3/
2/107.full.pdf+html.

Vajda:1964:BRM

[Vaj64] S. Vajda. Book review: Mathematical programming. The Computer Journal, 7(2):103, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/103.full.pdf+html.

Vickery:1965:PTD

[Vic65] J. C. Vickery. Principles of tunnel dioded circuits. The Computer Journal, 8(1):20, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/1/20.full.pdf+html.

Vickery:1967:BRa

[Vic67a] B. C. Vickery. Book review. The Computer Journal, 10(3):226, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/226.1.full.pdf+html.

Vickery:1967:BRb

[Vic67b] B. C. Vickery. Book review. The Computer Journal, 10(3):226, ???? 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/226.2.full.pdf+html.

Vince:1964:DSD

[Vin64] P. H. Vince. Documentation for systems design. The Computer Journal, 7(1):8–14, April 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/7/1/8.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_01/070008.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_01/tiff/10.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_
01/tiff/11.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_01/tiff/12.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/13.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_01/tiff/14.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_01/tiff/8.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_01/tiff/
9.tif.

Wallis:1962:BR

[Wal62] P. J. Wallis. Book reviews. The Computer Journal, 5(3):227, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/3/227.2.full.pdf+html.

Wallace:1964:CRE

[Wal64] C. S. Wallace. Correlated round-off errors in digital integrating differential analyzers. The Computer Journal, 7(2):131–134, July 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/7/2/131.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_07/Issue_02/070131.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
07/Issue_02/tiff/131.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_02/tiff/132.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_07/Issue_02/
tiff/133.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_07/Issue_02/tiff/134.tif.

Walter:1969:MTM

[Wal69] W. Grey Walter. Modelling of thinking and the mind. The Computer Journal, 12(3):207b–207, August 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/12/3/207-b.full.pdf+
html.

Warmington:1960:FYP

[War60] C. B. Warmington. The first year’s production on a computer, and future plans. The Computer Journal, 3(3):124–127, ???? 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/3/
3/124.full.pdf+html.

Watt:1961:C

[Wat61] J. M. Watt. Correspondence. The Computer Journal, 4(2):167, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/4/
2/167.full.pdf+html.

Watt:1963:RAP

[Wat63] J. M. Watt. The realization of ALGOL procedures and designational expressions. The Computer Journal, 5(4):332–337, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/4/332.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/050332.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/332.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/333.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/334.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/335.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/336.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/337.tif.

Watt:1967:ASP

[Wat67] J. M. Watt. Algorithm 21: Solution of the polynomial equation using the method of Bairstow. The Computer Journal, 10(2):207– 208, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_02/tiff/207.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_02/
tiff/208.tif.

Watt:1968:BR

[Wat68a] J. M. Watt. Book review. The Computer Journal, 11(3):338, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/338.full.pdf+html.

Watt:1968:CSD

[Wat68b] J. M. Watt. Convergence and stability of discretization methods for functional equations. The Computer Journal, 11(1):77–82, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/77.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110077.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/78.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
79.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
82.tif.

Watson:1969:BRB

[Wat69] G. L. Watson. Book review: Indices and Primitive Roots. The Computer Journal, 12(1):23, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/1/23.full.pdf+html.

Watkins:2008:QAR

[Wat08] David S. Watkins. The QR algorithm revisited. SIAM Review, 50 (1):133–145, ???? 2008. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). See [Fra61, Fra62, Wat11].

Watkins:2011:FA

[Wat11] David S. Watkins. Francis’s algorithm. American Mathematical Monthly, 118(5):387–403, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/info/10.4169/amer.math.monthly.118.05.

387. See [Fra61, Fra62, Wat08].

Wallace:1968:IMC

[WB68] C. S. Wallace and D. M. Boulton. An information measure for classification. The Computer Journal, 11(2):185–194, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/185.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110185.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/185.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/186.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/187.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/188.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/189.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/190.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/191.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/192.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/193.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/194.tif.

Waston:1963:C

[WD63] G. H. L. Waston and G. M. Davis. Correspondence. The Computer Journal, 6(1):56, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/6/1/56.1.full.pdf+html.

Wegner:1962:ZAC

[Weg62] P. Wegner. Zero-address computers. The Computer Journal, 5 (1):15, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/1/15.full.pdf+html.

Wells:1970:CTF

[Wel70] M. Wells. Correspondence: Towards Fortran VI? The Computer Journal, 13(1):120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/120.

tif. See [HL70].

West:1964:BRA

[Wes64] John C. West. Book review: Automatic control. The Computer Journal, 7(2):134, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/134.full.pdf+html.

Westcott:1969:BRB

[Wes69] J. Westcott. Book review: Optimisation in Control and Practice. The Computer Journal, 12(1):9, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/12/1/9.full.pdf+html.

Wetherfield:1966:TPM

[Wet66] M. R. Wetherfield. A technique for program monitoring by interruption. The Computer Journal, 9(2):161–166, August 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/161.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090161.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
161.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/162.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/163.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/164.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/165.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/166.tif.

Wexler:1969:CHR

[Wex69] P. J. Wexler. Computers in humanistic research. The Computer Journal, 12(3):243, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/3/243.full.pdf+html.

Wiseman:1968:RSP

[WH68] N. E. Wiseman and J. O. Hiles. A ring structure processor for a small computer. The Computer Journal, 10(4):338–346, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
4/338.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/100338.sgm.abs.html;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/tiff/338.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/339.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/342.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/343.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/344.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/345.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/346.tif.

Wilkinson:1960:HMS

[Wil60] J. H. Wilkinson. Householder’s method for the solution of the algebraic eigenproblem. The Computer Journal, 3(1):23– 27, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/1/23.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_01/030023.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/23.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/24.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/25.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/26.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/27.
tif.

Wilkes:1961:BR

[Wil61a] M. V. Wilkes. Book review. The Computer Journal, 3(4):219, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/3/
4/219.2.full.pdf+html.

Wilkes:1961:DTN

[Wil61b] M. V. Wilkes. Data transmission and the new outlook for the computer field. The Computer Journal, 4(1):1–9, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/1.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/040001.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_04/Issue_01/tiff/1.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/2.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/3.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/4.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/tiff/5.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/tiff/6.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/
8.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/9.tif.

Wilkinson:1961:REB

[Wil61c] J. H. Wilkinson. Rigorous error bounds for computed eigensystems. The Computer Journal, 4(3):230–241, October 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/3/230.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/040230.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/
230.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_03/tiff/231.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_03/tiff/232.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_03/tiff/233.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/234.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/235.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/236.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_03/tiff/237.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_03/tiff/239.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_03/
tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_03/tiff/241.tif.

Williams:1961:PFF

[Wil61d] M. B. Williams. Present and future facilities for data transmission. The Computer Journal, 4(2):88–95, July 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/2/88.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/

Issue_02/040088.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_02/tiff/88.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
02/tiff/89.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/90.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/91.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/92.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/93.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
94.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_02/tiff/95.tif.

Wilkes:1962:BR

[Wil62a] M. V. Wilkes. Book review. The Computer Journal, 5(3):192, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/3/192.full.pdf+html.

Wilkinson:1962:BRB

[Wil62b] J. H. Wilkinson. Book review: Solutions Numériques des Équations Algébraiques, by E. Durand (Paris: Masson et Cie., 1960, pp. 328, 65 NF). The Computer Journal, 5(1):32, April 1962. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/5/1/28.full.pdf;
http://comjnl.oxfordjournals.org/content/5/1/32.full.pdf+
html.

Wilkinson:1962:IEM

[Wil62c] J. H. Wilkinson. Instability of the elimination method of reducing a matrix to tri-diagonal form. The Computer Journal, 5(1): 61–70, April 1962. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/5/1/61.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_01/050061.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_01/tiff/62.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_
01/tiff/63.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_01/tiff/64.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/65.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

05/Issue_01/tiff/66.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/67.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_01/tiff/
68.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_01/tiff/70.tif.

Wilkinson:1963:BRB

[Wil63a] J. H. Wilkinson. Book review: Solutions Numériques des Équations Algébraiques, Volume II, by E. Durand (Paris: Masson et Cie., 1961, pp. 445, 90 NF). The Computer Journal, 5(4):263, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/5/4/258.
full.pdf;   http://comjnl.oxfordjournals.org/content/5/
4/263.full.pdf+html.

Wilson:1963:ETA

[Wil63b] J. F. Wilson. Experience in transmitting accounting data. The Computer Journal, 5(4):305–307, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/305.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050305.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
305.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/306.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/307.tif.

Wilkes:1964:PUF

[Wil64a] Maurice V. Wilkes. A programmer’s utility filing system. The Computer Journal, 7(3):180–184, October 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/3/180.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/070180.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_03/tiff/
180.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_03/tiff/181.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_03/tiff/182.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_03/tiff/183.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_07/Issue_03/tiff/184.tif.

Williams:1964:PES

[Wil64b] J. W. J. Williams. E.S.P. the Elliott Simulator Package. The Computer Journal, 6(4):328–331, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/328.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060328.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
328.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/329.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/330.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_04/tiff/331.tif.

Williams:1964:BRT

[Wil64c] R. C. Williams. Book review: Threshold decoding. The Computer Journal, 7(2):130, ???? 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/7/2/130.full.pdf+html.

Wilkes:1965:LWT

[Wil65a] M. V. Wilkes. Lists and why they are useful. The Computer Journal, 7(4):278–281, January 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/7/4/278.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/070278.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_07/Issue_04/tiff/
278.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_07/Issue_04/tiff/279.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_07/Issue_04/tiff/280.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_07/
Issue_04/tiff/281.tif.

Wilkinson:1965:CLP

[Wil65b] J. H. Wilkinson. The calculation of Lamé polynomials. The Computer Journal, 8(3):273–277, October 1965. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/8/3/273.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_03/080273.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_03/tiff/
273.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_08/Issue_03/tiff/274.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_03/tiff/275.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_08/
Issue_03/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_03/tiff/277.tif.

Wilkinson:1965:CRA

[Wil65c] J. H. Wilkinson. Convergence of the LR, QR and related algorithms. The Computer Journal, 8(1):77–84, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/1/77.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/080077.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/78.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
79.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_08/Issue_01/tiff/
82.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
08/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_08/Issue_01/tiff/84.tif.

Wilkinson:1965:ARS

[Wil65d] J. H. Wilkinson. The QR algorithm for real symmetric matrices with multiple eigenvalues. The Computer Journal, 8(1):85– 87, April 1965. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/8/1/85.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_08/Issue_01/080085.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_08/Issue_01/tiff/85.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_08/Issue_01/tiff/86.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_08/Issue_
01/tiff/87.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_08/Issue_01/tiff/88.tif.

Wilkes:1966:BRB

[Wil66a] M. V. Wilkes. Book review: The Algebraic Eigenvalue Problem. The Computer Journal, 8(4):382, January 1966. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/8/4/382.
full.pdf+html.

Williams:1966:BRB

[Wil66b] D. Williams. Book review: Digital Communications. The Computer Journal, 8(4):314, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/8/4/314.full.pdf+html.

Wilkes:1967:DMA

[Wil67a] M. V. Wilkes. The design of multiple-access computer systems. The Computer Journal, 10(1):1–9, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/10/1/1.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_01/100001.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_10/Issue_01/tiff/1.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
01/tiff/2.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_01/tiff/3.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_01/tiff/4.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_01/tiff/5.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/6.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
7.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/8.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/9.tif.

Wilkinson:1967:CST

[Wil67b] J. C. Wilkinson. Computer simulation techniques. The Computer Journal, 10(2):142, August 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/10/2/142.full.pdf+html.

Wilkinson:1967:BRB

[Wil67c] J. H. Wilkinson. Book review: The Matrix Analysis of Vibration, by R. E. D. Bishop, G. M. L. Gladwell, and S. Michaelson, 1965; 404 pages (London: Cambridge University Press, 100s). The Computer Journal, 10(1):77, May 1967. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/10/1/74.full.pdf;

http://comjnl.oxfordjournals.org/content/10/1/77.full.
pdf+html.

Willey:1967:DCE

[Wil67d] E. L. Willey. Discussion and correspondence: an established U.K. software development. The Computer Journal, 10(3):313– 314, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/3/313.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/100313.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_03/tiff/313.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_03/tiff/314.tif.

See correspondence [Tay68b] and [TH64].

Wilkes:1968:OIS

[Wil68a] M. V. Wilkes. The outer and inner syntax of a programming language. The Computer Journal, 11(3):260–263, November 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/260.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_03/110260.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
03/tiff/260.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_03/tiff/261.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/262.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_03/tiff/263.tif.

Williams:1968:SLN

[Wil68b] H. P. Williams. The synthesis of logical nets consisting of NOR units. The Computer Journal, 11(2):173–176, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
2/173.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/110173.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_
02/tiff/173.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/174.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/175.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_02/tiff/176.tif.

Wilkes:1969:BR

[Wil69] M. V. Wilkes. Book review. The Computer Journal, 12(4):316, ???? 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/316.full.pdf+html.

Windley:1960:TFR

[Win60] P. F. Windley. Trees, forests and rearranging. The Computer Journal, 3(2):84–88, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/2/84.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_02/030084.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_02/tiff/84.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
02/tiff/85.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/86.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/87.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/88.tif.

Winrich:1969:NCE

[Win69] L. B. Winrich. Note on a comparison of evaluation schemes for the interpolating polynomial. The Computer Journal, 12(2):154– 155, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/2/154.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/120154.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/154.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/155.tif.

Wiseman:1961:BR

[Wis61] N. E. Wiseman. Book review. The Computer Journal, 4(3): 221, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/3/221.full.pdf+html.

Wiseman:1964:ALP

[Wis64] N. E. Wiseman. Application of list-processing methods to the design of interconnections for a fast logic system. The Computer Journal, 6(4):321–327, January 1964. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/4/321.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/060321.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/
321.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_04/tiff/322.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_04/tiff/323.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_04/tiff/324.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/325.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/
tiff/326.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/327.tif.

Wiseman:1968:NCD

[Wis68] N. E. Wiseman. A note on compiling display file from a data structure. The Computer Journal, 11(2):141–147, August 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/2/141.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/110141.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/tiff/
141.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_02/tiff/142.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_02/tiff/143.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_11/
Issue_02/tiff/144.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_02/tiff/145.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_11/Issue_02/
tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_11/Issue_02/tiff/147.tif.

Woodward:1961:AL

[WJ61] P. M. Woodward and D. P. Jenkins. Atoms and lists. The Computer Journal, 4(1):47–53, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/1/47.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/040047.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/47.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/48.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/49.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/50.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
04/Issue_01/tiff/51.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_01/tiff/52.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/
53.tif.

Windley:1960:DPU

[WKRJ60] P. F. Windley, L. R. Kay, and A. Rowland-Jones. Data processing in university administration. The Computer Journal, 3(1): 15–20, April 1960. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/3/1/15.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_01/030015.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_01/tiff/15.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_01/tiff/16.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
01/tiff/17.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_01/tiff/18.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_01/tiff/19.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_01/tiff/20.tif.

Williams:1969:WLC

[WM69] P. E. Williams and D. C. Menzies. The Windscale on-line computing system. The Computer Journal, 12(3):244–250, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
3/244.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/120244.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/244.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/245.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/246.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/247.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/248.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/249.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/250.tif.

Wilkes:1968:DMA

[WN68] M. V. Wilkes and R. M. Needham. The design of multiple-access computer systems: part 2. The Computer Journal, 10(4):315– 320, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/10/4/315.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/100315.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_04/tiff/315.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_10/Issue_04/tiff/316.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/
Issue_04/tiff/317.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/318.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_04/
tiff/319.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/320.tif.

Wolfenden:1966:BRB

[Wol66] K. Wolfenden. Book review: Invariant imbedding and Time-Dependent Transport Processes. The Computer Journal, 8(4):318, January 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/8/4/318.full.pdf+html.

Wolfenden:1968:BR

[Wol68] K. Wolfenden. Book review. The Computer Journal, 10(4):351, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/4/351.full.pdf+html.

Wolfe:1969:NSN

[Wol69] M. A. Wolfe. The numerical solution of non-singular integral and integrodifferential equations by iteration with Chebyshev series. The Computer Journal, 12(2):193–196, May 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/2/193.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/120193.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/
193.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/194.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/195.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/196.tif.

Woodger:1960:IA

[Woo60] M. Woodger. An introduction to Algol 60. The Computer Journal, 3(2):67–75, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/2/67.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_02/030067.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_02/tiff/67.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
02/tiff/68.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/69.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/70.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/71.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/72.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/
73.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/74.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/75.tif.

Woodward:1961:C

[Woo61] P. M. Woodward. Correspondence. The Computer Journal, 4 (2):180, ???? 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/2/180.full.pdf+html.

Woodger:1962:BR

[Woo62] M. Woodger. Book review. The Computer Journal, 5(1):60, ???? 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/5/
1/60.full.pdf+html.

Woodward:1966:BR

[Woo66a] P. M. Woodward. Book review. The Computer Journal, 9(3):303a– 303, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/9/3/303-a.full.pdf+html.

Woodward:1966:IT

[Woo66b] P. M. Woodward. Information transmission. The Computer Journal, 9(3):303c–303, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/9/3/303-c.full.pdf+html.

Woodward:1966:PCF

[Woo66c] P. M. Woodward. Principles of coding, filtering and information theory. The Computer Journal, 9(3):303b–303, November 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/9/
3/303-b.full.pdf+html.

Woodward:1966:SMS

[Woo66d] P. M. Woodward. The synthesis of music and speech. The Computer Journal, 9(3):257–262, November 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/3/257.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/090257.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/tiff/
257.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_03/tiff/258.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_03/tiff/259.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_03/tiff/260.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_03/tiff/261.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_03/
tiff/262.tif.

Wood:1968:SCU

[Woo68] D. C. Wood. A system for computing university examination timetables. The Computer Journal, 11(1):41–47, May 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/11/1/41.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_11/Issue_01/110041.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/41.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/42.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
43.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_11/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_01/tiff/
46.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
11/Issue_01/tiff/47.tif.

Wood:1969:NFT

[Woo69a] D. Wood. The normal form theorem — another proof. The Computer Journal, 12(2):139–147, May 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/2/139.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/120139.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/
139.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_02/tiff/140.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_02/tiff/141.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_02/tiff/142.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_02/tiff/143.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_02/
tiff/144.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_02/tiff/145.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_02/tiff/146.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_02/tiff/147.tif.

Wood:1969:TLF

[Woo69b] D. Wood. The theory of left factored languages: Part 1. The Computer Journal, 12(4):349–356, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/349.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120349.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
349.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/350.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/351.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/352.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/353.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/355.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/356.
tif.

Wood:1969:TCG

[Woo69c] D. C. Wood. A technique for colouring a graph applicable to large scale timetabling problems. The Computer Journal, 12(4):317– 319, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/12/4/317.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_12/Issue_04/120317.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/317.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/318.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/319.tif.

Woodall:1969:ALR

[Woo69d] A. D. Woodall. Algorithm 43: a listed radix sort. The Computer Journal, 12(4):406, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.

co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/406.
tif.

Woodall:1970:NAS

[Woo70] A. D. Woodall. Note on Algorithms 25, 26. 25: Sort a section of the elements of an array by determining the rank of each element. 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 13(3):326, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_03/tiff/326.tif. See [Boo67f, Boo67g, Hil69].

Welsh:1967:UBC

[WP67] D. J. A. Welsh and M. B. Powell. An upper bound for the chromatic number of a graph and its application to timetabling problems. The Computer Journal, 10(1):85–86, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/10/1/85.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_10/Issue_01/100085.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
85.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/86.tif. See comment [Col67].

Wolfe:1968:SMS

[WP68] M. A. Wolfe and G. M. Phillips. Some methods for the solution of non-singular Volterra integro-differential equations. The Computer Journal, 11(3):334–336, November 1968. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/11/3/334.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/110334.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_11/Issue_03/tiff/
334.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_11/Issue_03/tiff/335.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_11/Issue_03/tiff/336.tif.

Whitby:1969:COB

[WP69] W. P. L. Whitby and R. M. Paine. Correspondence: OCR — benefits and pitfalls. The Computer Journal, 12(4):410–411, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/12/4/410.2.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/410.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/411.tif. See [Pai69].

Wallace:1967:CMA

[WR67] C. S. Wallace and B. G. Rowswell. Competition for memory access in the KDF9. The Computer Journal, 10(1):64–68, May 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/10/1/64.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_10/Issue_01/100064.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/64.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/65.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_10/Issue_01/tiff/
66.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_01/tiff/67.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_01/tiff/68.tif.

Wragg:1966:ULM

[Wra66] A. Wragg. The use of Lanczos τ-methods in the numerical solution of a Stefan problem. The Computer Journal, 9(1):106–109, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/9/1/106.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_09/Issue_01/090106.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_01/tiff/106.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_01/tiff/107.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_01/
tiff/108.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_01/tiff/109.tif.

Wright:1960:MLI

[Wri60] M. A. Wright. Mechanizing a large index. The Computer Journal, 3(2):76–83, July 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/3/2/76.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_02/030076.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_02/tiff/76.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_
02/tiff/77.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_02/tiff/78.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/79.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_02/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_02/tiff/
82.tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_02/tiff/83.tif.

Wright:1961:DCT

[Wri61a] E. P. G. Wright. Data collection and transmission. The Computer Journal, 4(2):103–109, July 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/2/103.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/040103.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/tiff/
103.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_02/tiff/104.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_02/tiff/105.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_02/tiff/106.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_02/tiff/107.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_02/
tiff/108.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_02/tiff/109.tif.

Wright:1961:MII

[Wri61b] M. A. Wright. Matching inquiries to an index. The Computer Journal, 4(1):38–41, April 1961. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/4/1/38.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_01/040036.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/38.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/39.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/40.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/41.
tif.

Wrigley:1963:AJM

[Wri63] H. E. Wrigley. Accelerating the Jacobi method for solving simultaneous equations by Chebyshev extrapolation when the eigenvalues of the iteration matrix are complex. The Computer Journal, 6(2):169–176, August 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/6/2/169.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/060169.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/
169.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_06/Issue_02/tiff/170.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_06/Issue_02/tiff/171.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_06/
Issue_02/tiff/172.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_02/tiff/173.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_02/
tiff/174.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_02/tiff/175.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_02/tiff/176.
tif.

Wright:1964:CCM

[Wri64] K. Wright. Chebyshev collocation methods for ordinary differential equations. The Computer Journal, 6(4):358–365, January 1964. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/6/4/358.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_06/Issue_04/060358.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/358.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/359.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/

tiff/360.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_06/Issue_04/tiff/361.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_06/Issue_04/tiff/362.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
06/Issue_04/tiff/363.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_06/Issue_04/tiff/364.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_06/Issue_04/
tiff/365.tif.

Wright:1966:CEC

[Wri66a] K. Wright. Correspondence: On “Error curves for Lanczos ‘selected points’ method”. The Computer Journal, 9(1):115, May 1966. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_01/tiff/115.tif. See [Kiz66b, Kiz66a].

Wright:1966:SMI

[Wri66b] K. Wright. Series methods for integration. The Computer Journal, 9(2):191–199, August 1966. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/9/2/191.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/090191.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/
191.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_02/tiff/192.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_09/Issue_02/tiff/193.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/
Issue_02/tiff/194.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_09/Issue_02/tiff/195.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_02/
tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_02/tiff/197.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_09/Issue_02/tiff/198.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
09/Issue_02/tiff/199.tif.

Wright:1966:BRB

[Wri66c] M. A. Wright. Book review: Computers in Biomedical Research. The Computer Journal, 9(1):114, May 1966. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/9/1/114.
full.pdf+html.

Watt:1962:ASL

[WY62] J. M. Watt and Andrew Young. An attempt to simulate the liver on a computer. The Computer Journal, 5(3):221–227, November 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/3/221.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_03/050221.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/221.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/222.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_03/
tiff/223.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_03/tiff/224.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_03/tiff/225.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_03/tiff/226.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_03/tiff/227.tif.

Young:1961:PDS

[YA61] A. Young and G. Almond. Predicting distributions of staff. The Computer Journal, 3(4):246–250, January 1961. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/3/4/246.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/030246.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_03/Issue_04/tiff/
246.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_03/Issue_04/tiff/247.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_03/Issue_04/tiff/248.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_03/
Issue_04/tiff/249.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_04/tiff/250.tif.

Yates:1962:CRP

[Yat62] F. Yates. Computers in research — promise and performance. The Computer Journal, 4(4):273–279, January 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/4/4/273.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/040273.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/tiff/
273.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_04/tiff/274.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_04/Issue_04/tiff/275.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_04/
Issue_04/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_04/Issue_04/tiff/277.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_04/
tiff/278.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_04/tiff/279.tif.

Yates:1963:SAA

[YGS63] F. Yates, J. C. Gower, and H. R. Simpson. A specialized Autocode for the analysis of replicated experiments. The Computer Journal, 5(4):313–319, January 1963. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/5/4/313.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/050313.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/tiff/
313.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_05/Issue_04/tiff/314.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_05/Issue_04/tiff/315.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_05/
Issue_04/tiff/316.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/317.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/318.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_04/tiff/319.tif.

Youle:1960:SFS

[You60] P. V. Youle. Simulation of full-scale multi-stage batchwise chemical plant. The Computer Journal, 3(3):150–157, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/3/150.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_03/030150.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_03/tiff/150.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/151.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/
tiff/152.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_03/tiff/153.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/154.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_03/tiff/155.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_03/Issue_03/tiff/156.tif;   http://

www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/
tiff/157.tif.

Young:1963:IUE

[You63a] A. Young. The impact on universities of the expansion in their computer facilities. The Computer Journal, 5(4):294–296, January 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/4/294.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_05/Issue_04/050294.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
05/Issue_04/tiff/294.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_05/Issue_04/tiff/295.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_05/Issue_04/
tiff/296.tif.

Young:1963:BR

[You63b] Andrew Young. Book review. The Computer Journal, 5(4):262, ???? 1963. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/5/4/262.1.full.pdf+html.

Young:1968:BRa

[You68a] A. Young. Book review. The Computer Journal, 11(1):56, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
1/56.full.pdf+html.

Young:1968:BRb

[You68b] A. Young. Book review. The Computer Journal, 11(3):298, ???? 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/11/
3/298.full.pdf+html.

Yates:1960:GPA

[YS60] F. Yates and H. R. Simpson. A general program for the analysis of surveys. The Computer Journal, 3(3):136–140, October 1960. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/3/3/136.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_03/Issue_03/030136.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
03/Issue_03/tiff/136.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_03/Issue_03/tiff/137.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_03/Issue_03/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_03/Issue_03/tiff/139.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_03/Issue_03/tiff/140.
tif.

Yates:1961:ASP

[YS61] F. Yates and H. R. Simpson. The analysis of surveys: processing and printed the basic tables. The Computer Journal, 4(1): 20–24, April 1961. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/4/1/20.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_04/Issue_01/040020.sgm.
abs.html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_04/Issue_01/tiff/20.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_04/Issue_01/tiff/21.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_04/Issue_
01/tiff/22.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_04/Issue_01/tiff/23.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_04/Issue_01/tiff/24.
tif.

Yule:1968:EHA

[Yul68] A. P. Yule. Extensions to the heuristic algorithm for University timetables. The Computer Journal, 10(4):360–364, February 1968. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
4/360.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/100360.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
04/tiff/360.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_04/tiff/361.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_04/tiff/362.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_04/tiff/363.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_10/Issue_04/tiff/364.tif.

Zangwill:1967:MFC

[Zan67] W. I. Zangwill. Minimizing a function without calculating derivatives. The Computer Journal, 10(3):293–296, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/10/
3/293.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_10/Issue_03/100293.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_
03/tiff/293.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_10/Issue_03/tiff/294.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_10/Issue_03/tiff/295.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
10/Issue_03/tiff/296.tif.

Zaremba:1969:SAI

[Zar69] W. A. Zaremba. A syntax for Algol input/output formats. The Computer Journal, 12(4):342–348, November 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://comjnl.oxfordjournals.org/content/12/4/342.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/120342.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/
342.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_04/tiff/343.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/344.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/
Issue_04/tiff/345.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/346.tif;   http://
www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
tiff/347.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/348.tif.